### v1

# Channel Choice and Customer Value

Abstract
The rapid growth of digital shopping channels has prompted many traditional retailers to invest in e-commerce websites and mobile apps. While prior literature shows that multichannel customers tend to be more valuable, it overlooks how the reason for adopting a new channel may shape post-adoption behavior. Using transaction-level data from a major Brazilian pet supplies retailer, we examine how adoption of online shopping—by previously offline-only customers—affects post-adoption spend, profitability, and channel usage. We study four distinct adoption pathways: organic adoption, adoption due to the COVID-19 pandemic, Black Friday promotions, and a newly launched loyalty program. We find that although all adopter groups increase their spending relative to offline-only customers, post-adoption behavior differs based on the reason for adoption. COVID-19 adopters behave similarly to organic adopters in terms of spending, but show greater offline stickiness consistent with consumer inertia and habit theory, yielding higher profits due to higher offline margins. In contrast, promotion-driven adopters spend less post-adoption due to forward buying and exhibit lower profitability. Our findings caution against treating all multichannel customers as equal and highlight the importance of incorporating behavioral theories into forecasting and targeting strategies. Firms should account for adoption motives when evaluating channel and promotional investments.
Keywords: Retailing, Multichannel, Online Shopping, Customer value
1 Introduction
Online shopping has been growing in popularity over the last two decades, with the proliferation of computers and smartphones. In response, many grocery and retail stores like Walmart, Target, Costco, etc., which earlier only operated physical retail stores, now sell their products through both physical stores and e-commerce channels such as websites and mobile apps. Given the size of the retail industry and the growth of digital channels (especially the acceleration of the adoption of online shopping during the COVID-19 pandemic), quantifying the effect of the adoption of digital shopping channels is important for managers and firms (Goyal 2023; Bloomberg 2020).
In this paper, we consider two related research questions. First, we examine whether and how the adoption of digital shopping channels (e-commerce website and/or mobile app) of a retailer affects the purchase behavior of consumers who were previously purely offline-only (brick-and-mortar) shoppers. We seek to document the impact of the adoption of online shopping on three metrics of interest for managers: (1) Spend, (2) Profitability, and (3) Stickiness to the offline channels post-adoption. The first metric, spend, quantifies the total revenue/sales (across offline and online channels) generated by the customer after the adoption of online shopping, the second metric quantifies the profits (sales minus cost) generated by the customer, and the third metric denotes the extent to which consumers use offline channels even after adopting online shopping. Second, we are interested in understanding whether and how post-adoption shopping behavior differs depending on consumers’ different reasons for adopting online shopping. This is important from a managerial perspective – if the changes to consumers’ purchase behavior (post-online adoption) are largely driven by the motivations/reasons for consumers choosing to adopt online shopping, then it will be challenging to extrapolate the behavior of organic adopters (who adopt online shopping without any external incentives) to other consumers who are incentivized to adopt by marketing campaigns or other external economic or environmental factors i.e., different motivations compared to organic adopters. Furthermore, if the differences in post-adoption behavior relative to organic adopters depend on the type of external factor (i.e., macroeconomic or environmental vs. firm-driven promotions); forecasting post-adoption behavior would be even more complicated. On the other hand, if the consumers who adopt online shopping due to different reasons behave similarly post-adoption, then managers can be more confident in extrapolating the behavior of those who adopt organically to a broader set of adopters (and be less concerned about the reason for the adoption of online shopping).
While a large stream of prior research suggests that multichannel consumers (defined as consumers who buy from multiple channels including catalogs, mobile apps, stores, websites, etc.) tend to spend more than single-channel consumers (Neslin and Shankar 2009), it is not obvious if the earlier results translate directly to our setting for three reasons. We briefly discuss the reasons here and refer the reader to Table 1 for a more detailed comparison to previous literature. First, they do not specifically study the effect of the adoption of online shopping from previously only purchasing from brick-and-mortar stores. Studying the physical store to online (e-commerce or mobile app) transition is both novel and interesting because many major retailers started out with brick-and-mortar stores, and only over the last couple of decades have been investing more heavily in developing the e-commerce and mobile app channels. E-commerce websites and mobile apps, as channels, are potentially quite different from other channels like mail order catalogs, phone shopping, or interacting with salespeople because both these channels reduce search costs, and it is easy for customers using these channels to visit other retailers’ websites and apps to comparison shop across retailers. Second, the vast majority of these papers adopt one of two approaches – (a) a cross-sectional analysis (Kumar and Venkatesan 2005; Thomas and Sullivan 2005; Kushwaha and Shankar 2013) comparing the spend of single-channel shoppers (any single channel, not necessarily just physical stores) vs. multichannel shoppers (which can include physical stores and online channels, but also mail-order catalogs, ordering through phone, salespeople, etc.), or, (b) a within-customer analysis (Venkatesan and Ravishanker 2007; Ansari, Mela, and Neslin 2008) comparing the pre- and post-adoption periods of customers who start using multiple channels from being single channel only (where again, these channels include websites, catalogs, physical stores etc.). However, cross-sectional comparisons cannot answer the question of how customer behavior changes after the adoption of multiple channels, and within-customer analyses require the strong assumption that the entire difference between the before and after periods can be fully attributed to the adoption of multiple channels, without any other time-varying factors. More recently, some studies have used Difference-In-Differences (DID) analyses to study the effects of the adoption of mobile apps (Narang and Shankar 2019; Wang, Malthouse, and Krishnamurthi 2015). However, before the adoption of the app, customers in these studies could have already been using the retailer’s website to shop online. As such, their findings may not generalize to our context, where the transition in focus is from physical stores to online channels. Third, the previous studies only examine customers who have voluntarily/organically chosen to use multiple channels for shopping or have chosen to adopt the firm’s mobile app. They completely ignore different channel adoption pathways and implicitly assume that the benefits of multichannel customers are uniform regardless of why they chose to adopt multiple channels or a particular channel. There are a couple of important theoretical reasons to expect differences between (a) organic adopters and externally incentivised adopters and (b) different types of external incentives i.e. macro-environmental shocks vs. firm driven promotions. Consumer inertia and habit theory (Verplanken and Wood 2006; Wood and Neal 2009) would predict that adoption driven by external shocks (e.g., macro shocks or firm-driven promotions) may lead to temporary disruptions in consumer behavior (i.e. adopting a new channels and using multiple channels to shop), but without reinforcement, consumers may revert to prior routines (i.e. not use the newly adopted channels much). Next, for promotion-induced adoption in particular, the theory of forward buying and demand distortion may apply. Promotions incentivizing online adoption may also induce forward buying, leading to reduced post-promotion consumption (Neslin, Henderson, and Quelch 1985; Ailawadi and Neslin 1998), in turn leading to lower revenues for the firm post-adoption. Thus, we extend the previous literature by (a) studying a particularly important type of transition from single-channel to multi-channel (i.e. offline-only to offline-plus-online) and (b) by examining the behavior or consumers who adopted online shopping due to different reasons, and challenging the assumption that all multichannel customers are uniformly beneficial and behave the same after adopting a new channel regardless of their reason for adoption.
We use data from a large pet supplies retailer in Brazil with a large network of brick-and-mortar stores and an online presence (website and mobile app) to answer our research questions. Our data consists of all consumer purchases across all channels for the period of 2019-2024. This time period allows us to construct four groups of adopters of online shopping who previously only shopped at the firm’s physical stores – (1) covid adopters who started shopping online during the onset of COVID-19, (2) Black Friday adopters who adopted online shopping due to deep discounts available online during Black Friday, (3) Loyalty Program adopters who adopted online shopping during the launch of a loyalty program which had to be activated on the retailer’s mobile app, and (4) organic adopters who started shopping online fully of their own volition, without any external stimuli. In addition, we also consider a base control group of non-adopters or offline-only customers (consumers who did not adopt online shopping during our observation period). We then construct consumer-level monthly panel data to quantify the revenue, profitability, and share of spend in the offline and online channels. Since we have four groups of adopters who chose to adopt for different reasons, in addition to non-adopters, we can address both our research questions – quantify the impact of the adoption of online shopping on consumers’ spend, profitability, and stickiness to offline channels, as well as comment on whether and how the different reasons for adoption affects the post-adoption behavior of adopters.
We now discuss our main findings. First, we find that relative to offline-only customers, adopters of online shopping, have higher spend and profitablility in the post-adoption period. However, the post-adoption behaviors of the different types of adopters diverge significantly from each other across the three key dimensions: total spend, profitability and the share of offline/online spending. While covid adopters spend similarly relative to organic adopters post-adoption, both promotion-driven adopters, Black Friday adopters and Loyalty Program adopters, spend less. This reduced spend is driven largely by forward buying during the promotional period, meaning they stocked up and then reduced future purchases, which doesn’t occur in the case of covid adopters. In terms of the share of offline/online spend post-adoption, consistent with consumer inertia and habit theory, all three adopter groups motivated by external stimuli (covid adopters, Black Friday adopters and Loyalty Program adopters) have a lower online spend share post-adoption relative to organic adopters. A combination of demographic characteristics and consumer inertia and habit theory also explain differences in how this share evolves over time. We find that covid adopters show a slower growth in online spend share over time. This is mainly driven by the fact that covid adopters are older and thus are more set in their habits. Black Friday adopters on the other hand, are younger and show a faster growth rate in the online spend share relative to organic adopters, although they still take time to catch up to the same levels. There is no difference in trends for the online spend share between Loyalty Program adopters and organic adopters. This may be due to the fact that the two adopter groups are quite similar in terms of demographics, and also the fact that the Loyalty Program involves monthly online activation, and therefore there is some reinforcement for the habit of online shopping for Loyalty Program adopters. The post-adoption profits for the different adopter groups are a function of both spend and the share of offline/online share. We find that despite similar spend, covid adopters are more profitable post-adoption, than organic adopters in our setting. This is because of their higher spend share in the offline channel, where profit margins are higher in our setting. In contrast, Black Friday adopters and Loyalty Program adopters are less profitable due to their lower overall spending in the post-adoption period.
Our findings suggest that the reason for a consumer’s decision to adopt online shopping significantly shapes their long-term behavior. Consumers who adopt online shopping organically behave very differently from those motivated by external events or firm-led promotions, due to a combination of their demographics, consumer inertia, and forward buying during promotions. Our findings have important managerial implications. For example, had the retailer assumed that Black Friday adopters would behave like organic adopters, it would have significantly underestimated the payback period for discounts offered during the promotion. While projections based on organic adopters would suggest a break-even point at 13 months, the actual break-even occurred only after 39 months. Even when overall spending appears similar—as in the case of covid adopters—differences in channel usage (online vs. offline) can have substantial implications for profitability. Firms should avoid using organic adopters as the benchmark when predicting the behavior of customers adopting a new channel for other reasons. More broadly, our empirical results indicate that managers can use their knowledge of expected customer demographics, and apply the theories of consumer inertia and forward buying to forecast spend, profit and channel usage outcomes in other channel adoption scenarios. Tailoring strategies to the unique characteristics of each type of adopter group is essential for maximizing both customer value and long-term profitability.
In summary, our paper makes both substantive and managerial contributions to the literature on channel choice and consumer value. From a substantive perspective, we show that the adoption of online shopping from only using physical stores, either voluntarily or prompted by external factors like a macro-environmental shock or firm-driven promotions, causes consumers to increase their total spend with the firm relative to consumers who stay offline-only. More interestingly, we show that consumers who adopt online shopping due to different reasons, have different post-adoption outcomes in terms of their spend, channel usage and profits, and these differences can be explained using demographics, consumer inertia theory, and forward buying during promotions. From a managerial perspective, we not only quantify the impact of the adoption of online shopping on revenues and profits but also provide a cautionary tale against naively extrapolating the profitability of one group of online shopping adopters to a broader/different set of adopters. We also provide theoretical explanations of why post-adoption behaviors of different adopter groups differ, which managers can use to forecast outcomes in other settings.
2 Related Literature
Paper Setting Channels Methodology Type of adopters
Our Paper Pet Supplies Physical stores only to Diff-in-diff Organic +
adding website and/or multiple external
mobile app stimuli
Narang and Shankar (2019) Video games Physical stores + website Diff-in-diff Organic only
and Electronics to adding mobile app
Montaguti, Neslin, and Valentini (2016) Books Mail order, fax, phone, Cross-sectional Organic +
(subscribers only) SMS, Internet, external stimulus
and bookstores
Wang, Malthouse, and Krishnamurthi (2015) Online grocery Website to adding Diff-in-diff Organic only
mobile app
Kushwaha and Shankar (2013) Multiple categories Website and catalog Cross-sectional Organic only
Ansari, Mela, and Neslin (2008) Apparel Website and catalog Within-customer Organic only
Venkatesan and Ravishanker (2007) Apparel Full price stores, Within-customer Organic only
Discount Stores,
and website
Kumar and Venkatesan (2005) Technology Salespersons, direct mail, Cross-sectional Organic only
telephone and website
Thomas and Sullivan (2005) Retail Catalog, physical stores, Cross-sectional Organic only
website
Table 1: Prior Empirical Literature on Multichannel Shopping in Retail Settings
There is a large empirical literature in marketing studying the effects of multichannel shopping vs. single-channel shopping. We provide a brief overview of the most closely related papers here in Table 1 and discuss how our paper differs from and contributes to this literature in the rest of this section. We refer readers to Neslin et al. (2006); Neslin and Shankar (2009); Verhoef (2012) for more comprehensive reviews of the broader multichannel literature.
As shown in Table 1, our paper differs from the earlier literature both substantively and methodologically. First, from a substantive perspective, we study a specific type of single-channel to multichannel transition that has not been studied before – from shopping at physical stores only to shopping both at physical stores and online through the retailer’s website/mobile app. Earlier findings from the multichannel literature largely focus on the transition from any single channel to some combination of multiple channels (including physical stores, catalogs, phones, mail orders, salespeople, mobile apps, and websites).
Studying the physical store to e-commerce transition is of particular interest to traditional brick-and-mortar retailers (Goyal 2023; Bloomberg 2020) like Walmart, Target, and Costco since these channels (websites and mobile apps) represent a key area of growth for them. However, since the earlier literature does not explicitly study this transition (and rather considers a broader set of channels), those findings may not generalize to this context. Indeed, there are some important differences among these different channels. For instance, websites and apps provide easy search and comparison between products within a retailer and also between competing retailers because it is easy to jump between retailer websites/apps on your device. This is not the case for other channels that have been studied before, like catalogs, salespersons, or phone shopping. Also, in many cases, prices on a retailer’s website or mobile app are lower than the prices in physical stores, which is not the case for other channels. Thus, the findings from other channels may not fully translate to the channels we study here.
Moreover, the setting of our study is a fast-moving consumer packaged goods retailer (pet supplies), whereas much of the earlier literature uses different settings like books, apparel, video games, electronics, technology, etc. These earlier settings differ from ours on many dimensions, e.g., they tend to be less frequently purchased, do not have as many competing stores, and the nature of the product is more hedonic (rather than utilitarian), which has been found to be important for the multichannel effect (Kushwaha and Shankar 2013).
Most importantly, the existing literature overlooks the heterogeneity in channel adoption pathways, implicitly assuming that the benefits derived from multichannel customers are consistent, irrespective of the motivations underlying their adoption decisions. However, theoretical frameworks suggest important distinctions between (a) organic adopters and those induced through external incentives, and (b) different types of external stimuli—namely, macro-environmental shocks versus firm-initiated promotions. Consumer inertia and habit theory (Verplanken and Wood 2006; Wood and Neal 2009) suggests that externally triggered adoption events—such as those prompted by environmental disruptions or marketing campaigns—may temporarily disrupt existing behavioral routines, resulting in short-term channel experimentation. Without reinforcement, however, such consumers are likely to revert to prior habits, thereby potentially limiting the long-term impact of multi-channel adoption. Moreover, when adoption is driven by promotions, the literature on forward buying and demand distortion (Neslin, Henderson, and Quelch 1985; Ailawadi and Neslin 1998) raises concerns about post-promotion gains for firms. Specifically, promotion-induced online channel adoption may lead to intertemporal demand shifts rather than incremental sales, reducing post-promotion revenues and potentially diminishing the value of these new multichannel users.
Thus, this paper extends the literature from a substantive perspective in two key ways: First, we focus on a key type of channel transition: when customers move from shopping only in offline stores to using both offline and online channels. Second, we examine how post-adoption behavior differs depending on why consumers started using the online channel— whether organically or due to external factors like promotions or macro-environmental shocks, and provide theoretical explanations for the differences. By doing so, we challenge the implicit assumption in prior research that all multichannel customers behave similarly post-adoption and are equally valuable to firms, regardless of how or why they adopted a new channel [^1] Like us, Montaguti, Neslin, and Valentini (2016) also consider an external stimulus for multichannel adoption (marketing campaign). However, the methods, goals, and estimands of interest of the two papers are quite different – we use difference-in-differences to study the change caused by adoption of online shopping, whereas they use a cross-sectional comparison to measure the effectiveness of a marketing campaign nudging consumers to use multiple channels. The substantive settings also differ – we focus on the transition from using only physical stores to physical stores + e-commerce in a competitive pet groceries setting, whereas they use data from a bookstore with a subscription model and focus on subscribers who use single or multiple channels (mail order, fax, phone, SMS, the Internet, and bookstores). Most importantly, they do not study the differences between organic multichannel customers vs. those who were nudged by the marketing campaign, which is a key area of focus in our paper..
Next, from a methodological perspective, prior research has considered three broad approaches to documenting the value of adopting new channels: cross-sectional, within-user, and DID. Below, we discuss each stream and explain how our research is distinct from the earlier work.
Cross-sectional: Papers in this stream (Thomas and Sullivan 2005; Kumar and Venkatesan 2005; Kushwaha and Shankar 2013; Montaguti, Neslin, and Valentini 2016) typically compare the revenue/profits of two distinct groups of consumers to document the value of multichannel shopping – those who use multiple channels (among catalog, brick-and-mortar, salespersons, direct mail, telephone sales, SMS or the internet) vs. those who use a single channel. A consensus finding in this literature is that multichannel customers tend to be more valuable than single-channel customers. However, a potential drawback of this methodology is that multichannel consumers are self-selected and could be systematically different from those using single channels. As such, cross-sectional comparisons cannot answer the question of how customer behavior changes after the adoption of multiple channels.
Within Customer: Papers in this stream (Venkatesan and Ravishanker 2007; Ansari, Mela, and Neslin 2008) typically compare how a customer’s purchase behavior changes after the adoption of multiple channels (which could be different types of stores, catalogs, internet etc.) compared to their behavior when they were a single-channel customer. The key takeaway here is that customers spend more after becoming multichannel compared to when they were using only a single channel. However, there are important drawbacks to this type of analysis. First, without an appropriate control group of non-adopters, we need to assume that the entire difference between the before and after periods can be fully attributed to the adoption of multiple channels, and there are no other time-varying factors affecting purchasing behavior – which is a strong assumption. Second, it is not obvious if the estimated gains of the adopters can be extrapolated to non-adopters because the adopters self-selected into the use of multiple channels.
Difference-in-Differences (DID): This more recent stream of literature tries to address some of the concerns with the above two approaches by using a DID methodology and using non-adopters as a control group. However, the papers in this stream (Wang, Malthouse, and Krishnamurthi 2015; Narang and Shankar 2019; Kim, Wang, and Malthouse 2015; Gu and Kannan 2021) focus only on the adoption of mobile apps rather than a broader set of channels. Specifically, they find that customers who adopted a firm’s mobile app (from previously using physical stores, website, or both) increase their purchases post-adoption [^2] Gu and Kannan (2021) shows the opposite result, i.e., a decrease in purchases post-app adoption for a hotel chain. While our paper also uses a DID method, our substantive focus is very different from these papers, as explained previously.
Finally, while the literature on multichannel vs. single-channel shopping described above is most closely related to our paper, there are a couple of other strands of the marketing literature that are also somewhat related. First, there is a literature that has studied the role of channels in non-retail settings, e.g., quantifying the effects of the adoption of new channels (like streaming services) on the downloading and owning in the context of entertainment goods like movies and music (Datta, Knox, and Bronnenberg 2018; Aguiar and Waldfogel 2018). Second, there is also literature that studies related questions like the value of physical stores in an omnichannel environment (Zhang, Chang, and Neslin 2022) and the effects of customer experiences in different channels on retention (Chang and Zhang 2016). See Zhang et al. (2010) and Cui et al. (2021) for a summary of the findings and challenges in the broader literature on omnichannel retail strategy.
3 Setting and Data
3.1 Setting
Our data come from one of the largest Brazilian pet supplies retailers, which sells a wide variety of products for pets, such as pet food and snacks, medicines/pharmacy items, hygiene and grooming products, and various pet toys and accessories. The firm has both offline and online channels. The retailer operates more than 200 physical stores in Brazil, which serve as the offline channel, and has a website as well as a mobile app, which serve as the online channel. Pet supplies is a large and competitive industry with about \$250 billion in annual revenue globally (Fortune 2024).
3.2 Data
Our data include transaction-level records of all individual customers from both online and offline channels during the 63-month period from January 1st, 2019 to March 31st, 2024. Each transaction record is uniquely identified by an order number with one or more purchased items. For each purchased item, we have information on the unit price, quantity purchased, brand, purchased channel (online/offline), subcategory, and category (e.g., food, hygiene).
The firm tracks customer purchases using a National Identification Number (similar to a Social Security Number), which is unique to each individual in Brazil. Customers who purchase online (either through the website or the app) must be logged in to their profile, and therefore, all online purchases can be traced back to the National ID. Customers are also asked for their National ID when making an in-store purchase; however, this is not mandatory. Discussions with the retailer revealed that while most offline transactions are also linked with a customer profile, about a quarter of in-store transactions are not linked to a National ID. This number has remained fairly steady during the entire observation period and was not affected by COVID-19 or other marketing events [^3] Because the number of offline transactions not linked to customer IDs remained similar through our observation period and exhibited no significant shifts in response to any of the events that spurred online adoption (including COVID-19), it is unlikely that this issue systematically biases our findings on the adoption of online shopping.
A potential concern with any channel analysis is that the results could be driven by assortment differences across the two channels. To ensure that this is not the case, we exclude data from four product categories that were not available in both channels for the entirety of the observation period; see Table A1 in Web Appendix $\S$§A for a summary of these categories and their availability information. Together, sales from these categories account for only 0.63% of the total sales during the observation period; as such we do not expect these to play any significant role in the findings.
The total number of consumers who shopped at least once during our sample period is 5,934,248444This is the set of consumers who purchased at least one item from one of the non-excluded categories.. For each of these consumers, we aggregate the transaction-level data into customer-month-level panel data over the customer’s entire purchase history within the observation period (January 2019 to March 2024), starting from his or her initial purchase month and ending with the last observed purchase month. For example, if a customer purchased for the first time in our data in August 2019 and the last observed purchase is in July 2020, then the panel for this customer spans 12 months, from August 2019 to July 2020. If there are intermediate months in which a customer did not make any purchase, the values for the monthly purchase-related variables are set to zero.
We now discuss the summary statistics of our data. In $\S$3.2.1, we summarize the consumer-level demographic data, and in $\S$3.2.2, we summarize the monthly purchase behaviors.
3.2.1 Time Invariant Consumer Attributes
Table 2: Summary Statistics of Customer-Level Categorical Demographic Variables
Gender Count Percentage (%)
Female 2,351,011 39.62
Male 1,521,157 25.63
No Information 2,062,080 34.75
Total 5,934,248 100.00
Table 3: Summary Statistics of Customer-Level Demographic Variables
Variable Mean SD 25th 50th 75th (Min, Max) Count Missing (%)
Age 40.96 12.75 31 39 49 (7, 99) 4,302,771 27.49
HouseholdIncome 1854.37 1471.55 824.54 1358.21 2448.73 (8, 30826) 4,377,957 26.23
PanelLength 19.76 19.39 1 14 34 (1, 63) 5,934,248 0.00
We first summarize the time-invariant attributes of the consumers in the data. We have data on five consumer-specific variables:
  • $\bullet$
    $Gender_{i}$: Categorical variable denoting $i^{\prime}s$ gender (Male, Female, or No Information).
  • $\bullet$
    $Age_{i}$: Customer $i$’s age.
  • $\bullet$
    $HouseholdIncome_{i}$: Customer $i^{\prime}s$ average monthly household income.
  • $\bullet$
    $PanelLength_{i}$: Period spanning the first and last purchases for customer $i$ (in months).
  • $\bullet$
    $Zipcode_{i}$: The zipcode of customer $i$.
We present the summary statistics for demographic variables in Tables 2 and 3. Out of the 5,934,248 customers, 2,351,011 (39.6%) are Female, 1,521,157 (25.6%) are Male, and the rest (34.8%) do not report gender information. The median consumer is 39 years old. In addition, the mean and median of monthly household incomes are 1854.37 and 1358.21 Brazilian Real, respectively. Next, we see that the average and median panel lengths are 19.76 and 14 months respectively, which suggests that there is consumer entry as well as churn during the observation period. Finally, we observe that customers in our data sample come from 300,774 distinct zip codes across all of the 28 states of Brazil (though 20.4% of total customers do not provide the zip code information).
3.2.2 Customer Monthly Purchase Behaviors
Table 4: Customer-Month Summary Statistics
Variables mean std min 25% 50% 75% max count
Transactions across both Channels
Spend 303.02 759.25 0 0 0 374.92 2,340,617.16 123,748,452
Quantity 2.44 10.27 0 0 0 2.00 44,158.00 123,748,452
Order 0.61 1.03 0 0 0 1.00 430.00 123,748,452
UniqueItem 1.42 2.93 0 0 0 2.00 4,191.00 123,748,452
UniqueBrand 1.15 2.09 0 0 0 2.00 332.00 123,748,452
UniqueSubcategory 1.00 1.70 0 0 0 2.00 64.00 123,748,452
UniqueCategory 0.85 1.35 0 0 0 1.00 18.00 123,748,452
Transactions in Offline Channel
Spend 195.07 631.18 0 0 0 125.93 2,340,617.16 123,748,452
Quantity 1.78 9.03 0 0 0 1.00 44,158.00 123,748,452
Orders 0.43 0.87 0 0 0 1.00 430.00 123,748,452
UniqueItems 1.11 2.72 0 0 0 1.00 4,191.00 123,748,452
UniqueBrands 0.90 1.94 0 0 0 1.00 332.00 123,748,452
UniqueSubcategories 0.78 1.59 0 0 0 1.00 64.00 123,748,452
UniqueCategories 0.66 1.26 0 0 0 1.00 18.00 123,748,452
Transactions in Online Channel
Spend 107.95 421.24 0 0 0 0 276,399.66 123,748,452
Quantity 0.66 4.68 0 0 0 0 4,800.00 123,748,452
Orders 0.18 0.56 0 0 0 0 130.00 123,748,452
UniqueItems 0.32 1.14 0 0 0 0 141.00 123,748,452
UniqueBrands 0.27 0.88 0 0 0 0 44.00 123,748,452
UniqueSubcategories 0.24 0.75 0 0 0 0 22.00 123,748,452
UniqueCategories 0.22 0.64 0 0 0 0 13.00 123,748,452
We quantify consumers’ purchase behavior on three dimensions: Volume, Frequency, and Variety.
  • $\bullet$
    Volume of purchase is operationalized using two variables – (1) $(\textit{Spend}_{it})$, which is the revenue generated from the customer $i$ in month $t$, and (2) $(\textit{Quantity}_{it})$, the total number of items purchased by customer $i$ in month $t$.
  • $\bullet$
    Frequency is measured by the number of unique orders per month $(\textit{Orders}_{it})$. This metric focuses on purchase frequency without considering the value of purchases, which differs from volume metrics.
  • $\bullet$
[…]
To preserve the confidentiality of the data provider, we mask the true spend value ($\textit{Spend}_{it}$) by multiplying it by an undisclosed scalar. All the descriptive and model estimation results are thus shown in terms of Masked Currency Units (MCU). The rest of the variables listed above are shown and used as is, without any masking.
Table 4 presents customer-month-level summary statistics of all variables discussed above for three cases: (1) Transactions across both channels, (2) Transactions in the offline channel, and (3) Transactions in the online channel. Table 4 shows the mean statistic of each variable, measured at the customer-month-channel level, which is calculated as $\frac{\sum_{i=1}^{N}\sum_{t=1}^{T_{i}}X_{itc}}{\sum_{i=1}^{N}T_{i}}$, where $X_{itc}$ is the customer-month-channel level variable, $T_{i}$ is the panel length of customer $i$, and $N$ is the total number of customers, which is 5,934,248. $\sum_{i=1}^{N}T_{i}$ is the total number of observations (referred to as count in Table 4), which is 123,748,452.
We find that, on average, consumers spend more offline than online. It is important to note that although the unconditional mean of monthly offline spend (195.07 MCU) is higher than online spend (107.95 MCU), we cannot conclude that, on average, customers spend more money offline than online. This is because some customers may only access one sales channel, or they strategically select the channel that benefits them the most, or choose different channels for different types of purchases. We will discuss these ideas in detail later in the paper. Further, we find that all the variety and frequency variables are also higher for offline purchases compared to online purchases. All the purchase metrics have over 50% of observations as zero, which suggests that purchase frequencies are lower than once per month for a large fraction of consumers, though some consumers purchase very frequently and buy a large variety of products.
4 Descriptive Analysis
As a first step, we conduct some descriptive analyses to examine (a) cross-sectional differences in demographics, monthly spending, and other purchase behaviors of multichannel consumers, online-only consumers, and offline-only consumers; and (b) within-customer changes to monthly spending and other purchase behaviors after adopting online shopping. We highlight the main takeaways from each analysis here and refer readers to Web Appendix B for additional details.
First, we observe that multichannel consumers are more likely to be female compared to customers who are online-only or offline-only. Of the three groups, online-only consumers are the youngest, and offline-only consumers are the oldest. Multichannel consumers have the highest income, and offline-only consumers have the lowest income. Next, we run simple cross-sectional regressions for the outcome variables, spend, quantity, unique items, brands and categories, on the explanatory variable of consumer-type (online-only, offline-only and multichannel), controlling for demographics and month fixed effects, to measure cross-sectional differences in spend and other purchase behaviors. We find that multichannel consumers spend more (and also purchase higher quantities, more unique items, brands, and categories) relative to both offline-only and online-only consumers. These patterns align with previous literature (Kumar and Venkatesan 2005; Thomas and Sullivan 2005; Kushwaha and Shankar 2013) documenting that multichannel consumers are more valuable than single-channel consumers (though the set of channels studied in that literature is different).
Next, we conduct a within-consumer analysis to examine whether consumers who were initially only shopping offline increase their spending after adopting online shopping. We do see a significant increase in monthly spending (by about 39%) post-adoption  [^5] Referring to Table A6 in Web Appendix B.2, this number is given by the incremental spend after online adoption (124.2 MCU) divided by the pre-adoption average spend (318.4 MCU). These results are consistent with earlier work (studying different channel transitions) that shows that consumers who transition to shopping on multiple channels from previously using only a single channel tend to spend more (Venkatesan and Ravishanker 2007; Ansari, Mela, and Neslin 2008) post multichannel adoption. We also see similar increases in quantity, unique items, brands, and categories.
While these descriptive analyses confirm existing studies, they nevertheless have several shortcomings. The cross-sectional analysis does not measure changes in behavior, while the within-customer analysis assumes that the before-after changes we observe are entirely attributable to the adoption of online shopping (since there is no control group of non-adopters). Moreover, these analyses do not shed light on whether the different reasons for the adoption of online shopping affect consumers’ purchase behavior in the post-adoption period differently. We address these issues and adopt a more comprehensive empirical framework to answer our research questions in the next two sections.
5 Firm’s Problem and Sample Construction
We start by formalizing the firm’s problem. As discussed in $\S$1, our first goal is to quantify how the adoption of digital shopping channels (e-commerce website and/or mobile app) affects the purchase behavior of consumers who were previously purely offline-only (brick-and-mortar) shoppers in retail markets. Specifically, we seek to document the impact of the adoption of online shopping on the following three metrics: (1) Total Spend, (2) Profitability, and (3) Stickiness to the offline channels post-adoption. Formally, we are interested in the Average Treatment Effect on Treated (ATT) of adopting online shopping on these three metrics. A second goal is to understand whether the different reasons for adopting online shopping affect consumers’ post-adoption purchase behavior, and to what extent. To that end, we consider four types of reasons behind online adoption: (1) a macro-environmental shock (COVID-19), (2) a short-term promotion (Black Friday sales), (3) a long-term loyalty program, and (4) organic adoption, not driven by any external events. The first event is outside the firm’s control, whereas the latter two were firm-driven marketing activities. For each event, we first compare consumers who adopted online shopping adopters with a control group of offline-only consumers to study the effect of online adoption. Next, we compare consumers who adopted online shopping due to the event (i.e., event adopters) with those who adopted organically (without external incentives) around the same time frame as the event (i.e., organic adopters) to understand how these different external events affect post-adoption behavior relative to organic adoption. In the process, we also seek to provide theoretical underpinnings for how and why the different reasons for online adoption affect post-adoption purchase behavior differently.
We first discuss the sample construction for the three focal events in $\S$5.1 and then present some summary statistics of the three samples in $\S$5.2.
5.1 Sample Construction
For each event, we first define three customer cohorts – event adopters, organic adopters, and offline-only customers. We then establish the pre-adoption, adoption, and post-adoption periods for each focal event. Specifically, for a given event:
  • $\bullet$
    The pre-adoption period is defined as the common period during which all three groups of customers purchase only through physical stores (i.e., before the onset of the focal event).
  • $\bullet$
    The adoption period is defined as the time during which adopters started purchasing online. Because each event has two sets of adopters, we have two adoption periods:
    • $\bullet$
      Adoption period for event adopters: This is the period during which the focal event took place, and induced customers to adopt online shopping. For instance, in the case of COVID-19, this is the first month and a half since the onset of COVID-19. Similarly, for the Black Friday event, the adoption period for event adopters is simply the day of the Black Friday sale.
    • $\bullet$
      Adoption period for organic adopters: Since organic adopters serve as a control group for event adopters, for each event, we choose consumers who organically adopted online shopping just before the onset of the focal event, and this period varies based on the event.
  • $\bullet$
    The post-adoption period refers to the months after the adoption periods of both event adopters and organic adopters.
See Table 5 for an overview of these periods for each of the three events. In our estimation sample for each event, we only include data from the pre- and post-adoption periods for that event. We exclude the adoption periods for both event and organic adopters from the estimation sample because consumers are in flux during that time (where some adopters have adopted while others have not). Furthermore, by excluding these observations, we avoid capturing any abnormal purchase behaviors associated with the initial adoption of online shopping and the onset of specific events (i.e., COVID-19, Black Friday, Loyalty Program). As such, it allows us to cleanly identify the post-online-adoption difference in customers’ purchase behaviors (Gu and Kannan 2021). We now clarify a few points on the choice of these three periods for each event below.
  • $\bullet$
    COVID-19 event: We use one year of data before the adoption period as the pre-adoption period (the calendar year of 2019). For post-adoption period, we use all the data from May 2020 till April 2023666Pilot studies for the firm’s Loyalty program started running from May 2023; so we cut off the post-adoption period with April 2023..
  • $\bullet$
    Black Friday event: We use eight months of data, starting from January 2019 to October 2020, as the pre-adoption period. For the post-adoption period, we use data from December 2019 (following the Black Friday event) until the next Black Friday.
  • $\bullet$
    Loyalty Program event: We use the data from five months before the adoption period, and all the data available (from the end of the adoption period till March 2024) as the post-adoption period.
Table 5: Pre-adoption, Adoption, and Post-adoption Periods Under Each Event
COVID-19 Black Friday Loyalty Program
Pre-adoption 2019-01-01 – 2019-12-31 2019-01-01 – 2019-10-31 2023-01-01 – 2023-05-31
Adoption
   Organic 2020-01-01 – 2020-02-29 2019-11-14 – 2019-11-28 2023-06-01 – 2023-07-31
   Event-Induced 2020-03-16 – 2020-04-30 2019-11-29 2023-08-01 – 2023-09-30
Post-adoption 2020-05-01 – 2023-04-30 2019-12-01 – 2020-10-31 2023-10-01 – 2024-03-31
We discuss each event and the definition of the three cohorts of consumers in detail below.
5.1.1 Event 1: COVID-19
Figure 1: Number of Offline-only Customers First Adopting Online Shopping due to COVID-19 (by month, Black Friday (by day), and Loyalty Program (by month)
(a) COVID-19
(b) Black Friday
(c) Loyalty Program
The COVID-19 pandemic was an exogenous macro-environmental shock that prompted many consumers to adopt online shopping. Figure 1(a) shows the number of customers who adopted online shopping through the retailer’s website and/or mobile app (from previously only shopping at physical stores) for each month for our observation period. As we can see, during March and April 2020, there was a large peak in the number of offline customers who first switched to online shopping; the number of adopters sharply increased from approx. 7,000 in January and February 2020 to approximately 20,000 in March and April 2020777In Figure 1(a), we also see smaller peaks associated with the second and third waves of COVID-19.. The customers who adopted online shopping during the onset of COVID-19 may have done so for reasons like health concerns, social distancing norms, etc., whereas the customers who adopted online shopping before that did so of their own volition, for other reasons (e.g., the convenience of online shopping), unrelated to COVID-19.
We consider the set of customers who made their first offline purchase before 2019-12-31 and were offline-only through the end of 2019 and continued shopping with the firm after the onset of COVID-19. Then, we categorize each consumer from this set into one of the following three groups (see second column in Table 5 for exact dates):
  • $\bullet$
    Covid adopters: Customers who made their first online purchase during the onset of COVID-19, i.e., from 2020-03-16 8882020-03-16 was the date when the Brazilian government started implementing physical distancing and confinement measures in response to COVID-19 (Faria de Moura Villela et al. 2021). to 2020-04-30.  [^9] The set of people who adopted during this period could also include consumers who would have organically adopted even without the COVID-19 shock. We have no clear way of distinguishing such potential organic adopters from those who adopted online shopping as a result of COVID-19. Nevertheless, based on Figure 1(a), it seems like approximately two-thirds of these consumers switched due to COVID-19. That is, a significant majority of these consumers switched due to COVID-19, and we refer to this cohort as covid adopters for simplicity. Please see $\S$8.1 for a detailed discussion on the implications of this potential mixing on the ATT estimates for covid and the other two events.
  • $\bullet$
    Organic adopters: Customers who adopted online shopping organically in the period just before the onset of COVID-19. They made their first online purchase between 2020-01-01 and 2020-02-29, i.e., the two months before the onset of COVID-19.
  • $\bullet$
    Offline-only customers: Customers who are active through both the pre- and post-adoption periods (i.e., at least one purchase before 2019-12-31 and after 2020-05-01), and never make online purchases during the estimation period.
In total, we have 33,246 covid adopters, 12,799 organic adopters, and 573,934 offline-only customers for this covid analysis. Note that a large fraction of consumers were able to remain offline-only during COVID-19 because the focal retailer’s stores were allowed to remain open, as pet supplies retailing was deemed an essential service by the Brazilian government.
5.1.2 Event 2: Black Friday Sales
Next, we discuss the retailer’s Black Friday sales event in 2019, characterized by a one-day, time-limited flash sales event offering discounts exclusively for online purchases. [^10] We focus on the Black Friday sales in 2019 and do not consider Black Friday sales in 2020 and afterwards because online adoption during those years may have been driven by COVID-19 or other marketing campaigns (e.g., the loyalty program), rather than purely driven by the sales event itself like 2019. Further, unlike 2019, the firm did not have major Black Friday sales in later years. Because of the exclusive online nature of the event, consumers interested in these deals were motivated to adopt online shopping. Figure 1(b) indicates a peak in the number of adopters on 2019-11-29, which is the date that Black Friday sales took place.
For this estimation sample, we consider the cohort of customers who made their first offline purchase before 2019-10-31 and continued to be the firm’s customers after the Black Friday sales date. Then, we categorize each consumer from this cohort into one of the following three groups (see Table third column in 5 for exact dates):
  • $\bullet$
    Black Friday adopters: Customers who made their first online purchase during the 2019 Black Friday sales.
  • $\bullet$
    Organic adopters: Customers who adopted online shopping organically in the two weeks preceding the 2019 Black Friday sales.
  • $\bullet$
    Offline-only customers: Customers who are active through both the pre- and post-adoption periods, and never make online purchases during the estimation period.
In total, we have 1380 Black Friday adopters, 3061 organic adopters, and 534,871 offline-only customers for the Black Friday analysis.
5.1.3 Event 3: Loyalty Program
Finally, we consider a retailer-led marketing campaign that incentivized offline-only customers to adopt the online channel. In August and September 2023111111Pilot studies for the loyalty program ran in two states (Rio Grande do Sul and Paraná) as early as May. Therefore, we exclude these two states from the estimation sample for this event., the retailer launched a loyalty program and extensively advertised it to customers through email and store marketing. Customers who wanted to join the LP had to use the mobile app (no offline options were available for activating the loyalty program). After completing the online activation, customers could access and activate monthly special offers and receive discounts on both online and offline purchases [^12] Note that customers had to activate each month for participation in each month’s special offers.. Figure 1(c) presents the number of adopters by month since January 2023 and shows clear peaks in the number of adopters during August and September 2023, coinciding with the launch of the loyalty program. This suggests that online-only activation of the newly available loyalty program prompted a significant number of customers to adopt online shopping during the roll-out period.
For this event, we consider the cohort of customers who made their first offline purchase before 2022-12-31 and continued to be the firm’s customers during the post-online-adoption period (i.e., after 2023-10-01). Then, we categorize each consumer from this cohort into one of the following groups (see last column of Table 5 for details):
  • $\bullet$
    Loyalty Program adopters: Customers who made their first online purchase in August or September 2023, with their first online purchase occurring after they activated the loyalty program online. In such cases, we can attribute the online loyalty activation as a reason for online adoption, as the activation precedes the first online purchase and both events occur within a similar time frame.
  • $\bullet$
    Organic adopters: Customers who made their first online purchase in June or July 2023, prior to the launch of loyalty program. [^13] Note that these consumers adopt online shopping organically, but may still activate loyalty offers after their online adoption. We will take this into account in our empirical analysis.
  • $\bullet$
    Offline-only customers: Customers who are active through both the pre- and post-adoption periods (i.e., at least one purchase before 2022-12-31 and after 2023-10-01), and never make online purchases during the estimation period.
In total, we have 1567 loyalty program adopters, 16,746 organic adopters, and 396,396 offline-only customers for this analysis.
5.2 Summary Statistics
For each event (Covid, Black Friday, and Loyalty Program), we generate customer-month-level panel data for the pre- and post-adoption periods. These three panel datasets will form the basis of our empirical analysis going forward. We present the detailed summary statistics on demographics (gender, age, household income, and locations), and pre-adoption purchase behaviors (spend, quantity, orders, brands, categories, etc.) for the three groups of customers (offline-only, organic adopters and event adopters) from each event (Covid, Black Friday and Loyalty Program) in Web Appendix $\S$§C. We discuss some key takeaways here.
Across our three samples for COVID-19, Black Friday, and the Loyalty Program, we find consistent differences between the two groups of adopters (organic adopters and event adopters) and offline-only customers. For instance, the adopter groups are consistently younger, higher income, and have higher spend with the retailer pre-adoption, relative to offline-only customers.
However, there are some important differences between the two groups of adopters (event adopters and organic adopters) across all three events. For instance, in the case of COVID-19, event/covid adopters tend to be older, have higher incomes, and have been shopping with the retailer for a longer time compared to organic adopters. However, these two groups have similar pre-adoption monthly spending. In the case of Black Friday, event(or Black Friday) adopters are younger, have been shopping with the retailer for a longer time, and have lower monthly spend with the retailer in the pre-adoption period relative to organic adopters. In contrast, in the Loyalty program case, event adopters tend to be slightly older, have been shopping with the retailer for a shorter period, and have a higher monthly spend with the retailer in the pre-adoption period compared to organic adopters. However, the differences in characteristics between the two groups of adopters for each event are much smaller in magnitude relative to the differences between adopters and non-adopters (i.e. offline-only customers).
6 Empirical Specification and Results
We now present our DID models in $\S$6.1 and the results in $\S$6.2.
6.1 Model
To estimate the impact of adopting online shopping on post-adoption purchase behavior, we use the difference-in-differences approach. Unlike our descriptive analyses, this approach allows us to account for unobserved heterogeneity across customers and time-specific shocks. For each event, we use customer-month-level panel data for the estimation sample corresponding to that event, and we utilize the customer-month observations from the pre- and post-adoption periods. We exclude the adoption periods from the estimation for the reasons discussed in $\S$5.1.
Specifically, we use the following DID specification to measure the ATT of online adoption on an outcome variable of interest, $y_{it}$, for customer $i$ in month $t$:
$y_{it}$ $=\alpha_{0}+\alpha_{1}\mathbbm{I}\{Adopter_{i}\}+\alpha_{2}\mathbbm{I}\{Post_{t}\}+\alpha_{3}\mathbbm{I}\{Adopter_{i}\}\times\mathbbm{I}\{Post_{t}\}+\boldsymbol{{\alpha}}_{4}^{\prime}\boldsymbol{{x}}_{it}+\epsilon_{it}.$ (1)
We have three main outcome variables of interest: spend, profitability and share of offline spend. We now discuss the interpretation of the coefficients and variables in the above equation for the case when spend is the outcome variable of interest. The interpretations are similar for the other outcome variables.
The intercept $\alpha_{0}$ captures the average monthly spend of offline-only customers in the pre-adoption period. $\mathbbm{I}\{Adopter_{i}\}$ is a binary variable that takes the value 1 for both organic adopters and event adopters, and 0 for offline-only customers. $\alpha_{1}$ denotes the pre-adoption period spend of adopters (including both organic adopters and event adopters), compared to offline-only customers. $\mathbbm{I}\{Post_{t}\}$ is a binary variable that takes the value 1 for the post-adoption periods, and 0 for pre-adoption periods. $\alpha_{2}$ captures the average change in spend of offline-only customers in the post-adoption period, relative to the pre-adoption period. $\alpha_{3}$ represents the average change in spend per month for adopters in the post-adoption period, respectively, relative to offline-only customers. $\alpha_{3}$ is the coefficient of interest here, representing the ATT estimate. We also include other time-varying control variables $\boldsymbol{{x}}_{it}$ to better isolate the causal effect of online adoption from other confounding factors that affect the outcome variables for one of our events – the loyalty program. Specifically, we construct control variables that capture customers’ participation in the loyalty program, such as whether they activate the offer or not [^14] We describe the list of control variables in Web Appendix $\S$§C.4..
We also use a second specification including customer fixed effects ($\gamma_{i}$) and time (month) fixed effects ($\eta_{t}$) as follows:
$y_{it}$ $=\gamma_{i}+\eta_{t}+\beta_{0}\mathbbm{I}\{Post_{t}\}\times\mathbbm{I}\{Adopter_{i}\}+\boldsymbol{{\beta}}_{1}^{\prime}\boldsymbol{{x}}_{it}+\epsilon_{it}$ (2)
Here, $\beta_{0}$ is the coefficient of interest, representing the ATT estimates.
Finally, we have a specification where we focus on the different types of adopters and use a model that compares organic adopters and event adopters, as follows:
$y_{it}=\delta_{0}+\delta_{1}\mathbbm{1}\{Event\_Adopter_{i}\}+\delta_{2}\mathbbm{1}\{Post_{t}\}+\delta_{3}\mathbbm{1}\{Event\_Adopter_{i}\}\times\mathbbm{1}\{Post_{t}\}+\boldsymbol{{\delta_{4}}}^{\prime}\boldsymbol{{x}}_{it}+\epsilon_{it}$ (3)
Here, $(\delta_{3})$, the coefficient of interest, represents the difference in the change in spend between event adopters and organic adopters in the post-adoption period relative to pre-adoption. We also include a second specification, Equation (4) with customer fixed effects ($\gamma_{i}$) and time (month) fixed effects ($\eta_{t}$).
$y_{it}=\gamma_{i}+\eta_{t}+\beta_{2}\mathbbm{1}\{Event\_Adopter_{i}\}\times\mathbbm{1}\{Post_{t}\}+\boldsymbol{{\beta_{3}}}^{\prime}\boldsymbol{{x}}_{it}+\epsilon_{it}.$ (4)
Here $(\beta_{2})$ is the coefficient of interest.
The validity of all the DID models discussed here relies on the parallel trends assumption. We discuss the validity of this assumption for all the models and outcome variables in $\S$8.2.
6.2 Results
We estimate the DID models specified above for two of our main outcome variables of interest: spend and profit, and present the results in this section. We discuss our third variable of interest, share of offline/online spend, in the next section. We also consider other outcome variables, e.g., quantity, orders, and number of unique brands purchased, in Web Appendix $\S$§E.2.
6.2.1 Effect on Total Spend
First, we present the results from the DID models specified in Equations (1) and (2) with offline-only customers as the control group and spend as the outcome variable in Table 6. Across all three events (COVID-19, Black Friday, and Loyalty Program), we find that adopters (pooling organic and event adopters) increase their spending in the post-adoption period (relative to pre-adoption) compared to offline-only customers.
Next, to examine the differences between each of the different event adopters and organic adopters, we estimate the DID models specified in Equations (3) and (4). The results from Equation (3) are in column (1), (3), and (5) of Table 7 for Covid, Black Friday, and Loyalty Program, respectively. We see that there is no significant difference in the post-adoption spend between covid adopters and organic adopters. However, the incremental spending of Black Friday adopters and Loyalty Program adopters is lower than organic adopters, post-adoption. We also control for customer and time fixed effects as shown in Equation (4) (in column (2), (4), and (6) of Table 7) and confirm the robustness of these findings to these additional controls.
In summary, our findings suggest that the different reasons for the adoption of online shopping do matter for the post-adoption incremental spending. While adopters of online shopping increased their spending with the retailer in the post-adoption period, relative to offline-only customers; those who adopt due to a macro-environmental shock (COVID-19, in our case) spend similarly post-adoption, relative to organic adopters, and those who adopt due to a promotional activity (Black Friday and Loyalty Program) by the firm, spend less.
Table 6: DID Main Analysis - Spend (Adopters vs. Offline-only)
Study: Covid Black Friday Loyalty Program
DV: Spend
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 280.1 (0.557) 291.7 (0.586) 332.6 (0.800)
[0.000] [0.000] [0.000]
Adopter 177.2 (2.885) 177.8 (9.177) 123.4 (4.42)
[0.000] [0.000] [0.000]
Post -62.90 (0.496) -96.95 (0.426) -51.36 (0.632)
[0.000] [0.000] [0.000]
Adopter * Post 140.4 (2.843) 140.1 (2.974) 81.29 (7.543) 78.37 (7.078) 10.02 (5.191) 84.89 (4.238)
[0.000] [0.000] [0.000] [0.000] [0.053] [0.000]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 22,204,608 22,204,608 9,505,999 9,505,999 4,403,449 4,403,449
R2 0.01905 0.38214 0.00980 0.41557 0.00918 0.48127
Within R2 0.00102 $6.84\times 10^{-5}$ 0.00294
Clustered (Customer) standard‐errors are presented in standard brackets, while p‐values are shown in square brackets.
Table 7: DID Main Analysis - Spend (Event Adopters vs. Organic Adopters)
Study: Covid Black Friday Loyalty Program
DV: Spend
Model: (1) (2) (3) (4) (5) (6)
Variables
Organic_Adopter 432.8 (5.325) 459.1 (10.48) 444.7 (4.415)
[0.000] [0.000] [0.000]
Covid_Adopter 32.84 (6.281)
[0.000]
Post 85.85 (5.34) 10.54 (9.334) -29.92 (5.041)
[0.000] [0.257] [0.000]
Covid_Adopter * Post -10.64 (6.271) -3.095 (6.465)
[0.089] [0.632]
BlackFriday_Adopter 32.01 (20.72)
[0.122]
BlackFriday_Adopter * Post -82.25 (15.47) -69.88 (14.26)
[0.000] [0.000]
Loyalty_Adopter 131.7 (19.08)
[0.000]
Loyalty_Adopter * Post -254.7 (20.82) -53.22 (19.42)
[0.000] [0.006]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 76,842 76,842 197,658 197,658
R2 0.00131 0.39193 0.00081 0.41147 0.05635 0.47126
Within R2 $5.37\times 10^{-7}$ 0.00068 0.02692
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
6.2.2 Effect on Profitability
In this section, we examine whether and how the adoption of online shopping affects the profitability of adopters relative to offline-only customers, and whether there are differences between the different groups of event adopters relative to organic adopters.
In our data, we observe item-level profit margins for each purchase [^15] We observe this starting January 2020. We obtain the 2019 profit margins of each item in a given channel by extrapolating the item-channel level profit margins from 2020–2023 to the 2019 time-frame based on Brazil’s Extended National Consumer Price Index (referred to as IPCA). We refer interested readers to Web Appendix $\S$§D for details. From our discussions with managers at the focal retailer, we do not have any reason to believe that the profit margin structures changed between 2019 and the rest of the data period.. We first provide some summary statistics of the differences in prices and profit margins across the two channels. To highlight the differences in these three variables across channels, we first generate a panel of prices at the item-channel-month level [^16] For instance, in the case of price, an observation in this panel is the average price of an item $j$ sold through a channel $c$ in time period $t$, and is calculated by the total sales of item $j$ in channel $c$ in month $t$ divided by the total quantity of that item sold in that channel-month.. Next, we calculate the offline-online differences at the item-month level to obtain item-specific offline-online differences in price and profit margin. We then average over months and obtain item-level average differences in price and profit margin between offline and online channel. Table 8 provides item-level summary statistics for the 22,803 items that were sold in both offline and online channels in our data (and were purchased at least once in the same month). For confidentiality, we multiply the profit margin by the same undisclosed multiplier used for the price. We see from Table 8, that on average, prices in physical stores are higher than online prices, and profit margins are higher in physical stores vs. online. Thus, consumers who spend more in physical stores may be more profitable (assuming they buy similar assortments of products).
Table 8: Item-Level Offline Price and Profit Margin Relative to Online
mean std 25% 50% 75% (min, max) count
Offline - Online Price 6.01 91.07 -0.84 4.61 17.71 (-7,598.46,1,430.91) 22,803
Offline - Online Profit Margin 5.37 78.49 -1.77 3.47 14.05 (-6,586.62, 2,163.31) 22,803
Table 9: DID Main Analysis - Profit (Adopters vs. Offline-only)
Study: Covid Black Friday Loyalty Program
DV: Profit
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 92.57 (0.177) 96.13 (0.187) 125.8 (0.288)
[0.000] [0.000] [0.000]
Adopter 61.35 (0.925) 60.94 (3.041) 41.57 (1.607)
[0.000] [0.000] [0.000]
Post -10.56 (0.174) -22.83 (0.152) -16.81 (0.238)
[0.000] [0.000] [0.000]
Adopter * Post 35.87 (0.974) 37.37 (1.070) 19.61 (2.547) 18.44 (2.506) -7.231 (1.820) 18.41 (1.506)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 22,204,608 22,204,608 9,505,999 9,505,999 4,403,449 4,403,449
R2 0.01228 0.31903 0.00450 0.34750 0.00651 0.44546
Within R2 0.00047 $2.48\times 10^{-5}$ 0.00193
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Next, we estimate the DID models in Equations (1)–(2) with profit as the dependent variable, and present the results in Table 9. We see that across all the events, adopters show an increase in profitability post-adoption, relative to offline-only customers.
Table 10: DID Main Analysis - Profit (Event Adopters vs. Organic Adopters)
Study: Covid Black Friday Loyalty Program
DV: Profit
Model: (1) (2) (3) (4) (5) (6)
Variables
Organic_Adopter 145.7 (1.698) 154.7 (3.519) 164.5 (1.627)
[0.000] [0.000] [0.000]
Covid_Adopter 10.96 (2.007)
[0.000]
Post 20.27 (1.788) 3.315 (3.163) -21.05 (1.788)
[0.000] [0.295] [0.000]
Covid_Adopter * Post 7.248 (2.117) 9.079 (2.274)
[0.000] [0.000]
BlackFriday_Adopter 7.338 (6.79)
[0.28]
BlackFriday_Adopter * Post -20.52 (5.228) -16.04 (5.094)
[0.000] [0.002]
Loyalty_Adopter 34.13 (6.413)
[0.000]
Loyalty_Adopter * Post -86.83 (7.352) -15.41 (6.835)
[0.000] [0.024]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 76,842 76,842 197,658 197,658
R2 0.00134 0.29792 0.00039 0.35806 0.04568 0.44335
Within R2 $2.79\times 10^{-5}$ 0.00025 0.02248
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Finally, we investigate the differences between each group of event adopters relative to organic adopters by running the DID models specified in Equation (3) and Equation (4). The results are shown in Table 10. First, we see that covid adopters are significantly more profitable than organic adopters post-adoption. The differences are economically meaningful and large in magnitude – the post-adoption gain in profitability is over 30% (7.248/20.27) higher for covid adopters relative to organic adopters. The fact that covid adopters are more profitable than organic adopters despite there being no significant differences in spend is an interesting finding. This may be due to the differences in prices across channels in our setting and potential differences in the share of offline/online spending. We explore this mechanism more in the following section. Also interestingly, we find that Black Friday adopters and Loyalty Program adopters are less profitable relative to organic adopters in the post-adoption period by 10.2% (16.04 / (154.7 + 3.3)) and 10.7% (15.41 / (164.5 - 21.05)), respectively, and this is likely driven by their lower spending post-adoption.
To summarize, in our empirical setting, we find that adopters due to the macro environmental shock of COVID-19 are more profitable post-adoption, relative to organic adopters despite similar spending. However, those that adopted due to firm driven promotions like Black Friday and the Loyalty Program are less profitable. An important managerial implication is that the firm should not naively assume that users who adopt online shopping due to external factors will show the same levels of profitability post-adoption relative to organic adopters. We discuss the mechanisms and theoretical underpinnings behind our findings in the next section, and we also expand on the managerial implications.
7 Mechanisms, Theoretical Underpinnings and Managerial Implications
We now discuss the mechanisms and theoretical underpinnings behind our findings on spend and profit; if and how managers can translate these findings to other settings, and some general takeaways from our analysis.
7.1 Share of offline/online spend and the theory of Consumer Inertia
One of our interesting findings from $\S$6.2 is that covid adopters are more profitable compared to organic adopters even though their overall spending is similar. We now examine if this is due to differences in the post-adoption share of offline/online spending between different groups of adopters in more detail. We also discuss the theoretical underpinnings behind these differences and derive implications for managers based on theory and our empirical findings.
Consumer inertia and habit theory (Verplanken and Wood 2006; Wood and Neal 2009)would predict that adoption driven by external shocks (e.g., COVID-19, firm driven promotions) may lead to temporary disruptions in consumer habits (adopting a new channel and multichannel shopping), but without reinforcement, consumers may revert to prior routines (not use the newly adopted channels much). We test this prediction in our empirical setting, by comparing post-adoption channel usage between consumers who adopted online shopping organically and those who did so in response to the COVID-19 shock, a short-term one-time firm-driven promotion (Black Friday) and a more long-term and recurring firm driven promotion (Loyalty Program). In other words, we examine if there are differences in how the total spend with the retailer is distributed across the offline and online channels for the different groups of event adopters relative to organic adopters in the post-adoption period.
As a first step, we run the DID models specified in Equation (3) and (4) with the share of offline spend, and offline spend as the outcome variables. The results are presented in Web Appendix $\S$§E.1. We find that covid adopters spend a larger fraction of their spend offline relative to organic adopters (by about 7 percentage points) and also have a higher offline spend in the post-adoption period  [^17] Since covid adopters are more active in the offline channel, we similarly see a corresponding positive effect on the number of orders and variety of products purchased offline; see Table A26 in Web Appendix $\S$§E). Black Friday adopters also spend a larger fraction of their spending offline relative to organic adopters (by about 9 percentage points). Loyalty Program adopters also have a higher share of offline spending relative to organic adopters.
The results above represent the averages for the entire post-adoption period and mask any trends. So we also investigate whether there are differential trends in the share of offline spend between each type of event adopter relative to organic adopters in the post-adoption period. In Figure 2, we show the trend lines for share of offline spend for each type of event adopter vs. organic adopters (and the differences).
Figure 2: Share of offline spend relative to total spend – Event Adopters vs. Organic Adopters
(a) Covid
(b) Black Friday
(c) Loyalty Program
From Figure 2(a), in the case of COVID-19, we see that covid adopters not only have a higher offline share of spending, post-adoption, relative to organic adopters, but their trend of moving towards more online shopping is much slower than that of organic adopters181818see Web Appendix $\S$§E.1 for formal tests of the magnitude of the difference. This is fully in-line with what consumer inertia and habit theory would predict. Further, we know from $\S$5.2 that covid adopters are older and thus are more likely to be more set in their habits and exhibit higher inertia (Yang and Ching 2014). To test whether the difference in the rate of growth of online shopping is primarily driven by differences in age, we perform a median split of all the covid adopters and organic adopters and confirm that the differences in trends only show up among older adopters and there is no difference for younger adopters [^19] See Figure A3 and Table A24 in Web Appendix E.1 for more details.
In the case of Black Friday promotion, we see from Figure 2(b) that the level of offline share of spending post-adoption is higher relative to organic adopters, again consistent with what one would expect from consumer inertia and habit theory. However, Black Friday adopters show a faster rate 202020with p-value 0.109; see Web Appendix E.1 Table A23. Extending the post-adoption period further reduces p-value. of increasing their online share of spend relative to organic adopters. This may be explained by the fact that Black Friday adopters are younger than organic adopters, and thus are less set in their habits and more open and willing to explore the conveniences of online shopping. However, since they adopted online shopping due to an external incentive, it still takes some time to catch up organic adopters in terms of post-adoption share of online spending.
Finally, in the case of the Loyalty Program, from Figure 2(c) we see again, that Loyalty Program adopters have a slightly higher share of offline spend, and there is no difference in trends relative to organic adopters. The level difference again is what we might expect from consumer inertia theory, and the fact that there is no difference in trends may be due to the fact that the two adopter groups are quite similar in terms of demographics, and also the fact that the Loyalty Program involves monthly online activation, and therefore there is some reinforcement for the habit of online shopping for Loyalty Program adopters.
Thus, based on consumer inertia theory and our empirical findings, managers should not expect consumers who adopt online shopping for external reasons to have as high a share of online spending after adoption as organic adopters. Moreover, whether the rate of growth in online share of spend for consumers who adopt online shopping due to external reasons is higher or lower relative to organic adopters depends on whether the external event/incentive is more likely to attract younger or older consumers (relative to organic adopters) to adopt online shopping [^21] For interested readers, we also explore the online-offline differences in other metrics of interest (e.g. Quantity purchased, Number of orders, Unique items, Brands, Categories, and Subcategories) in Web Appendix E.2..
7.2 Lower post-adoption spend from firm-driven promotions and the theory of forward buying
Our other interesting finding from $\S$6.2 is both groups of firm driven promotion adopters i.e. Black Friday adopters and Loyalty Program adopters exhibit lower post-adoption spend and profit relative to organic adopters. We know that promotions are an important mechanism for firms to induce various kinds of consumer behavior including the adoption of online shopping. A plausible theoretical explanation for this finding is the theory of forward buying and demand distortion (Neslin, Henderson, and Quelch 1985; Hendel and Nevo 2006; Van Heerde and Neslin 2017). This theory indicates that that time-limited promotions may induce forward buying, leading to demand distortion and reduced post-promotion consumption .
In our context, Black Friday adopters and Loyalty Program adopters could be increasing their spending during the promotion (i.e., the adoption period) and effectively pulling forward their future demand, resulting in a “post-promotion dip” in spending in the post-adoption period. To examine if this is the case, we first present a visual depiction of the trendlines for monthly spending of Black Friday adopters and Loyalty Program adopters relative to organic adopters and offline-only customers in Figure 3. We see that there is a spike in spend during the month of adoption, for all three types of adopters  [^22] We see a similar pattern for covid adopters as well. and the increase in spending during the month of adoption is higher for Black Friday adopters and Loyalty Program adopters relative to organic adopters.
To formally examine the differences between the two groups of event adopters relative to organic adopters during each month around adoption, we estimate the following event study regression:
$y_{it}$ $=\gamma_{i}+\eta_{t}+\sum_{j}\theta_{j}(PRE_{it}(j)\cdot\mathbbm{I}\{Event\_Adopter_{i}\})$ (5)
$+\sum_{k}\theta_{k}(POST_{it}(k)\cdot\mathbbm{I}\{Event\_Adopter_{i}\})+\boldsymbol{\lambda}^{\prime}\boldsymbol{x}_{it}+\epsilon_{it}$
In this model, $PRE_{it}(j)$ is an indicator that equals 1 if period $t$ is $j$ months prior to the month of adoption. The coefficient $\theta_{j}$ for $j=-J,-(J-1),\ldots,-1$ measures the difference in pre-adoption trends between a type of event adopter and organic adopters. These coefficients allow us to test for the validity of the parallel trends assumption, which is discussed in detail in §8.2. Next, $POST_{it}(k)$ is an indicator that equals 1 if period $t$ is $k$ months after online adoption. The coefficient $\theta_{k}$ captures the differences in spend between event adopters and organic adopters during each month following adoption. This event study regression allows us to measure: (1) the differences spending between event adopters and organic adopters during the month of adoption, $\theta_{0}$; and (2) the dynamic post-adoption effects $\theta_{k}$ for $k=1,2,\ldots$, which reflect the spend differences in the subsequent months post-adoption. We compare each group’s spending in a time period to their most recent pre-adoption outcome, i.e., spend in the month immediately prior to adoption.
We estimate the event study model in Equation (5) for Black Friday and the Loyalty Program and present the estimation results in Web Appendix §E.3. Figure 4(a) shows the point estimates and 95% confidence intervals for the dynamic treatment effects. We see a clear pattern consistent with the notion of forward buying among both Black Friday adopters and Loyalty Program adopters, relative to organic adopters. During the month of adoption, they spend more than organic adopters, and in the months following adoption, they spend less [^23] Notably, if we only look at offline spending (see Figure 4(a) in Web Appendix E.3), there are no significant differences in offline spending between Black Friday adopters and organic adopters, suggesting that the forward buying effect is specific to the online channel for Black Friday. But for the Loyalty Program, we see the effect across both channels.. Finally, we confirm 242424see Figure A5 in Web Appendix E.3. that for COVID-19, which is not a firm-driven promotion, we do not see evidence of forward buying.
Figure 3: Monthly Total Spend Visualization – Black Friday/Loyalty Program vs. Organic Adopters.
(a) Black Friday
(b) Loyalty Program
Figure 4: Event Study for Forward Buying: Both Black Friday adopters and Loyalty Program adopters spend more than organic adopters during the month of adoption. However, they spend less than organic adopters in the following months
(a) Black Friday
(b) Loyalty Program
Thus, based on the theory of forward buying during promotions and our empirical findings, managers should not expect as high a post-adoption spend for promotion-driven adopters of online shopping relative to organic adopters.
7.3 Managerial Implications
Our main findings indicate that consumers who adopt online shopping due to different reasons (organic, macro environmental shocks, or firm-driven promotional activities) behave differently in the post-adoption period. We have discussed above that our findings are explained and supported by the theories of consumer inertia and forward buying, and we believe that our main takeaways can be abstracted to other similar situations where these same theoretical concepts may apply. An important and practically useful managerial implication is that firms should not naively use the post-adoption behavior of organic adopters as a proxy to forecast the behavior of consumers who adopt online shopping for other reasons such as environmental or macroeconomic shocks, or firm-run marketing campaigns, and we highlight this implication for managers evaluating promotional campaigns to increase adoption of online shopping with some numbers from our empirical context.
To assess whether to run a promotional campaign (like the Black Friday online discounts) to induce customers to adopt online shopping, firms have to forecast the expected benefits in terms of incremental sales from consumers who choose to adopt due to the promotion, and compare the benefits to the costs of running the campaign. If managers naively use organic adopters’ incremental profit as a proxy for adopters resulting from the promotional campaign (Black Friday adopters, in this case), they could grossly overestimate the expected benefits. This in turn would lead them to wrongly compute metrics like the time taken to break-even from such a promotional campaign. To illustrate this point, we use the Black Friday promotional campaign as an example and show some back-of-the-envelope calculations.
First, we quantify the cost of running the Black Friday promotion, by computing the difference between actual consumer spending under promotional prices on Black Friday and what consumers would have paid for the same quantities at regular prices. We find the average discount across categories is around 28%, with dry food categories discounted around 20% and hygiene categories around 33%. The average promotional cost of Black Friday sales is 299.18 MCU per customer for that day. Then, from our DID results, we obtain the average monthly incremental profit of organic adopters (23.57 MCU) and Black Friday adopters (7.81 MCU)  [^25] Numbers are taken from Column 3 of Table A34 in Web Appendix $\S$§E.4.. By using the behavior of organic adopters to predict the behavior of Black Friday adopters, managers would predict that the incremental profits from Black Friday adopters would result in a break-even duration of 13 months post-adoption (i.e. 299.18 / 23.57). However, our analysis shows that it should actually take at least 39 months (e.g., 299.18 / 7.81) to break even, nearly three times longer than the managers’ initial forecast. Therefore, managers should appropriately calibrate their expectations of increased consumer spending when running promotions that encourage the adoption of online shopping.
8 Robustness Checks
We now present a series of robustness checks to provide additional credibility for our findings.
8.1 Effect size measurement
A potential concern with our ATT measurement, especially for our Covid analysis, is that covid adopters may be “contaminated” with some organic adopters – customers who would have adopted the online channel organically during the onset of COVID-19, had COVID-19 not occurred. As a result, the treatment group may include a mix of “true” covid adopters and organic adopters. If such contamination exists, the ATT estimates reported in §6.2 would be a lower bound on the true ATT for covid-driven adoption (relative to organic), and we can back out the ATT for the subset of customers of interest (i.e., the “true” covid adopters). The true ATT for covid adopters relative to organic adopters can be obtained by dividing the estimated ATT (reported in §6.2) by the fraction of “true” covid adopters, given by Equation 6:
$\textrm{True}~ATT_{covid,organic}=\frac{\textrm{Estimated}~ATT_{covid,organic}}{w_{covid}}$ (6)
where $w_{covid}$ is the fraction of “true” covid adopters in the sample. [^26] Assuming the monthly incidence of organic adopters remains constant, we estimate that around 38.5% of the covid adopters could actually be organic adopters (12799/33246)(12799/33246)( 12799 / 33246 ). Hence, $w_{covid}=61.5\%$ can be the fraction of “true” covid adopters. Therefore, the “true” ATT of spend for covid-driven adopters relative to organic adopters would be $\frac{-3.095}{0.615}=-5.033$ (std. error: 10.512), and the “true” ATT of profit would be $\frac{-9.079}{0.615}=14.763$ (std. error: 3.698) [^27] Numbers are taken from the column 2 of Tables 7 and 10.. This confirms that even with potential contamination, the difference in spend relative to organic adopters is small and with a confidence interval that includes zero, while the difference in profit is larger and with a confidence interval that does not include zero. As such, our substantive conclusions remain unchanged.
Note that we are less concerned about the contamination problem for our other two events for a few reasons, which we briefly discuss below.
  • $\bullet$
    First, since the Black Friday sale occurs only on a single day (and not over a prolonged period, unlike COVID-19), and there is a large spike in adopters on that single day compared to the days around it, the extent of potential contamination is very small. In this case, the results are very similar to those from $\S$6.2 since historical data suggest that close to 85.7% of adopters during the Black Friday event adopt due to the event (and not organically). Nevertheless, we still conduct a similar analysis by backing out the “true” ATT using the estimated ATT of Black Friday adopters relative to organic adopters divided by $w_{black\_friday}=85.7\%$. The “true” ATT of spend for Black Friday adopters relative to organic adopters would be $\frac{-69.88}{0.857}=-81.54$ (std. error: 16.639), and the ATT of profit would be $\frac{-16.04}{0.857}=-18.716$ (std. error: 5.944). [^28] Numbers are taken from the column 4 of Tables 7 and 10. This confirms that even with such potential contamination, Black Friday adopters have lower spend and profit relative to organic adopters.
  • $\bullet$
    Next, recall that the Loyalty Program adopters are customers who specifically went online for the first time to activate the loyalty program on their mobile app and only made their first online purchase after activation, so contamination with organic adopters (those adopted online shopping before the loyalty program was launched and then activated the loyalty program after it was launched) is unlikely.
8.2 Parallel pre-trends
We test whether the parallel trends assumption holds for our DID analyses comparing spend and profitability [^29] Note that for our analyses of share of offline spend, parallel trends are satisfied by definition since all customers have 100% offline spend pre-adoption. We also present the results for our other minor metrics. between event adopters and organic adopters, and between adopters and offline-only customers, using both visual inspection and more formal event study regressions in Web Appendix $\S$§F.1. The parallel trends assumption appears to be satisfied for both spend and profit when comparing event adopters and organic adopters for all three events. However, the parallel pre-trends assumption is not satisfied for the comparisons between adopters and offline-only customers, across all three events. Further, specifically for the case of covid adopters and organic adopters, there seems to be a difference in pre-trends visually, although the differences in pre-adoption outcomes are not statistically different for most of the pre-adoption months. For these cases (spend and profits for adopters vs. offline-only customers for all three events, and covid adopters vs. organic adopters), we bound our treatment effects under different assumptions of the extent of the violation using the Honest Difference-in-Differences approach proposed by Rambachan and Roth (2023) in Web Appendix $\S$§F.2. In all instances but one, we conclude that even if the post-adoption trend violation is twice as large as the pre-adoption trend, our substantive conclusions remain unchanged in terms of both direction and statistical significance. The only exception is that for the Loyalty Program event, the difference in post-adoption profit for adopters relative to offline-only customers has a confidence interval that includes zero. However, even for this event, the difference in post-adoption spend is large and with a confidence interval that does not include zero.
8.3 Doubly robust DID, Propensity Score Matching and IPW approaches
In $\S$5.2, we saw that there are some differences in the demographics and other pre-adoption variables between organic adopters and event adopters. These differences should not affect the validity of our results as long as the parallel trends assumption holds (Angrist and Pischke 2008; Roth et al. 2023). Nevertheless, to provide additional credibility to our findings and ensure that they are robust to any factors correlated with observable characteristics that may affect the propensity to adopt online shopping organically versus due to external events, we also consider doubly robust DID (Sant’Anna and Zhao 2020) and propensity score-based approaches (Rosenbaum and Rubin 1985; Stuart et al. 2011). These methods provide better comparability between groups by improving covariate balance by discarding poorly matched customers. We compute propensity scores using a logistic regression model, with variables including demographics (e.g., gender, age, and household income) and pre-adoption purchase variables (i.e., average monthly spend, number of unique orders, unique brands, and unique categories). With the estimated propensity scores, we use our DID estimator using both IPW and PSM methods, and also run a doubly robust DID estimator. The propensity score model specification and detailed results from the IPW, PSM and doubly-robust DID methods are shown in Web Appendix $\S$§F.3. We find that our substantive conclusions remain the same with these approaches.
8.4 Alternate definitions for the organic adopter and covid adopter cohorts
To check whether our results are sensitive to the particular months we chose to define our covid adopter and organic adopter cohorts in our COVID-19 analysis, and to test whether our conclusions would change if we used “later” covid adopters vs. “early” covid adopters, we repeat our analysis with alternative definitions for both cohorts and present results in Web Appendix $\S$§F.4. To start with, we redefine organic adopters as those customers who were offline-only before 2019-10-31 and made their first online purchase between 2019-11-01 and 2020-02-29 (i.e., we extend the period we previously used to define organic adopters). We re-run the same set of DID models and find that all results are consistent with our main analysis. Next, we redefine the cohort of covid adopters as those customers who were offline-only before 2019-12-31 and made their first online purchase between 2020-05-01 and 2020-06-30 (i.e., two months later than we had used in our main analysis). Again, we find that all the results are consistent with the main analysis. Finally, we compare “early” covid adopters (who adopted online shopping between 2020-03-16 and 2020-04-30) and “later” covid adopters (who adopted online shopping between 2020-05-01 and 2020-06-30). Again, we run the same set of DID models and do not find any significant differences except that the “later” covid adopters spend slightly higher by 2.48%.303030p-value = 0.048.
These analyses also addresses another potential concern with our main analysis, which is that organic adopters adopt for two extra months relative to covid adopters before our common post-adoption period, and this may drive some of the results. The fact that our conclusions don’t change when we change the set of months used to define our cohorts also allays this concern. Finally, our definitions for Black Friday adopters and Loyalty Program adopters are much more precise than for covid adopters, so we don’t repeat this robustness check for those analyses.
8.5 Using Log Scale for Dependent Variables
To test whether the positive skewness in purchase outcomes affects our results, we re-estimate all DID models using a natural log transformation for the dependent variables ($\ln(y+1)$) and report the results in Web Appendix $\S$§F.5. Across all specifications, the results remain consistent with our main findings, confirming the robustness of our analysis to log transformation of the outcomes.
8.6 Handling Zero-Inflated Dependent Variables
A potential concern in our setting is that customers do not make purchases every month, and many customer-month spending observations are zero. Therefore, we consider an alternative specification that accommodates zero-inflated dependent variable. Specifically, we estimate a two-part (hurdle) model at the customer-month level (Cameron and Trivedi 2005). This approach involves: (1) a linear probability model with a DID specification to estimate changes in the likelihood of making a purchase during the post-adoption periods [^31] Here, the dependent variable of linear probability model is $\mathbbm{I}\{Spend_{it}\}$, a binary indicator for whether customer $i$ made purchase during month $t$, and the DID specification is Equations (1) and (2) for comparing adopters vs. offline-only customers, and Equations (3) and (4) for comparing organic adopters vs. event adopters); (2) a DID model that estimates the changes in spend conditional on making a purchase. [^32] We estimate the same equations as the main analysis, but use the customer-month observations with non-zero spend. Overall, we find that the results from these specifications are consistent with our main results; see Web Appendix $\S$§F.6 for detailed estimation results.
9 Conclusions
In conclusion, this paper examines the impact of the adoption of online shopping, i.e., shopping on a retailer’s e-commerce website and/or mobile app, on customers who previously only shopped at physical stores, using data from a large pet supplies retailer in Brazil spanning the period of 2019-2024. We study four groups of consumers who adopted online shopping for different reasons during our data period: (a) covid adopters who started shopping online during the onset of COVID-19, (b) Black Friday adopters who adopted online shopping due to deep discounts available online during Black Friday, (c) Loyalty Program adopters who adopted online shopping during the launch of a loyalty program which had to be activated on the retailer’s mobile app, and (d) organic adopters who started shopping online fully of their own volition, without any external stimuli. We find that for all four types of adopters of online shopping, there is an increase in spend with the retailer post-adoption, relative to consumers who stayed offline-only.
However, there are important differences between the four groups of adopters in terms of their post-adoption behavior along three dimensions - spend, profitability and share of offline/online spend. Compared to organic adopters, covid adopters, who adopted online shopping due to a macro-environmental shock, spend similarly in the post-adoption period. However, Black Friday adopters and Loyalty Program adopters, who adopted online shopping due to firm-driven promotional activities, spend less. We show that this is explained by the theory of forward buying during promotions - Black Friday adopters and Loyalty Program adopters increase their spending during the promotion, (when they adopt online shopping) by pulling forward their future demand, resulting in a “post-promotion dip” in their demand. In terms of channel usage post-adoption, consistent with consumer inertia and habit theory, we find that all three groups of externally driven adopters have a lower share of online spend post-adoption relative to organic adopters. Further, we find that covid adopters also have a slower growth in the share of online spending relative to organic adopters. This is mainly driven by the fact that covid adopters are older on average, compared to organic adopters, and thus are more set in their habits. In contrast, the younger Black Friday adopters show a faster rate of growth for online spend share relative to organic adopters. Loyalty Program adopters do not differ significantly from organic adopters in terms of the growth in the share of online spending, probably because their demographics are quite similar and the fact that consumers have to activate the loyalty program online every month. Finally, we find that covid adopters are more profitable than organic adopters due to their higher share of offline spending, where margins are higher in our setting. Black Friday adopters and Loyalty Program adopters are less profitable relative to organic adopters, due to their lower spending.
Taken together, our findings indicate that consumers who adopt online shopping due to different reasons (organic, macro environmental shocks or different types of firm driven promotional activities) behave differently post-adoption, and the differences can be explained using consumer inertia and habit theory, the theory of forward buying during promotions, and the differences in demographics between these groups of consumers. Managers can use their knowledge of consumer demographics, and the theories of consumer inertia and promotional forward buying to predict spend, profit and channel usage outcomes in other situations where these same theories apply.
An important managerial implication is that firms should not naively use the behavior of organic adopters of a new channel as a proxy to forecast the behavior of consumers who adopt the new channel for other reasons such as environmental or macroeconomic shocks, or firm-run marketing campaigns. As an example, we find that the retailer only breaks-even on running the Black Friday promotion to induce more customers to adopt online shopping, after 39 months, instead of the expected 13 months it would have predicted based on the post-adoption behavior of organic adopters. Further, even if post-adoption spending is similar (like in the case of covid adopters), differences in the usage of online channels vs. offline could cause significant differences in profitability.
Finally, we note that our study has limitations that serve as avenues for future research. First, as highlighted above, the demographics of the different types of adopters are an important input along with the theories of consumer inertia and forward buying, to predict post-adoption behavior. The external stimuli for online adoption in our setting are limited to the COVID-19 pandemic, and two specific promotional campaigns run by our retailer. Future research could investigate whether other types of events (e.g. store closures or other types of promotions or macro-shocks) lead to adopters with different demographics (and therefore potentially different interactions with these behavioral forces) than what we see in our setting. Second, we focus on one retail category (pet supplies), and it is unclear if consumers in other categories and contexts, with other forces at play, show similar behavior. For instance, recent research in the non-retail and more competitive hotel industry has found that adopters of a hotel chain’s mobile app decrease their spending at the focal hotel because they tend to also adopt competitors’ apps at the same time (Gu and Kannan 2021). Thus, future research that studies these questions in a wider set of categories and contexts can further enhance our understanding of the consequences of consumer migration to online channels.
Funding and Competing Interests
The authors were not paid by the data provider. A preliminary version of this manuscript was reviewed by the data provider to ensure that data was accurately depicted, but they did not review/comment on the content of the manuscript.
Appendix A Data cleaning
Table A1: Item categories with substantially different availability time via online and offline channels (YY-MM-DD)
Item Category First Offline Purchase Date First Online Purchase Date
Large size food 2020-09-12 2020-02-28
Human food 2019-01-02 2021-02-12
Large size pharmacy 2021-02-03 2020-12-09
Pet Cleaning Supplies 2019-01-03 2022-03-13
There are four product categories with significant discrepancies in their initial availability time on the offline and online channels. For example, the product category Large size food (e.g. pack size greater than 10KG) was available online in February 2020 but only became available offline in September 2020. In addition, the product category Pet Cleaning Supplies was only available offline before COVID-19 but became available online in August 2022.
Including these food categories in our analysis can be problematic since it also reflects retailers’ assortment and the channel-specific effects (i.e. consumers can only buy some products exclusively from one channel). Therefore, we exclude these four categories from our analysis. The fraction of sales of these four categories only accounts for 0.63% of the total revenue during the sample period, and as such unlikely to have any meaningful impact on our findings.
Appendix B Appendix to Descriptive Analysis
In this section, we provide some descriptive analyses summarizing the relationship between channel choice and consumers’ purchase behaviors. In particular, we are interested in quantifying whether the value/revenue generated by a customer is systematically correlated with their channel choice. In $\S$§B.1, we first present some cross-sectional evidence linking channel choice to customer value. Next, in $\S$§B.2, we present a within-user analysis that can control for consumer-specific variables.
B.1 Cross-sectional Comparison
Table A2: Frequency distribution of gender – offline-only/online-only/multichannel
Gender Online-only (%) Offline-only (%) Multichannel (%) Overall (%)
No Information 41.95 37.24 24.50 34.75
Female 37.75 35.23 49.08 39.62
Male 20.30 27.53 26.42 25.63
Total 100 100 100 100
Figure A1: Cumulative distribution function (CDF) of Age/Monthly Household Income – offline-only, online-only,multichannel
We start by categorizing consumers into three mutually exclusive groups:
  • $\bullet$
    Multichannel consumers – consumers who have purchased from both offline and online channels at least once during the observation period.
  • $\bullet$
    Offline-only consumers – consumers who have exclusively purchased from physical stores during the observation period.
  • $\bullet$
    Online-only consumers – consumers who have only purchased using the online channels (website and/or mobile app) during the observation period.
In total, our data consists of 1,642,859 (27.7%) multichannel consumers, 2,986,541 (50.3%) offline-only consumers, and 1,304,848 (22.0%) online-only consumers.
Table A3: Customer-Month Summary Statistics By Customer Types – Multichannel/Offline-only/Online-only
Variables Mean Std Min 25% 50% 75% Max Count
Multichannel Customers
Spend 398.16 805.24 0 0 0 548.07 301,134.12 56,151,374
Quantity 3.15 9.69 0 0 0 3.00 4,301.00 56,151,374
Order 0.76 1.18 0 0 0 1.00 236.00 56,151,374
UniqueItem 1.74 3.20 0 0 0 2.00 355.00 56,151,374
UniqueBrand 1.39 2.32 0 0 0 2.00 127.00 56,151,374
UniqueSubcategory 1.20 1.87 0 0 0 2.00 43.00 56,151,374
UniqueCategory 1.01 1.46 0 0 0 2.00 16.00 56,151,374
Offline-only Customers
Spend 221.46 738.65 0 0 0 226.34 2,340,617.16 56,259,534
Quantity 1.96 11.38 0 0 0 2.00 44,158.00 56,259,534
Order 0.50 0.91 0 0 0 1.00 430.00 56,259,534
UniqueItem 1.25 2.82 0 0 0 2.00 4,191.00 56,259,534
UniqueBrand 1.01 1.96 0 0 0 1.00 332.00 56,259,534
UniqueSubcategory 0.89 1.62 0 0 0 1.00 64.00 56,259,534
UniqueCategory 0.76 1.29 0 0 0 1.00 18.00 56,259,534
Online-only Customers
Spend 236.49 539.39 0 0 0 311.22 110,682.95 11,337,544
Quantity 1.33 6.21 0 0 0 1.00 4,800.00 11,337,544
Order 0.42 0.68 0 0 0 1.00 130.00 11,337,544
UniqueItem 0.72 1.49 0 0 0 1.00 141.00 11,337,544
UniqueBrand 0.62 1.15 0 0 0 1.00 68.00 11,337,544
UniqueSubcategory 0.56 0.98 0 0 0 1.00 25.00 11,337,544
UniqueCategory 0.51 0.84 0 0 0 1.00 11.00 11,337,544
We find that these three groups are systematically different from each other in terms of demographics (see Table A2 and Figure A1). Multichannel consumers are more likely to be female compared to offline and online consumers. Online-only consumers tend to be younger than both offline and multichannel consumers. Further, offline-only consumers tend to be lower income while multichannel consumers tend to be the wealthiest. There are also significant differences in channel choice across geographic locations (see Figure A2). Overall, these differences suggest that there is significant self-selection in channel choice based on consumer attributes.
Figure A2: Fraction of multichannel customers to total customers within states
Next, we summarize the purchase behaviors of each of these groups in Table A3. The means shown in this table are calculated for each group $g$ (where $g$ is either multichannel, offline-only, or online-only) as follows: $\frac{\sum_{i=1}^{N_{g}}\sum_{t=1}^{T_{i}}X_{it}}{\sum_{i=1}^{N_{g}}T_{i}}$, where $i$ denotes a customer belonging to group $g$, $t$ denotes the year-month index, $T_{i}$ is the panel length of customer $i$, and $N_{g}$ is the total number of consumers in group $g$. As we can see, multichannel customers tend to spend more, and buy higher quantities and larger varieties of products compared to offline-only or online-only consumers. If we compare consumers who exclusively buy offline with those who exclusively buy online, we see that online-only consumers tend to spend more but purchase lower quantities and lower variety of items. This is mainly because online-only consumers tend to buy larger pack sizes of pet foods and more expensive products compared to offline-only consumers.
However, these patterns could be explained by the differences in observed consumer attributes across the three channels. To examine if this is the case, we estimate the following linear regression:
$y_{it}=CustomerType_{i}+Gender_{i}+Age_{i}+HouseholdIncome_{i}+Zipcode_{i}+YearMonth_{t}+\epsilon_{it}$ (A1)
where $y_{it}$ can denote one of the purchase metrics (e.g. Spend, Orders, UniqueCategories, etc.) of customer $i$ in month $t$. The main explanatory variable of interest is $CustomerType_{i}$, which is a dummy variable that denotes the customer’s channel choice: multichannel, offline-only, or online-only. We specify online-only as the baseline category. In addition, we include other control variables related to customer demographics (Gender, Age, Zipcode, and HouseholdIncome) for the reasons discussed earlier, i.e., to control for individual characteristics that can affect purchase behavior. Finally, we also control for time-varying unobserved shocks in period $t$ that can affect purchase behavior (such as seasonality and economic factors) using a period fixed effect ($YearMonth_{t}$).
Table A4: Cross-sectional Comparison of Customers’ Monthly Spend
DV: Spend
Model: (1) (2)
Variables
Constant 236.5 (0.412)
[0.000]
Multichannel 161.7 (0.595) 149.0 (0.610)
[0.000] [0.000]
Offline-Only -15.0 (0.509) -31.2 (0.561)
[0.000] [0.000]
Control Variables
Age Yes
Gender Yes
Zipcode Yes
Household Income Yes
Fixed-effects
YearMonth Yes
Fit statistics
Observations 123,748,697 123,748,697
R2 0.01307 0.04867
Within R2 0.01165
Clustered (Customer) standard‐errors are presented in standard brackets,
while p‑values are shown in square brackets.
Table A4 presents the difference in total spend of the multichannel and offline-only customers relative to the baseline category (i.e., online-only customers). In column (1), we use only customer type as the explanatory variable and find that spending patterns across customer types are the same as those discussed above. Next, in column (2), we estimate the same model with all the controls/fixed effects and find that the effects persist even after controlling for selection using the observed seasonality and demographic variables. Specifically, we find that, compared to offline-only consumers, multichannel consumers spend about 180.2 MCU more per month. Interestingly, the magnitude of this effect is almost the same in both models (1) and (2), i.e., the addition of controls and fixed effects did not materially change this result. Further, we replicate the specification in column (2) of Table A4 with other purchase metrics in Table A5 and find that the results are consistent with the observations from Table A3.
Table A5: Cross-sectional Comparisons of Multichannel and Offline-only relative to Online-only customers
DVs: Quantities Orders Unique Items Unique Brands Unique Subcategories Unique Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Multichannel 1.57 (0.009) 0.294 (0.001) 0.901 (0.002) 0.686 (0.001) 0.564 (0.001) 0.439 (0.001)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Offline-Only 0.260 (0.008) 0.006 (0.001) 0.373 (0.002) 0.277 (0.001) 0.225 (0.001) 0.162 (0.001)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Fixed-effects
YearMonth Yes Yes Yes Yes Yes Yes
Age Yes Yes Yes Yes Yes Yes
Gender Yes Yes Yes Yes Yes Yes
Zip code Yes Yes Yes Yes Yes Yes
Household Income Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 123,748,697 123,748,697 123,748,697 123,748,697 123,748,697 123,748,697
R2 0.02914 0.06216 0.04918 0.05339 0.05514 0.05695
Within R2 0.00382 0.01715 0.01004 0.01162 0.01193 0.01229
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
These findings align with previous literature such as Kumar and Venkatesan (2005); Thomas and Sullivan (2005); Kushwaha and Shankar (2013), who find that consumers who use multiple channels are more valuable than consumers who only purchase via one channel (though the set of channels studied in that literature are different). Given that offline-only consumers are the lowest spenders and multichannel consumers are the heaviest spenders, a natural question for the firm is whether it can increase revenues and profits by nudging offline-only consumers to also sign up for its online channels, or if more consumers sign up for online channels due to economic or environmental changes. However, the results from Table A4 could be a combination of the effect of having access to multiple channels as well as unobserved but systematic differences across consumers who use the two channels, and the reasons why they chose to adopt the online channels. For example, users who are loyal to the firm and/or heavy consumers are more likely to use both offline and online channels. As such, converting an offline-only consumer to use both offline and online channels may not necessarily lead to the same incremental gains. Further, a cross-sectional comparison doesn’t inform us about the changes in purchase behavior that occur after a consumer starts using the online channels in addition to offline.
B.2 Within-consumer Comparison
Table A6: Customers’ Spend Before/After Adopting the Online Channel
DV: Spend
Model: (1) (2)
Variables
Constant 318.4 (0.621)
[0.000]
Post_online_adoption 147.4 (0.764) 124.2 (0.782)
[0.000] [0.000]
Fixed-effects
Customer Yes
YearMonth Yes
Fit statistics
Observations 34,226,864 34,226,864
R2 0.00754 0.34036
Within R2 0.00275
Clustered (Customer) standard‐errors are presented in standard brackets,
while p‑values are shown in square brackets.
Table A7: Customers’ Purchase Behaviors Before/After Adopting Online Shopping
DVs: Quantities Orders Unique Items Unique Brands Unique Subcategories Unique Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Post_online_adoption 1.05 (0.011) 0.289 (0.001) 0.279 (0.003) 0.214 (0.002) 0.179 (0.002) 0.168 (0.001)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 34,226,864 34,226,864 34,226,864 34,226,864 34,226,864 34,226,864
R2 0.41355 0.32593 0.31713 0.29285 0.27776 0.27099
Within R2 0.00141 0.00663 0.00078 0.00085 0.00091 0.00131
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Given the shortcomings of the cross-sectional analysis discussed above, we now perform a within-consumer comparison that allows us to control for some of the selection issues, and also study the changes in consumer purchase behavior after online shopping adoption. We focus on consumers who were offline-only at the beginning of our observation period and adopted the online channel sometime during the observation period. 898,413 consumers satisfy this criterion, and we use their data for our analysis. To cleanly identify the difference in spending before and after the adoption of online shopping, we exclude the observations of the month a customer decides to adopt online shopping. We then run the following regression on these consumers’ monthly spending:
$y_{it}=Customer_{i}+\beta_{0}Post\_online\_adoption_{it}+YearMonth_{t}+\epsilon_{it},$ (A2)
where $y_{it}$ refers to an outcome variables of interest, $Customer_{i}$ is a consumer fixed-effect,
$Post\_online\_adoption_{it}$ is an indicator for whether consumer $i$ had adopted online shopping before period $t$, and $YearMonth_{t}$ is a time period fixed effect, as before. Table A6 presents the results from this regression for Spend as the outcome variable. Model (1) shows the results without the customer and year-month fixed effects and model (2) shows the results with the fixed effects. We see that, on average, customers spend 124.2 MCU more after they adopt online shopping. Note that this result is significantly smaller than the 180.2 MCU from the previous cross-sectional analysis. These results are consistent with earlier work (studying different channel transitions) that shows that consumers who transition to shopping on multiple channels from previously using only a single channel tend to spend more (Venkatesan, Kumar, and Ravishanker 2007; Ansari, Mela, and Neslin 2008) post multichannel adoption.
We also perform a similar analysis on other metrics of interest (Quantity, No. of orders, Unique items, Brands, Categories, and Subcategories) and find that all of them show a positive increase after the adoption of online shopping (see Table A7). However, as with spend, the magnitude of these effects is much smaller than what we saw in the cross-sectional analysis.
However, while the within-consumer analysis allows us to study changes in behavior for the same consumers, there is still a strong assumption that the changes we observe between the before and after periods are entirely attributable to the adoption of online shopping, and there are no other time varying factors in play causing some of the changes in behavior we observe. As such, without including an appropriate control group of non-adopters, the effect size here could still be an inflated measure of the incremental value from converting an offline-only consumer to a multichannel consumer who also uses online channels. Moreover, these analyses do not shed light on whether the different reasons for the adoption of online shopping have any significant impact on a consumer’s purchase behavior in the post-adoption period.
Appendix C Appendix to Summary Statistics
In this section, we summarize the demographics, pre-adoption, and post-adoption purchase behaviors of three customer groups for each focal event: event adopters, organic adopters, and offline-only customers.
To characterize the pre-online-adoption purchase behavior of each group of customers, we compute several consumer-level metrics, such as AvgSpendPerMonth, AvgOrdersPerMonth, AvgUniqueCategoriesPerMonth, using data from the pre-adoption period specific to each case study. We summarize these variables for each group – event adopters, organic adopters, and offline-only – by calculating the group mean as follows: $\frac{1}{N_{g}}\sum_{i=1}^{N_{g}}(\frac{1}{T_{i}^{pre}}\sum_{t=1}^{T_{i}^{pre}}X_{it})$, where $i$ denotes a customer belonging to group $g$, $t$ denotes the year-month index, $T_{i}^{pre}$ denotes the length of the consumer $i$’s panel during the pre-adoption period, and $N_{g}$ is the number of customers in group $g$. We next summarize the post-online adoption behaviors in a similar way using the post-online adoption data under each case study. The post-online adoption means for each group $g$event adopters, organic adopters, and offline-only – are calculated as follows: $\frac{1}{N_{g}}\sum_{i=1}^{N_{g}}(\frac{1}{T_{i}^{post}}\sum_{t=1}^{T_{i}^{post}}X_{it})$, where $i$ denotes a customer belonging to group $g$, $t$ denotes the year-month index, $T_{i}^{post}$ denotes the length of the consumer $i$’s panel during the post-adoption period, and $N_{g}$ is the number of customers in group $g$.
C.1 Event 1: COVID-19
In this section, we present the descriptive statistics for customers in the sample for the COVID-19. We first summarize demographics in Table A8. Next, in Tables A9 and A10, we show customer-level summary statistics of average purchase variables over the pre-adoption and post-adoption period, respectively. Finally, in Table A11, we present more detailed summary statistics (e.g., median and quartiles) for all three groups during the post-online-adoption periods (i.e., between May 1st, 2020 and April 30, 2023). Note that the counts of organic adopters and covid adopters from Table A11 slightly differ from the raw count based on our cohorts’ definition because 469 (3.66% of) organic adopters and 933 (2.81% of) covid adopters churned before 2020-05-01, and therefore are not observed in the post-COVID-19 panel. Including these customers does not affect our identification of post-online-adoption ATT, and the estimation results remain the same.
Table A8: Customer Demographics – Summary Statistics of Gender (Panel A) and Geographic State (Panel B).
Categories Covid Adopter (%) Organic Adopter (%) Offline-only (%)
Panel A: Gender
No Information 12.09 13.52 15.63
Female 59.59 57.36 46.51
Male 28.31 29.13 37.86
Total 100.00 100.00 100.00
Panel B: Geographic State
São Paulo 58.78 49.68 63.70
Minas Gerais 8.58 12.13 4.81
Rio de Janeiro 6.69 6.95 4.30
Distrito Federal 6.21 5.48 4.26
Zip_State_Missing 4.07 4.28 10.34
Santa Catarina 3.33 3.46 2.90
Goiás 2.37 3.36 2.13
Bahia 1.74 3.31 0.69
Rio Grande do Sul 3.21 2.77 2.20
Espírito Santo 1.27 2.43 0.91
Mato Grosso do Sul 0.93 1.88 1.05
Paraná 1.68 1.86 2.31
Other States 1.14 2.41 0.39
Total 100.00 100.00 100.00
Table A9: Customer-Level Summary Statistics of Age and Monthly Household Income (Panel A), Pre-online-adoption Purchase Behaviors between January 1st, 2019 and December 31st, 2019 (Panel B). Columns 1 - 3 describe statistics for Covid adopters, and Organic adopters, respectively. Column 4 calculates the difference between Covid adopters and Organic adopters, and columns 5 and 6 show the two-sample t-test and p-value results. The variable, Tenure, is the number of days between a customer’s first offline purchase in the retailer and the last date of the pre-adoption period of the covid study (i.e., 2019-12-31).
Variable Covid Adopters (Mean) Organic Adopters (Mean) Diff t-stats p-value Offline-only (Mean)
Panel A: Numerical Demographic Variables and Other Pre-online-adoption Variable
Age 43.36 41.44 1.92 15.65 0.00 45.62
Monthly Household Income 2432.11 2200.69 231.41 12.30 0.00 2033.01
Tenure(Days) 245.12 215.63 29.49 23.99 0.00 239.70
Panel B: Pre-online-adoption Monthly Purchase Behaviors
AvgSpendPerMonth 463.46 453.26 10.20 1.80 0.07 284.38
AvgQuantitiesPerMonth 4.98 4.43 0.55 6.15 0.00 2.96
AvgOrdersPerMonth 0.95 0.95 -0.00 -0.38 0.71 0.74
AvgUniqueItemsPerMonth 2.92 2.89 0.03 0.97 0.33 1.91
AvgUniqueBrandsPerMonth 2.32 2.34 -0.02 -0.92 0.36 1.55
AvgUniqueSubcategoriesPerMonth 2.02 2.04 -0.02 -1.10 0.27 1.39
AvgUniqueCategoriesPerMonth 1.65 1.66 -0.01 -0.62 0.54 1.17
AvgMarginPerMonth 162.12 162.36 -0.24 -0.12 0.91 97.52
Table A10: Customer-Level Summary Statistics of Post-online-adoption Purchase Behaviors between May 1st 2020 and April 30th 2023 (Panel A). Columns 1 - 2 describe statistics for Covid adopters and Organic adopters, respectively. Column 3 calculates the difference between Covid adopters and Organic adopters, and columns 4 and 5 show the two-sample t-test and p-value results.
Variable Covid Adopters (Mean) Organic Adopters (Mean) Diff t-stats p-value Offline-only (Mean)
Panel A: Post-online-adoption Monthly Purchase Behaviors
AvgSpendPerMonth 514.47 489.47 25.00 3.75 0.00 202.29
AvgQuantitiesPerMonth 4.88 4.11 0.77 7.58 0.00 1.85
AvgOrdersPerMonth 0.89 0.89 -0.00 -0.08 0.93 0.45
AvgUniqueItemsPerMonth 2.28 2.04 0.23 9.46 0.00 1.13
AvgUniqueBrandsPerMonth 1.78 1.63 0.15 9.18 0.00 0.91
AvgUniqueSubcategoriesPerMonth 1.52 1.39 0.13 10.35 0.00 0.80
AvgUniqueCategoriesPerMonth 1.27 1.16 0.11 10.82 0.00 0.68
AvgMarginPerMonth 175.95 157.42 18.53 8.18 0.00 76.73
Table A11: Customer-Level Summary Statistics for All Three Groups – Covid adopters, Organic Adopters, and Offline-only Customers for post-online-adoption period (i.e., between May 1st, 2020 and April 30th, 2023).
Variables Mean Std. Min 25% 50% 75% Max Count
Covid adopters
AvgSpendPerMonth 514.47 647.83 0.00 155.32 338.44 655.25 40153.38 32,313
AvgQuantitiesPerMonth 4.88 10.11 0.00 0.97 2.18 4.89 414.11 32,313
AvgOrdersPerMonth 0.89 0.81 0.00 0.36 0.69 1.17 17.81 32,313
AvgUniqueItemsPerMonth 2.28 2.41 0.00 0.78 1.57 2.97 58.62 32,313
AvgUniqueBrandsPerMonth 1.78 1.61 0.00 0.67 1.33 2.41 18.61 32,313
AvgUniqueSubcategoriesPerMonth 1.52 1.23 0.00 0.62 1.21 2.11 12.86 32,313
AvgUniqueCategoriesPerMonth 1.27 0.93 0.00 0.56 1.06 1.78 8.81 32,313
AvgMarginPerMonth 175.95 222.05 -95.36 54.56 116.51 224.52 15880.26 32,313
Organic Adopters
AvgSpendPerMonth 489.47 582.43 0.00 141.99 317.95 629.33 10709.68 12,330
AvgQuantitiesPerMonth 4.11 8.10 0.00 0.83 1.89 4.22 245.86 12,330
AvgOrdersPerMonth 0.89 0.83 0.00 0.33 0.67 1.19 11.03 12,330
AvgUniqueItemsPerMonth 2.04 2.15 0.00 0.67 1.39 2.67 30.58 12,330
AvgUniqueBrandsPerMonth 1.63 1.51 0.00 0.58 1.19 2.19 16.31 12,330
AvgUniqueSubcategoriesPerMonth 1.39 1.16 0.00 0.54 1.08 1.92 9.47 12,330
AvgUniqueCategoriesPerMonth 1.16 0.88 0.00 0.50 0.94 1.61 6.81 12,330
AvgMarginPerMonth 157.42 191.05 -24.34 46.69 101.13 203.75 5028.99 12,330
Offline-only Customers
AvgSpendPerMonth 202.29 330.96 0.00 35.98 99.68 242.19 24573.87 573,934
AvgQuantitiesPerMonth 1.85 4.59 0.00 0.30 0.77 1.85 487.17 573,934
AvgOrdersPerMonth 0.45 0.53 0.00 0.12 0.28 0.58 40.11 573,934
AvgUniqueItemsPerMonth 1.13 1.56 0.00 0.25 0.63 1.39 123.22 573,934
AvgUniqueBrandsPerMonth 0.91 1.09 0.00 0.23 0.55 1.17 48.44 573,934
AvgUniqueSubcategoriesPerMonth 0.80 0.87 0.00 0.21 0.50 1.06 19.90 573,934
AvgUniqueCategoriesPerMonth 0.68 0.70 0.00 0.19 0.44 0.94 10.00 573,934
AvgMarginPerMonth 76.73 120.02 -491.73 14.17 38.99 93.21 9834.37 573,934
C.2 Event 2: Black Friday
In this section, we present the descriptive statistics for customers in the sample for the Black Friday event. We first summarize demographics in Table A12. Next, in Tables A13 and A14, we show customer-level summary statistics of average purchase variables over the pre-adoption and post-adoption period, respectively. Finally, in Table A15, we present more detailed summary statistics (e.g., median and quartiles) for all three groups during the post-online-adoption periods (i.e., between December 1st, 2019 and October 31, 2020). Note that the counts of organic adopters and Black Friday adopters from Table A15 slightly differ from the raw count based on our cohorts’ definition because 79 (2.58% of) organic adopters and 24 (1.74% of) Black Friday adopters churned before 2019-12-01, and therefore are not observed in the post-Black Friday panel. Including these customers does not affect our identification of post-online-adoption ATT, and the estimation results remain the same.
Table A12: Customer Demographics – Summary Statistics of Gender (Panel A) and Geographic State (Panel B).
Categories Black Friday Adopter (%) Organic Adopter (%) Offline-only (%)
Panel A: Gender
No Information 14.13 13.59 15.42
Female 59.42 57.89 46.46
Male 26.45 28.52 38.12
Total 100.00 100.00 100.00
Panel B: Geographic State
São Paulo 57.97 51.91 64.83
Minas Gerais 8.48 13.10 4.69
Rio de Janeiro 7.90 6.99 4.50
Distrito Federal 4.78 4.64 4.38
Zip_State_Missing 4.93 3.76 9.90
Santa Catarina 2.54 3.40 2.60
Bahia 2.61 3.20 0.50
Goiás 2.83 3.17 2.19
Rio Grande do Sul 2.17 3.17 2.10
Espírito Santo 1.96 2.52 0.73
Paraná 1.74 1.99 2.39
Mato Grosso do Sul 1.52 1.83 1.10
Other States 0.58 0.33 0.09
Total 100.00 100.00 100.00
Table A13: Customer-Level Summary Statistics of Age and Monthly Household Income (Panel A), Pre-online-adoption Purchase Behaviors between January 1st 2019 and October 31st 2019 (Panel B). Columns 1 - 3 describe statistics for Black Friday adopters, and Organic adopters, respectively. Column 4 calculates the difference between Black Friday adopters and Organic adopters, and columns 5 and 6 show the two-sample t-test and p-value results. The variable, Tenure, is the number of days between a customer’s first offline purchase in the retailer and the last date of the pre-adoption period of the Black Friday study (i.e., 2019-10-31).
Variable Black Friday Adopters (Mean) Organic Adopters (Mean) Diff t-stats p-value Offline-only (Mean)
Panel A: Numerical Demographic Variables and Other Pre-online-adoption Variable
Age 41.86 42.72 -0.86 -2.49 0.01 47.31
Monthly Household Income 2278.60 2254.31 24.29 0.42 0.68 2046.46
Tenure(Days) 203.25 188.93 14.32 4.51 0.00 204.45
Panel B: Pre-online-adoption Monthly Purchase Behaviors
AvgSpendPerMonth 484.56 472.50 12.06 0.67 0.50 291.87
AvgQuantitiesPerMonth 5.00 4.49 0.50 2.07 0.04 3.00
AvgOrdersPerMonth 1.00 0.99 0.02 0.55 0.58 0.76
AvgUniqueItemsPerMonth 3.06 2.95 0.10 0.99 0.32 1.93
AvgUniqueBrandsPerMonth 2.44 2.39 0.05 0.72 0.47 1.57
AvgUniqueSubcategoriesPerMonth 2.13 2.08 0.04 0.74 0.46 1.41
AvgUniqueCategoriesPerMonth 1.72 1.69 0.03 0.77 0.44 1.19
AvgMarginPerMonth 165.90 168.36 -2.46 -0.36 0.72 99.52
Table A14: Customer-Level Summary Statistics of Post-online-adoption Purchase Behaviors between December 1st 2019 and October 31st 2020 (Panel A). Columns 1 - 2 describe statistics for Black Friday adopters and Organic adopters, respectively. Column 3 calculates the difference between Black Friday adopters and Organic adopters, and columns 4 and 5 show the two-sample t-test and p-value results.
Variable Black Friday Adopters (Mean) Organic Adopters (Mean) Diff t-stats p-value Offline-only (Mean)
Panel A: Post-online-adoption Monthly Purchase Behaviors
AvgSpendPerMonth 417.53 468.23 -50.70 -2.94 0.00 195.21
AvgQuantitiesPerMonth 4.37 4.55 -0.17 -0.59 0.56 2.06
AvgOrdersPerMonth 0.88 0.97 -0.10 -3.36 0.00 0.48
AvgUniqueItemsPerMonth 2.28 2.39 -0.11 -1.33 0.18 1.24
AvgUniqueBrandsPerMonth 1.83 1.90 -0.08 -1.25 0.21 0.99
AvgUniqueSubcategoriesPerMonth 1.56 1.63 -0.07 -1.55 0.12 0.87
AvgUniqueCategoriesPerMonth 1.29 1.35 -0.06 -1.66 0.10 0.74
AvgMarginPerMonth 144.17 157.64 -13.47 -2.31 0.02 73.46
Table A15: Customer-Level Summary Statistics for All Three Groups – Black Friday adopters, Organic Adopters, and Offline-only customers for post-online-adoption period (i.e., between December 1st, 2019 and October 31st, 2020).
Variables Mean Std. Min 25% 50% 75% Max Count
Black Friday adopters
AvgSpendPerMonth 417.53 485.46 0.00 118.34 280.31 536.91 6029.03 1,356
AvgQuantitiesPerMonth 4.37 10.24 0.00 0.82 2.00 4.64 219.18 1,356
AvgOrdersPerMonth 0.88 0.82 0.00 0.27 0.64 1.18 6.91 1,356
AvgUniqueItemsPerMonth 2.28 2.51 0.00 0.64 1.50 3.09 20.27 1,356
AvgUniqueBrandsPerMonth 1.83 1.82 0.00 0.55 1.27 2.55 13.82 1,356
AvgUniqueSubcategoriesPerMonth 1.56 1.40 0.00 0.55 1.18 2.18 8.91 1,356
AvgUniqueCategoriesPerMonth 1.29 1.06 0.00 0.45 1.00 1.91 7.00 1,356
AvgMarginPerMonth 144.17 167.61 -56.75 40.64 96.10 184.10 2184.19 1,356
Organic Adopters
AvgSpendPerMonth 468.23 543.57 0.00 127.57 315.86 605.36 6608.86 2,982
AvgQuantitiesPerMonth 4.55 8.55 0.00 0.82 2.09 4.91 162.82 2,982
AvgOrdersPerMonth 0.97 0.93 0.00 0.36 0.73 1.27 7.82 2,982
AvgUniqueItemsPerMonth 2.39 2.66 0.00 0.67 1.64 3.09 27.00 2,982
AvgUniqueBrandsPerMonth 1.90 1.88 0.00 0.64 1.36 2.55 17.82 2,982
AvgUniqueSubcategoriesPerMonth 1.63 1.45 0.00 0.55 1.27 2.27 12.45 2,982
AvgUniqueCategoriesPerMonth 1.35 1.09 0.00 0.55 1.09 1.91 7.73 2,982
AvgMarginPerMonth 157.64 182.93 -105.01 42.74 105.73 203.51 2349.98 2,982
Offline-only Customers
AvgSpendPerMonth 195.21 330.14 0.00 14.31 86.21 243.99 16616.77 534,871
AvgQuantitiesPerMonth 2.06 5.51 0.00 0.18 0.73 2.09 522.36 534,871
AvgOrdersPerMonth 0.48 0.62 0.00 0.09 0.27 0.64 19.73 534,871
AvgUniqueItemsPerMonth 1.24 1.84 0.00 0.18 0.64 1.55 52.91 534,871
AvgUniqueBrandsPerMonth 0.99 1.31 0.00 0.12 0.55 1.36 23.18 534,871
AvgUniqueSubcategoriesPerMonth 0.87 1.06 0.00 0.09 0.55 1.25 16.09 534,871
AvgUniqueCategoriesPerMonth 0.74 0.85 0.00 0.09 0.45 1.09 9.82 534,871
AvgMarginPerMonth 73.46 120.07 -331.15 5.58 33.33 93.25 6313.29 534,871
C.3 Event 3: Loyalty Program
In this section, we present the descriptive statistics for customers in the sample for the loyalty program analysis. We first summarize demographics in Table A16. Next, in Tables A17 and A18, we show customer-level summary statistics of average purchase variables over the pre-adoption and post-adoption period, respectively. Finally, in Table A19, we present more detailed summary statistics (e.g., median and quartiles) for all three groups during the post-online-adoption periods (i.e., between October 1st, 2023 and March 31, 2024)
Table A16: Customer Demographics – Summary Statistics of Gender (Panel A) and Geographic State (Panel B). Our data providers label around 45% of overall customers as no gender information in this study.
Categories Loyalty Program Adopter (%) Organic Adopter (%) Offline-only (%)
Panel A: Gender
No Information 50.16 48.30 42.85
Female 30.57 32.52 31.02
Male 19.27 19.17 26.13
Total 100.00 100.00 100.00
Panel B: Geographic State
São Paulo 35.99 43.29 55.28
Zip_State_Missing 32.99 19.92 21.93
Minas Gerais 5.30 6.69 3.84
Rio de Janeiro 3.89 4.39 3.15
Santa Catarina 6.25 4.14 3.04
Distrito Federal 2.62 3.86 3.09
Espírito Santo 2.23 2.59 1.28
Ceará 2.04 2.37 0.83
Pernambuco 1.08 2.10 0.94
Bahia 1.21 1.99 1.12
Goiás 1.34 1.83 1.63
Mato Grosso do Sul 0.70 1.22 0.95
Other States 4.34 5.62 2.93
Total 100.00 100.00 100.00
Table A17: Customer-Level Summary Statistics of Age and Monthly Household Income (Panel A), Pre-online-adoption Purchase Behaviors between January 1st, 2023 and May 31st, 2023 (Panel B). Columns 1 - 3 describe statistics for Loyalty Program Adopters, and Organic adopters, respectively. Column 4 calculates the difference between Loyalty Program Adopters and Organic adopters, and columns 5 and 6 show the two-sample t-test and p-value results. The variable, Tenure, is the number of days between a customer’s first offline purchase in the retailer and the last date of the pre-adoption period of the loyalty program study (i.e., 2023-05-31).
Variable Loyalty Program Adopters (Mean) Organic Adopters (Mean) Diff t-stats p-value Offline-only (Mean)
Panel A: Numerical Demographic Variables and Other Pre-online-adoption Variable
Age 44.57 43.43 1.14 2.78 0.01 47.01
Monthly Household Income 1959.34 2072.83 -113.49 -2.24 0.02 1975.35
Tenure(Days) 786.58 827.54 -40.96 -3.09 0.00 915.96
Panel B: Pre-online-adoption Monthly Purchase Behaviors
AvgSpendPerMonth 576.42 444.70 131.71 8.49 0.00 332.57
AvgQuantitiesPerMonth 4.31 3.25 1.06 6.44 0.00 2.62
AvgOrdersPerMonth 1.04 0.83 0.20 9.64 0.00 0.70
AvgUniqueItemsPerMonth 2.55 2.05 0.50 7.96 0.00 1.64
AvgUniqueBrandsPerMonth 2.00 1.64 0.36 8.48 0.00 1.31
AvgUniqueSubcategoriesPerMonth 1.70 1.41 0.29 8.76 0.00 1.14
AvgUniqueCategoriesPerMonth 1.42 1.20 0.22 8.57 0.00 0.98
AvgMarginPerMonth 198.59 164.46 34.12 6.04 0.00 125.81
Table A18: Customer-Level Summary Statistics of Post-online-adoption Purchase Behaviors between October 1st, 2023 and March 31st, 2024 (Panel A). Columns 1 - 2 describe statistics for Loyalty Program Adopters and Organic adopters, respectively. Column 3 calculates the difference between Loyalty Program Adopters and Organic adopters, and columns 4 and 5 show the two-sample t-test and p-value results.
Variable Loyalty Program Adopters (Mean) Organic Adopters (Mean) Diff t-stats p-value Offline-only (Mean)
Panel A: Post-online-adoption Monthly Purchase Behaviors
AvgSpendPerMonth 696.00 524.96 171.04 9.42 0.00 284.45
AvgQuantitiesPerMonth 5.22 3.58 1.64 7.70 0.00 2.18
AvgOrdersPerMonth 1.36 1.02 0.34 12.57 0.00 0.59
AvgUniqueItemsPerMonth 2.75 2.05 0.69 10.81 0.00 1.36
AvgUniqueBrandsPerMonth 2.16 1.65 0.51 11.62 0.00 1.09
AvgUniqueSubcategoriesPerMonth 1.84 1.42 0.41 12.33 0.00 0.96
AvgUniqueCategoriesPerMonth 1.54 1.22 0.32 12.27 0.00 0.83
AvgMarginPerMonth 223.51 180.19 43.32 6.86 0.00 110.25
Table A19: Customer-Level Summary Statistics for All Three Groups – Loyalty Program Adopters, Organic Adopters, and Offline-only Customers for post-online-adoption period (i.e., between October 1st, 2023 and March 31st, 2024).
Variables Mean Std. Min 25% 50% 75% Max Count
Loyalty Program Adopters
AvgSpendPerMonth 696.00 810.94 0.00 210.82 452.64 876.14 9712.41 1,567
AvgQuantitiesPerMonth 5.22 11.23 0.00 1.00 2.67 5.45 262.33 1,567
AvgOrdersPerMonth 1.36 1.31 0.00 0.50 1.00 1.83 20.50 1,567
AvgUniqueItemsPerMonth 2.75 2.89 0.00 1.00 2.00 3.67 43.33 1,567
AvgUniqueBrandsPerMonth 2.16 1.97 0.00 0.83 1.67 2.83 27.67 1,567
AvgUniqueSubcategoriesPerMonth 1.84 1.48 0.00 0.83 1.50 2.67 15.17 1,567
AvgUniqueCategoriesPerMonth 1.54 1.11 0.00 0.67 1.33 2.17 9.00 1,567
AvgMarginPerMonth 223.51 272.37 -44.82 66.43 144.49 280.82 3800.20 1,567
Organic Adopters
AvgSpendPerMonth 524.96 674.20 0.00 132.98 336.48 678.11 21482.62 16,746
AvgQuantitiesPerMonth 3.58 7.69 0.00 0.67 1.67 3.83 358.33 16,746
AvgOrdersPerMonth 1.02 0.99 0.00 0.33 0.83 1.33 15.17 16,746
AvgUniqueItemsPerMonth 2.05 2.38 0.00 0.67 1.33 2.67 41.67 16,746
AvgUniqueBrandsPerMonth 1.65 1.64 0.00 0.50 1.17 2.17 22.33 16,746
AvgUniqueSubcategoriesPerMonth 1.42 1.25 0.00 0.50 1.17 2.00 14.33 16,746
AvgUniqueCategoriesPerMonth 1.22 0.97 0.00 0.50 1.00 1.67 9.00 16,746
AvgMarginPerMonth 180.19 235.65 -1078.75 45.07 114.10 233.52 7693.62 16,746
Offline-only Customers
AvgSpendPerMonth 284.45 484.92 0.00 36.70 146.34 356.93 37203.48 396,396
AvgQuantitiesPerMonth 2.18 5.39 0.00 0.33 1.00 2.17 859.33 396,396
AvgOrdersPerMonth 0.59 0.69 0.00 0.17 0.40 0.83 50.67 396,396
AvgUniqueItemsPerMonth 1.36 1.90 0.00 0.33 0.83 1.80 114.83 396,396
AvgUniqueBrandsPerMonth 1.09 1.29 0.00 0.25 0.67 1.50 56.00 396,396
AvgUniqueSubcategoriesPerMonth 0.96 1.02 0.00 0.20 0.67 1.33 22.67 396,396
AvgUniqueCategoriesPerMonth 0.83 0.82 0.00 0.17 0.67 1.17 10.67 396,396
AvgMarginPerMonth 110.25 180.70 -690.21 14.02 56.76 138.84 11025.70 396,396
C.4 Loyalty Program Control Variables
Below, we provide the definition of time-varying control variables related to customers’ participation in the loyalty program. These variables ($\boldsymbol{x}_{it}$) are included in the DID and event study analysis for the loyalty program study.
  • $\bullet$
    $isActivated_{it}$: 1 if customer i activates the offer during month $t$ and 0 otherwise. Use to compare within-activators, how much spend differs between months with/without activation.
  • $\bullet$
    $isFirstActivation_{it}$: A binary variable equal to 1 if customer $i$ activates the offer for the first time in month $t$, and 0 otherwise. This variable isolates the novelty (first-time) in the month of initial activation, relative to subsequent activation months.
  • $\bullet$
    $Months\_Since\_First\_Activation_{it}$: A non-negative integer indicating the number of months elapsed since customer $i$’s first activation of the offer, as of month $t$; equals 0 for months prior to the first activation. For example, if the first activation occurs in August 2023, then
    $Months\_Since\_First\_Activation_{i,September\_2023}=1$, and so on.
  • $\bullet$
    %$\%Month\_WithActivation_{it}$: The proportion of months in which customer $i$ has activated the offer up to (but not including) month $t$, calculated as the total number of activation months up to $t$ divided by the number of months since first activation (excluding $t$ itself). For the first month of activation, this variable is defined as 0 (since the denominator is 0); all first-month effects are captured by $isFirstActivation_{it}$. Starting from the second month post-activation, this variable reflects the historical activation rate prior to month $t$.
Appendix D Appendix to the Profitability Analysis: Missing Value Imputation
As discussed in the main text, we have data on prices, margins, and costs for each item sold since January 2020; however, we lack cost data for 2019 due to issues with the data provider’s upgrade to a new data management platform in 2020. Thus, we need to impute the missing costs and margins for the 2019 data. We do this in two steps.
Step 1: We impute the cost of each item in 2019 via an inflation-adjusting approach by using the unit cost and the inflation data. Based on conversations with our retail collaborator, we learned that procurement contracts are typically updated on an annual basis and that unit cost increases for the majority of items are largely due to inflation (or in sync with inflation rates). Hence, we calculate the average cost for each item $j$ in each channel $c$ (i.e., online and offline) and each year (i.e., 2020, 2021, and 2022), denoted as $cost_{jc,2020},cost_{jc,2021}$, $cost_{jc,2022}$.  [^33] We focus on cost value over a 36-month period between 2020 and 2022. We do not consider the 2023 cost value as the basis for cost imputation, as we do not have complete cost data for the full year of 2023 and cannot compute the yearly average cost.
We choose Brazil’s Extended National Consumer Price Index (IPCA) in June (the midpoint of the year) as the average level of IPCA for each year, and obtain a set of unit cost estimates for each item $j$ in channel $c$, denoted as $s_{jc}$, by adjusting the ratio of IPCA index in Equation A3.343434We obtain Brazil’s Broad National Consumer Index (IPCA) from the Brazil Institute of Geography and Statistics (IBGE). The IPCA data is available under the Historic Serie section of the following URL:https://www.ibge.gov.br/en/statistics/economic/prices-and-costs.
$s_{jc}=\{cost_{jc,2020}\times\frac{ICPA_{2019}}{ICPA_{2020}},cost_{jc,2021}\times\frac{ICPA_{2019}}{ICPA_{2021}},cost_{jc,2022}\times\frac{ICPA_{2019}}{ICPA_{2022}}\}$ (A3)
We can then take the average of the elements of set $s_{jc}$ for each item-channel combination and calculate the imputed cost of item $j$ in channel $c$ in 2019, as $cost_{jc,2019}$.
Step 2: Since we observe the price, we can then calculate the margin for each item-channel combination in 2019 as
$margin_{jc}=price_{jc}\times(1-AvgTaxRate)-cost_{jc,2019}.$ (A4)
Here, we impute missing margin using an average sales tax rate of 16% for imputing the missing margin. This average sales tax rate, 16% (or one minus average sales tax rate) is obtained by calculating the retailer’s total monthly margin plus cost divided by total monthly sales over a 36-month period between 2020-01-01 and 2022-12-31. This resulting 16% closely matches Brazil’s average sales tax rate of 17% Trading Economics (2025), and our retail collaborator has confirmed that this method is appropriate for imputing missing margin. Finally, we aggregate these margins at the customer-month level to get the customer-month panel data for the profitability measures in 2019.
Appendix E Appendix to Mechanisms, Theoretical Underpinnings and Managerial Implications
E.1 Appendix to the Share of Offline-Online Spend
Table A20: DID Analysis - Offline Spend (Adopters vs. Offline-only)
Study: Covid Black Friday Loyalty Program
DV: Offline Spend
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 280.1 (0.557) 291.7 (0.586) 332.6 (0.800)
[0.000] [0.000] [0.000]
Adopter 177.2 (2.885) 177.8 (9.177) 123.4 (4.420)
[0.000] [0.000] [0.000]
Post -62.90 (0.495) -96.95 (0.426) -51.36 (0.632)
[0.000] [0.000] [0.000]
Adopter_Post -114.3 (2.442) -110.5 (2.553) -116.7 (7.217) -118.0 (6.661) -189.8 (4.585) -116.9 (3.827)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 22,204,608 22,204,608 9,505,999 9,505,999 4,403,449 4,403,449
R2 0.00501 0.36735 0.00899 0.41478 0.00480 0.47768
Within R2 0.00068 0.00016 0.00165
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A21: DID Main Analysis of Offline Spend – Event Adopters vs. Organic Adopters
Study: Covid Black Friday Loyalty Program
DV: Offline Spend
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 432.8 (5.325) 459.1 (10.48) 444.7 (4.415)
[0.000] [0.000] [0.000]
Covid_Adopter 32.84 (6.281)
[0.000]
Post -199.9 (4.462) -205.8 (8.770) -229.8 (4.490)
[0.000] [0.000] [0.000]
Covid_Adopter * Post 32.24 (5.283) 36.13 (5.391)
[0.000] [0.000]
BlackFriday_Adopter 32.01 (20.72)
[0.122]
BlackFriday_Adopter * Post -23.82 (15.26) -7.116 (13.96)
[0.118] [0.611]
Loyalty_Adopter 131.7 (19.08)
[0.000]
Loyalty_Adopter * Post -210.9 (19.02) -12.09 (18.59)
[0.000] [0.515]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 76,842 76,842 197,658 197,658
R2 0.01177 0.36501 0.02459 0.40155 0.05133 0.43555
Within R2 0.00012 $9.15\times 10^{-6}$ 0.01910
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A22: DID Main Analysis of Share of Offline Spend – Event Adopters vs. Organic Adopters
Study: Covid Black Friday Loyalty Program
DV: Share of Offline Spend
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 1.000 ($2.28\times 10^{-15}$) 1.000 ($1.69\times 10^{-16}$) 1.000 ($3.29\times 10^{-17}$)
[0.000] [0.000] [0.000]
Covid_Adopter 0.000 (0.000)
[0.000]
Post -0.5109 (0.0034) -0.4332 (0.0073) -0.4531 (0.0039)
[0.000] [0.000] [0.000]
Covid_Adopter * Post 0.0811 (0.0040) 0.0698 (0.0042)
[0.000] [0.000]
BlackFriday_Adopter 0.000 (0.000)
[0.000]
BlackFriday_Adopter * Post 0.0977 (0.0124) 0.0922 (0.0120)
[0.000] [0.000]
Loyalty_Adopter 0.000 (0.000)
[0.000]
Loyalty_Adopter * Post 0.0002 (0.0115) 0.0018 (0.0114)
[0.987] [0.889]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 949,448 949,448 42,299 42,299 110,339 110,339
R2 0.18119 0.50739 0.25515 0.53685 0.28417 0.57059
Within R2 0.00140 0.00528 0.00904
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A23: Differential Trends in the Share of Offline Spend (Estimation Results for Equation A5) – Event Adopters vs. Organic Adopters
DV: Share of Offline Spend
Model: Covid Black Friday Loyalty Program
Variables
Organic_Adopter 0.4833 (0.0084) 0.5930 (0.0285) 0.5662 (0.0093)
[0.000] [0.000] [0.000]
Event_Adopter 0.0807 (0.0119) 0.1424 (0.0404) 0.0306 (0.0132)
[0.000] [0.000] [0.020]
t -0.0020 (0.0004) -0.0108 (0.0048) -0.0020 (0.0031)
[0.000] [0.025] [0.518]
Event_Adopter * t 0.0012 (0.0006) -0.0109 (0.0068) 0.0013 (0.0043)
[0.046] [0.109] [0.763]
Fit statistics
Observations 72 22 12
R2 0.81922 0.69953 0.72620
Adjusted R2 0.81124 0.64945 0.62353
Standard errors in parentheses; p‑values in square brackets.
Table A24: Differential Trend in the Share of Offline Spend by Age – Covid Adopters vs. Organic Adopters
DV: Share of Offline Spend
Model: (1) Senior Adopters (2) Young Adopters
Variables
Organic_Adopter 0.4947 (0.0092) 0.4728 (0.0081)
[0.000] [0.000]
Covid_Adopter 0.0706 (0.0130) 0.0882 (0.0114)
[0.000] [0.000]
t -0.0019 (0.0005) -0.0022 (0.0004)
[0.001] [0.000]
Covid_Adopter * t 0.0020 (0.0006) 0.0004 (0.0006)
[0.002] [0.523]
Fit statistics
Observations 72 72
R2 0.79964 0.82429
Adjusted R2 0.79080 0.81654
Standard errors are shown in parentheses; p-values are shown in square brackets.
Table A20 presents the DID results of offline spend between adopters and offline-only customers. Table A21 and A22 shows the DID results of offline spend and share of offline spend between event adopters and organic adopters. Table A23 presents the estimation results for the linear trend model by comparing a type of event adopter with organic adopters. Table A24 presents the changes in share of offline spend by different age groups, showing senior adopters tend to have a slower rate of shifting their spend to the online channel.
In §7.1, we find that the share of offline spending is persistently higher for covid adopters compared to organic adopters in the post-adoption period, and the downward trend is steeper for organic adopters, indicating that covid adopters shift their spending to online channels slower than organic adopters. To formally test whether these differential trends in share of offline spend are systematic, we estimate the following model with linear time trend on the post-adoption monthly panel data for both groups:
$y_{gt}=\zeta+\rho_{0}t+\rho_{1}\mathbbm{I}\{Event\_Adopter_{g}\}+\rho_{2}\mathbbm{I}\{Event\_Adopter_{g}\}\times t+\epsilon_{gt}.$ (A5)
Here, $y_{gt}$ is the fraction of the offline spend relative to the total spend (ranging from 0 to 1) for group $g$ in period $t$. $t$ is the numerical time index, and we set the first post-adoption month of each study as $t=0$. The results for covid, Black Friday and loyalty program studies from this regression are shown in Table A23. The estimate of the difference in trend ($\rho_{2}$) between covid and organic adoption cohorts is 0.0012 (or 0.12%). This finding suggests that although the different reasons (i.e., organic vs. exogenous shock) for adopting online shopping do not seem to have much impact on how much consumers spend with the firm in total, they do cause differences in where they spend it. Figure A3 plots the fraction of offline spending in the post-treatment period separately for older and younger adopters.
Figure A3: Differential Trend Visualization by Age – Covid Adopters vs. Organic Adopters
(a) Senior Adopters (Age 41$\geq$ 41≥ 41)
(b) Young Adopters (Age <41<41< 41)
E.2 Appendix to the Other Metrics
Finally, to further explore the differences in the purchase behavior of different groups of adopters relative to organic adopters post-online-adoption, we estimate the same DID model (as shown in Equation (3) and (4)) on other metrics of interest (Quantity, Number of orders, Unique items, Brands, Categories, and Subcategories). For each study (Covid, Black Friday, and Loyalty Program), we present results for both total and offline metrics. Specifically, total metrics are reported in Tables A25, A27, and A29, while offline metrics appear in Tables A26, A28, and A30, respectively.
To begin with, we find some differences between covid and organic groups. Covid adopters tend to place fewer orders but purchase a wider variety of items, brands, and categories, compared to organic adopters. This is consistent with their heavier use of the offline channel since consumers who use offline channels generally tend to buy more frequently and also buy a wider variety of products (Chintala, Liaukonytė, and Yang 2023). Next, we find Black Friday adopters consistently exhibit lower purchase overall across all total metrics than organic adopters, but show insignificant differences in offline metrics. This suggests, short-term online promotional event are insufficient to raise adopters’ purchase levels in online channel to those observed among organic adopters. Lastly, we do not observe significant differences in total and offline purchase metrics between loyalty program adopters and organic adopters, after controlling for loyalty program’s participation. This suggests that when adopters’ motivation for online adoption is to access long-term discounts (available for both online and offline), their purchase variety and channel utilization remain similar to organic adopters. However, they may be leveraging the loyalty program to maintain similar consumption levels while spending less, thus potentially becoming less profitable to the firm.
Appendix to Additional Covid Analysis
Table A25 and A26 presents DID results for the total and offline metrics under the covid analysis.
Table A25: Covid adopters vs. Organic Adopters Results – Other Total Metrics
DVs: Quantity Orders Unique Items Unique Brands Unique Subcategories Unique Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Covid_Adopter * Post 0.026 (0.099) -0.029 (0.010) 0.039 (0.028) 0.038 (0.020) 0.039 (0.016) 0.032 (0.012)
[0.792] [0.004] [0.163] [0.057] [0.015] [0.008]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949
R2 0.49396 0.37110 0.36977 0.33648 0.31704 0.30595
Within R2 $2.22\times 10^{-7}$ $2.38\times 10^{-5}$ $4.7\times 10^{-6}$ $8.66\times 10^{-6}$ $1.42\times 10^{-5}$ $1.54\times 10^{-5}$
Clustered (Customer) standard‐errors in parentheses; p‑values in square brackets.
Table A26: Covid adopters vs. Organic Adopters Results – Other Offline Metrics
DVs: Quantity Orders Unique Items Unique Brands Unique Subcategories Unique Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Covid_Adopter * Post -0.028 (0.082) 0.031 (0.008) 0.090 (0.027) 0.087 (0.019) 0.074 (0.015) 0.059 (0.012)
[0.627] [0.000] [0.000] [0.000] [0.000] [0.000]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949
R2 0.45211 0.36623 0.34965 0.32551 0.31257 0.30808
Within R2 $4.03\times 10^{-7}$ $4.13\times 10^{-5}$ $2.94\times 10^{-5}$ $5.17\times 10^{-5}$ $5.74\times 10^{-5}$ $6.07\times 10^{-5}$
Clustered (Customer) standard‐errors in parentheses; p‑values in square brackets.
Appendix to Additional Black Friday Analysis
Table A27 and A28 presents the DID estimates of other total metrics and offline metrics of interest under the Black Friday analysis.
Table A27: Black Friday Adopters vs. Organic Adopters Results - Other Total Metrics
DVs: Quantity Orders Unique Items Unique Brands Unique Subcategories Unique Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
BlackFriday_Adopter * Post -0.653 (0.219) -0.114 (0.025) -0.232 (0.076) -0.140 (0.056) -0.125 (0.044) -0.099 (0.033)
[0.002] [0.000] [0.003] [0.014] [0.005] [0.004]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 76,842 76,842 76,842 76,842 76,842 76,842
R2 0.52977 0.42114 0.41307 0.39483 0.37340 0.35736
Within R2 0.00035 0.00070 0.00029 0.00019 0.00023 0.00025
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A28: Black Friday Adopters vs. Organic Adopters Results - Other Offline Metrics
DVs: Quantity Orders Unique Items Unique Brands Unique Subcategories Unique Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
BlackFriday_Adopter * Post -0.211 (0.205) -0.001 (0.022) -0.039 (0.076) 0.013 (0.056) 0.011 (0.044) 0.017 (0.033)
[0.308] [0.967] [0.610] [0.812] [0.804] [0.591]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 76,842 76,842 76,842 76,842 76,842 76,842
R2 0.51966 0.44015 0.40718 0.39489 0.37993 0.37063
Within R2 $5.29\times 10^{-5}$ $1.13\times 10^{-7}$ $9.13\times 10^{-6}$ $1.93\times 10^{-6}$ $1.92\times 10^{-6}$ $8.1\times 10^{-6}$
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Appendix to Additional Loyalty Program Analysis
Table A29 and A30 presents the DID estimates of the other total metrics and offline metrics of interest under the loyalty program analysis.
Table A29: Loyalty Program Adopters vs. Organic Adopters Results - Other Total Metrics
DVs: Quantity Orders Unique Items Unique Brands Unique Subcategories Unique Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Loyalty_Adopter * Post -0.048 (0.201) -0.047 (0.032) -0.071 (0.077) -0.061 (0.053) -0.040 (0.041) -0.026 (0.031)
[0.811] [0.142] [0.356] [0.250] [0.330] [0.402]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 197,658 197,658 197,658 197,658 197,658 197,658
R2 0.56438 0.44676 0.44267 0.40084 0.37005 0.35296
Within R2 0.01429 0.03068 0.02502 0.02647 0.02701 0.02610
Clustered (Customer) standard‐errors are presented in standard brackets; p‑values are shown in square brackets.
Table A30: Loyalty Program Adopters vs. Organic Adopters Results - Other Offline Metrics
DVs: Quantity Orders Unique Items Unique Brands Unique Subcategories Unique Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Loyalty_Adopter * Post 0.141 (0.206) 0.018 (0.026) 0.090 (0.075) 0.087 (0.051) 0.074 (0.039) 0.059 (0.030)
[0.492] [0.488] [0.708] [0.752] [0.608] [0.406]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 197,658 197,658 197,658 197,658 197,658 197,658
R2 0.52637 0.44029 0.43107 0.39543 0.36797 0.35526
Within R2 0.01064 0.02050 0.01867 0.02049 0.02178 0.02205
Clustered (Customer) standard‐errors are presented in standard brackets; p‑values are shown in square brackets.
E.3 Appendix to Forward buying
Now, we discuss how to set up the event study regression for testing the forward buying under Black Friday and loyalty program analysis. For the Black Friday analysis, we define event time 0 as November 2019, given that Black Friday occurred on 2019-11-29. The pre-adoption period is from 2019-01-01 to 2019-10-31 (event time -10 to -1), and the post-adoption period is from 2019-12-01 to 2020-10-31 (event time 1 to 11). By estimating the event study regression in Equation (5), we present the estimation result in Table A31 and show that Black Friday adopters spend significantly higher than organic adopters during the adoption month, but spend significantly lower during the post-adoption periods, which indicates a pattern of forward buying.
Table A31: Event Study for Spend – Black Friday Adopters vs. Organic Adopters
DV: Spend
Model: (1)
Variables
BlackFriday_Adopter $\times$ Lag –10 –39.54 (50.73)
[0.436]
BlackFriday_Adopter $\times$ Lag –9 –44.26 (38.91)
[0.255]
BlackFriday_Adopter $\times$ Lag –8 –26.81 (35.61)
[0.452]
BlackFriday_Adopter $\times$ Lag –7 –25.93 (30.61)
[0.397]
BlackFriday_Adopter $\times$ Lag –6 5.206 (34.45)
[0.880]
BlackFriday_Adopter $\times$ Lag –5 –4.364 (27.55)
[0.874]
BlackFriday_Adopter $\times$ Lag –4 –18.65 (27.25)
[0.494]
BlackFriday_Adopter $\times$ Lag –3 –9.695 (25.51)
[0.704]
BlackFriday_Adopter $\times$ Lag –2 –38.55 (23.94)
[0.107]
BlackFriday_Adopter $\times$ Adoption Month 128.6 (32.46)
[0.000]
BlackFriday_Adopter $\times$ Lead 1 –134.3 (25.36)
[0.000]
BlackFriday_Adopter $\times$ Lead 2 –114.4 (25.65)
[0.000]
BlackFriday_Adopter $\times$ Lead 3 –105.5 (25.58)
[0.000]
BlackFriday_Adopter $\times$ Lead 4 –88.21 (28.63)
[0.002]
BlackFriday_Adopter $\times$ Lead 5 –55.82 (28.28)
[0.048]
BlackFriday_Adopter $\times$ Lead 6 –126.9 (27.95)
[0.000]
BlackFriday_Adopter $\times$ Lead 7 –88.91 (28.21)
[0.002]
BlackFriday_Adopter $\times$ Lead 8 –69.56 (27.24)
[0.011]
BlackFriday_Adopter $\times$ Lead 9 –25.79 (28.13)
[0.359]
BlackFriday_Adopter $\times$ Lead 10 –90.03 (28.42)
[0.002]
BlackFriday_Adopter $\times$ Lead 11 –67.00 (28.53)
[0.019]
Fixed-effects
Customer Yes
YearMonth Yes
Fit statistics
Observations 81,283
R2 0.42086
Within R2 0.00211
Clustered (Customer) standard-errors are presented in standard brackets,
while p‑values are shown in square brackets.
Regarding the loyalty program analysis, for both loyalty program adopters and organic adopters, the month of adoption is labeled as event time 0. For example, event time 0 is June 2023 for organic adopters who adopted online shopping in that month, and September 2023 for loyalty program adopters who adopted online shopping in that month.
We compare each group’s spending at event time 0 to their most recent pre-adoption outcome, specifically those observed in May 2023. This period serves as a common baseline for pre-treatment behaviors across all adopter groups. As defined in §5.1.3, the post-adoption periods span from October 2023 and March 2024, and we label October 2023 as the event time 1 for both organic and loyalty program adopters and label all the following months until March 2024 as event time 6. The event study coefficient at time 1 captures the difference in the monthly spend at October 2023 between two groups relative to the pre-adoption difference in the reference period May 2023.
We first compare the spend of organic and loyalty program adopters during their adoption month relative to the pre-adoption reference period (i.e., spend in May 2023), and Table A32 shows that loyalty program adopters tend to spend 177.7 MCU more than organic adopters during the adoption month. Next, we notice that both loyalty program and organic adopters can activate the monthly loyalty offer during the post-adoption period. To test the post-adoption differences in purchase behavior, we estimate the event study regression in Equation (5) with control variables for the customers’ activations. Table A33 shows that loyalty program adopters spend significantly less than organic adopters in most of the post-adoption periods.
Table A32: Event Study for Spend during the Adoption Month - Loyalty Program Adopters vs. Organic Adopters
DV: Spend
Model: (1)
Variables
Loyalty_Adopter $\times$ Lag -5 -11.43 (26.52)
[0.672]
Loyalty_Adopter $\times$ Lag -4 -29.64 (25.37)
[0.245]
Loyalty_Adopter $\times$ Lag -3 0.7290 (24.62)
[0.976]
Loyalty_Adopter $\times$ Lag -2 -25.12 (26.40)
[0.340]
Loyalty_Adopter $\times$ Adoption Month 177.7 (31.54)
[0.000]
Fixed-effects
Customer Yes
Event_Time Yes
Fit statistics
Observations 109,857
R2 0.53391
Within R2 0.00111
Clustered (Customer) standard-errors are presented in standard brackets,
while p‑values are shown in square brackets.
Table A33: Event Study for Spend during the Post-Online-Adoption Periods – Loyalty Program Adopters vs. Organic Adopters
DV: Spend
Model: (1)
Variables
Loyalty_Adopter $\times$ Lag -5 -11.45 (26.52)
[0.670]
Loyalty_Adopter $\times$ Lag -4 -29.64 (25.37)
[0.245]
Loyalty_Adopter $\times$ Lag -3 0.7526 (24.62)
[0.975]
Loyalty_Adopter $\times$ Lag -2 -25.10 (26.40)
[0.340]
Loyalty_Adopter $\times$ Lead 1 -102.5 (29.56)
[0.001]
Loyalty_Adopter $\times$ Lead 2 -34.48 (28.49)
[0.224]
Loyalty_Adopter $\times$ Lead 3 -45.26 (31.24)
[0.144]
Loyalty_Adopter $\times$ Lead 4 -73.56 (30.45)
[0.015]
Loyalty_Adopter $\times$ Lead 5 -69.94 (33.10)
[0.037]
Loyalty_Adopter $\times$ Lead 6 -75.00 (35.29)
[0.032]
Loyalty program controls Yes
Fixed-effects
Customer Yes
YearMonth Yes
Fit statistics
Observations 197,658
R2 0.47130
Within R2 0.02699
Clustered (Customer) standard-errors are presented in standard brackets,
while p‑values are shown in square brackets.
Next, we run the event study regression for offline spend, and Figure A4 presents the event study figure for the offline spend of Black Friday and loyalty program study.
Finally, we follow a similar approach described above to test for covid adopters. Figure A5 shows that there is no forward buying for covid adopters relative to organic adopters.
Figure A4: Event Study for Offline Spend
(a) Black Friday
(b) Loyalty Program
Figure A5: Event Study for Spend – Covid Adopters vs. Organic Adopters during the month of adoption
E.4 Appendix to Managerial Implications
Table A34 presents an additional DID result for profit under Black Friday analysis, which are inputs for the discussion of managerial implications. Instead of estimating Equation (1) and (2), with a single indicator $\mathbbm{I}\{Adopter_{i}\}$, we now include two separate dummies – $\mathbbm{I}\{Event\_Adopter_{i}\}$ and $\mathbbm{I}\{Organic\_Adopter_{i}\}$ – each interacted with the post-adoption indicator. This allows us to estimate the ATT of online adoption on profit for organic adopters and event adopters, relative to offline-only customers.
Table A34: DID Results of Profit – Black Friday Adopters, Organic Adopters, and Offline-only Customers
DV: Profit
Model: (1) (2)
Variables
Constant 96.13 (0.187)
[0.000]
Organic_Adopter 58.55 (3.523)
[0.000]
Post -22.83 (0.152)
[0.000]
Organic_Adopter * Post 26.14 (3.166) 23.57 (3.160)
[0.000] [0.000]
BlackFriday_Adopter 65.89 (5.809)
[0.000]
BlackFriday_Adopter * Post 5.621 (4.164) 7.813 (4.014)
[0.177] [0.051]
Fixed-effects
Customer Yes
YearMonth Yes
Fit statistics
Observations 9,505,999 9,505,999
R2 0.00451 0.34750
Within R2 $2.88\times 10^{-5}$
Clustered (Customer) standard-errors are presented in standard brackets,
while p‑values are shown in square brackets.
Appendix F Appendix to Validity and Robustness Checks
F.1 Appendix to Parallel Trend Tests
F.1.1 Parallel Trend Test for Covid Analysis - Total Spend
Figure A6: Parallel Trend Assumption - Total Spend
(a) Visualization
(b) Adopters vs. Non-Adopters
(c) Covid Adopters vs. Organic Adopters
Figure A6 provides parallel trend visualization and test for total spend. Figure 6(a) visualizes the average monthly spend of all three groups in both the pre- and post-treatment periods. For each group, this average monthly spend is calculated by dividing the total spend of customers belonging to the group in a month by the number of customers who have not churned by the end of that month. As we can see, the spending of all three groups evolves parallelly during the pre-adoption period from 2019-01-01 to 2019-12-31, and there are no differential trends in spend across the three groups. [^35] Each line in the parallel trend plots shows the average monthly spending for a given group of customers (organic adopters, covid adopters, and offline-only) over the observation period. At any given period $t$, the averaging is done over the set of active consumers in that period. Both at the beginning and end of the observation period, we have a (relatively) small number of customers in all three groups because of customer acquisition and churn: (1) Not all customers in each group initiate their first purchase at the beginning of the time series (e.g., Jan and Feb 2019), i.e., many customers joined after early 2019, and (2) Customers in all three groups continue to churn in the post-treatment period, and many are no longer in our panel towards the end of the observation period (e.g., March and April 2023). Further, the consumers who were with the firm from early 2019 and/or those who did not churn by mid-2023 tend to be loyal consumers (who spend more on average). Since the early and late observations largely consist of such consumers, we observe higher average spending in these time periods. However, the existence of such spikes does not invalidate the parallel trends observed in the pre-treatment period.
Then, Figure 6(b) presents the pre-trend between adopters (both covid and organic adopters) and offline-only customers (i.e., Non-Adopters), which shows some violation of parallel pre-trend. We use the honest DID approach from Rambachan and Roth (2023) and provide bounded ATT estimates for the total spend in Web Appendix §F.2. The incremental spend due to online adoption is still significant and robust, even in the presence of a large violation of parallel trend.
Finally, Figure 6(c) presents the event study estimated coefficients between organic adopters and covid adopters. As we can see, the estimated coefficients are not significantly different from zero, i.e., the parallel pre-trend assumption seems valid.  [^36] For the parallel trend assumption to hold, it is permissible for a small number of pre-treatment coefficients to be statistically significant (provided the majority of these coefficients are statistically insignificant), see Gopalan, Rosinger, and Ahn (2020) and Chapter 5 of Angrist and Pischke (2008). This suggests that the DID estimation of treatment effects was not falsely inflated by any trends before the treatment. As a further robustness check, we also provide bounded ATT estimates for the total spend result between organic adopters and covid adopters in Web Appendix §F.2.
F.1.2 Parallel Trend Tests for Covid Analysis - Profitability Analysis
Figure A7 presents the parallel trend plot for the profit. Figure 7(a) visualizes the trends. Figure 7(b) presents the pre-trend between adopters (both covid and organic adopters) and offline-only customers (i.e., Non-Adopters), which shows some violation of parallel pre-trend. We use the honest DID approach from Rambachan and Roth (2023) and provide bounded ATT estimates for profit in Web Appendix §F.2. The incremental profit due to online adoption is still significant and robust, even in the presence of a large violation of parallel trend.
Finally, Figure 7(c) confirms that the parallel trend assumption for total profit between covid adopters and organic adopters is satisfied since almost all pre-treatment coefficients are insignificant and not statistically different from 0. As a further robustness check, we also provide bounded ATT estimates for the profit result between organic adopters and covid adopters in Web Appendix §F.2.
Figure A7: Parallel Trend Assumption - Profit
(a) Visualization
(b) Adopters vs. Non-Adopters
(c) Covid Adopters vs. Organic Adopters
F.1.3 Additional Parallel Trend Plots for Covid Analysis – Other Metrics
In this section, we validate the parallel trend assumption for DID results in other metrics between organic adopters and covid adopters, presented in Table A25 and A26. Because both organic and covid adopters only purchase offline during the pre-adoption period, the pre-adoption values for total and offline metrics are exactly the same. So here we test for total metrics and same conclusions apply to offline metrics as well. In Figure A8, we present the parallel trend plots for the metrics shown in Table A25. Figure A9 shows the event study estimates test for pre-trend. We can see some significant pre-treatment coefficients for these metrics, but the magnitude of these coefficients are small.
Figure A8: Monthly Average Metrics – Covid, Organic adopters and Offline-only Customers (i.e., Non-Adopters)
(a) Quantity
(b) Orders
(c) Unique Items
(d) Unique Brands
(e) Unique Subcategories
(f) Unique Categories
Figure A9: Parallel Trend for Other Metrics - Covid Adopters vs. Organic Adopters
(a) Quantity
(b) Orders
(c) Unique Items
(d) Unique Brands
(e) Unique Subcategories
(f) Unique Categories
F.1.4 Parallel Trend for Black Friday Analysis
First, we provide a pre-trend test for two main metrics (i.e. spend and profit) between adopters (including both Black Friday adopters and Organic adopters) relative to offline-only customers. In total, we have 10 pre-treatment periods (i.e., between January 2019 and October 2019), labeled as -10 to -1, and 11 post-adoption periods (i.e., between December 2019 and October 2020), labeled as 1 to 11. The base period is Octorber 2019, and the event time 0 refers to the adoption month (i.e., November 2019). We exhibit a minor violation during the early stage of the pre-adoption period, but most pre-adoption coefficients are significant. To show the robustness of our main finding in the presence of violation of parallel pre-trend, we show the bounded ATT in Web Appendix §F.2 and confirm that the incremental spend and profit are still positive and significant in the presence of violation of parallel trend.
Figure A10: Event Study – Adopters (pooled by Organic and Black Friday) vs. Offline-only
(a) Spend
(b) Profit
Next, we provide a pre-trend test between Black Friday adopters and Organic adopters. Figure A11 visualizes the monthly average metric values for Black Friday adopters, organic adopters and offline-only customers. Figure A12 presents the event study regression estimates for the parallel pre-trend of the Black Friday study.
Figure A11: Monthly Average Metrics - Black Friday, Organic adopters, and Offline-only Customers (i.e., Non-Adopters)
(a) Spend
(b) Profit
(c) Quantity
(d) Orders
(e) Unique Items
(f) Unique Brands
(g) Unique Subcategories
(h) Unique Categories
Figure A12: Parallel Trend for Black Friday Analysis - Total Metrics (Black Friday Adopters vs. Organic Adopters). We present the estimates of event study regression. The pretreatment outcomes of all offline metrics are the same as the total metrics, since all customers only shop offline.
(a) Spend
(b) Profit
(c) Quantity
(d) Orders
(e) Unique Items
(f) Unique Brands
(g) Unique Subcategories
(h) Unique Categories
F.1.5 Parallel Trend for Loyalty Program Analysis
First, we provide a pre-trend test for two main metrics (i.e. spend and profit) between adopters (including both Loyalty Program adopters and Organic adopters) relative to offline-only customers (see Figure A13). We exhibit a violation of parallel trend assumption. To show the robustness of our main finding in the presence of violation of parallel pre-trend, we show the bounded ATT in Web Appendix §F.2 and confirm that the incremental spend is still positive and significant in the presence of violation of parallel trend.
Figure A13: Event Study – Adopters (pooled by Organic and Loyalty Program) vs. Offline-only
(a) Spend
(b) Profit
Next, we provide a pre-trend test between Loyalty Program adopters and Organic adopters. Figure A14 visualizes the monthly average metric values for loyalty program adopters, organic adopters and offline-only customers. Figure A15 present the event study regression estimates for the parallel pre-trend of the loyalty program online activation study.
Figure A14: Monthly Average Metrics - Loyalty Program, Organic adopters, and Offline-only Customers (i.e., Non-Adopters).
(a) Spend
(b) Profit
(c) Quantity
(d) Orders
(e) Unique Items
(f) Unique Brands
(g) Unique Subcategories
(h) Unique Categories
Figure A15: Parallel Trend for Loyalty Program Analysis - Total Metrics. We present the estimates of event study regression for Loyalty Program Adopters. The pretreatment outcomes of all offline metrics are the same as the total metrics, since all customers only shop offline.
(a) Spend
(b) Profit
(c) Quantity
(d) Orders
(e) Unique Items
(f) Unique Brands
(g) Unique Subcategories
(h) Unique Categories
F.2 Honest DID
In F.1, we see that when comparing adopters and offline-only customers, there are some violations of parallel trend assumption. There are also some concerns about the parallel trend assumption between covid adopters and organic adopters. To ensure that estimated ATTs of online adoption are robust to the violations in the parallel trends assumption, we perform the Honest Difference-in-Differences (Honest DID) estimation to quantify the extent to which violations in parallel trends can nullify the treatment effect (Rambachan and Roth 2023).
To conduct this sensitivity analysis, we impose the restriction that the change in slope of the differential trend between treatment and control customers across two periods is bounded by:
$\Delta^{SD}:=\{\zeta:|(\zeta_{t+1}-\zeta_{t})-(\zeta_{t}-\zeta_{t-1})|\leq M,~\forall t\}$ (A6)
Here, $\zeta_{t}$ is the difference in trends between two groups at time $t$, and $M$ is the magnitude of maximum allowable change in trend between consecutive periods.
First, we estimate a linear trend model, including all treatment and control customers using all customer-month observations during the pre-adoption periods:
$y_{it}=\gamma_{i}+\eta_{t}+\phi_{e,v}\mathbbm{I}\{Treatment\}_{i}\times t+\epsilon_{it}$ (A7)
Here, $\mathbbm{I}\{Treatment\}_{i}$ refers to the treatment customers. When we compare adopters with offline-only customers, $\mathbbm{I}\{Treatment\}_{i}=1$ is customer $i$ is an adopter. When we compare covid adopters with organic adopters, $\mathbbm{I}\{Treatment\}_{i}=1$ if customer $i$ is a covid adopter. $t$ is the year-month for the pre-adoption period, and $\phi_{e,v}$ is the linear trend coefficient for event $e$ variable $v$. We test a wide range of $M$ values, up to and beyond $\phi_{e,v}$, for spend and profit under each event, and assess to what extent our ATT will be nullified.
We now present the Honest DID estimates for three events.
F.2.1 COVID-19
Table A35: Linear Trend – Adopters (Organic and Covid) vs. Offline-only
DVs: Spend Profit
Model: (1) (2)
Variables
Adopter * t 4.170 (0.377) 1.128 (0.141)
[0.000] [0.000]
Fixed-effects
Customer Yes Yes
YearMonth Yes Yes
Fit statistics
Observations 5,146,359 5,146,359
R2 0.48244 0.39923
Within R2 $6.28\times 10^{-5}$ $3.05\times 10^{-5}$
Clustered (Customer) standard-errors are presented in standard brackets,
while p‑values are shown in square brackets.
Figure A16: Honest DID Estimation - Adopters (Organic and Covid) vs. Offline-only
(a) Total Spend
(b) Total Profit
We first present the Honest DID estimates with varying $M$ values for two main metrics, monthly spend and profit between adopters (i.e., including both organic and covid adopters) and offline-only customers. Figure 16(a) presents the honest DID 95% confidence intervals when we allow for a violation of parallel trends between consecutive periods up to $M$. We test a wide range of $M$ values and conclude that even if the post-treatment violation of parallel trend is twice as large as the pre-treatment trend (i.e., estimated coefficient from the linear trend model in Table A35), our ATT estimates and main effects still hold: adopters spend more and are more profitable than offline-only customers during the post-adoption periods.
Table A36: Linear Trend – Covid Adopters vs. Organic Adopters
DVs Spend Profit
Model: (1) (2)
Variables
Covid_Adopter * t 4.393 (0.8605) 1.415 (0.3183)
[0.000] [0.000]
Fixed-effects
Customer Yes Yes
YearMonth Yes Yes
Fit statistics
Observations 378,387 378,387
R2 0.51305 0.42830
Within R2 0.00012 $7.71\times 10^{-5}$
Clustered (Customer) standard-errors are presented in standard brackets,
while p‑values are shown in square brackets.
Figure A17: Honest DID Estimation - Covid Adopters vs. Organic Adopters
(a) Total Spend
(b) Total Profit
Next, we present the Honest DID estimates between organic adopters and covid adopters. Figure A17 presents the honest DID 95% confidence intervals when we allow for a violation of parallel trends between consecutive periods up to $M$. We test a wide range of $M$ values and conclude that even if the post-treatment violation of parallel trend is twice as large as the pre-treatment trend (i.e., estimated from the linear trend model in Table A36), our ATT estimates and main effects still hold for spend: covid adopters have a similar incremental spend as the organic adopters post-online-adoption.
For the profit, we demonstrate that the main finding remains valid in the presence of a linear pre-trend (i.e., a significant 95% confidence interval when $M=0$ and a similar magnitude of confidence interval as the original effect). This also confirms the robustness of the result, indicating covid adopters are more profitable than organic adopters in the post-adoption period. However, the effect becomes weaker when we assume a more severe post-treatment violation of the parallel trend.
F.2.2 Black Friday
Table A37: Linear Trend – Adopters (Organic and Black Friday) vs. Offline-only
DVs Spend Profit
Model: (1) (2)
Variables
Adopter * t 4.244 (1.571) 1.081 (0.6015)
[0.007] [0.072]
Fixed-effects
Customer Yes Yes
YearMonth Yes Yes
Fit statistics
Observations 3,855,908 3,855,908
R2 0.48754 0.40597
Within R2 $5.24\times 10^{-6}$ $2.25\times 10^{-6}$
Clustered (Customer) standard-errors are presented in standard brackets,
while p‑values are shown in square brackets.
Figure A18: Honest DID Estimation - Adopters (Organic and Black Friday) vs. Offline-only
(a) Total Spend
(b) Total Profit
We now present the Honest DID estimates between Adopters (i.e., including both organic and Black Friday) and Offline-only customers. Figure 18(a) presents the honest DID 95% confidence intervals when we allow for a violation of parallel trends between consecutive periods up to $M$. We test a wide range of $M$ values and conclude that even if the post-treatment violation of parallel trend is twice as large as the pre-treatment trend (i.e., estimated from the linear trend model in Table A37), our ATT estimates and main effects for spend still hold: adopters spend more than offline-only customers during the post-adoption periods. For the profit, we do not see a significant linear pre-trend (see column 2 of Table A37). Also, we demonstrate that the main finding remains valid in the presence of a linear pre-trend and some deviation (i.e., a magnitude of 0.4) from the linear pre-trend (i.e., 1.081 from Table A37). This also confirms the robustness of the result, indicating adopters are more profitable than offline-only customers after the online adoption.
F.2.3 Loyalty Program
Table A38: Linear Trend – Adopters (Organic and Loyalty Program) vs. Offline-only
DVs: Spend Margin
Model: (1) (2)
Variables
Adopter * t 17.51 (1.504) 6.811 (0.578)
[0.000] [0.000]
Fixed-effects
Customer Yes Yes
YearMonth Yes Yes
Fit statistics
Observations 2,073,524 2,073,524
R2 0.55397 0.51055
Within R2 0.00012 0.00012
Clustered (Customer) standard-errors are presented in standard brackets,
while p‑values are shown in square brackets.
Figure A19: Honest DID Estimation – Adopters (Organic and Loyalty Program) vs. Offline-only
(a) Total Spend
(b) Total Profit
We present the Honest DID estimates between adopters (including organic and loyalty program adopters) and offline-only customers. Figure 19(a) presents the honest DID 95% confidence intervals when we allow for a violation of parallel trends between consecutive periods up to $M$. We test a wide range of $M$ values and conclude that even if the post-treatment violation of parallel trend is twice as large as the pre-treatment trend (i.e., estimated from the linear trend model in Table A38), our ATT estimates and main effects still hold: adopters spend more than offline-only customers after the online adoption. However, under this robustness check, the incremental profit associated with adopters (pooled by organic adopters and loyalty program adopters) is not significantly greater than that of offline-only customers during the post-adoption period, even when we assume a more severe post-treatment violation of the parallel trend.
F.3 Appendix to DID Models with Propensity Score Matching and Inverse Propensity Score Weighting
In this section, we provide a set of robustness check under PSM and IPTW. To conduct this analysis, we label the event adopters as 1 and organic adopters as 0 and model the propensity score of customer $i$ being a COVID adopter. The propensity score model is specified as follows:
$PropensityScore_{i}=f(X_{i}^{T}\theta),$ (A8)
[…]
Next, we use these propensity scores to augment our DID anaysis in two ways.
  • $\bullet$
    First, we perform a customer-level propensity score matching and then re-run the DID analysis. Matching based on propensity score is much more efficient than matching based on a set of covariates (Rosenbaum and Rubin 1985). Based on the predicted propensity score of each customer $i$, we use MatchIt package in R (Stuart et al. 2011) to perform nearest-neighbor matching for each event adopter. We then run the DID analysis on the newly matched sample (see Table A41, A42 and A43 in Web Appendix F.3.2). We find that the results from this exercise are consistent with the main findings.
  • $\bullet$
    Next, we use a slightly different approach to control for any potential selection issues. Specifically, we use an inverse propensity of treatment weight-adjusted (IPTW) DID regression. Here, the sample is the same as the one from the main analysis, but each observation is inversely weighted by its propensity score. We then perform a two-sample t-test on all the pre-treatment variables based on the weighted sample and confirm that the two samples are statistically insignificant on all the variables (see Table A40 for the weighted sample of organic adopters and event adopters, respectively in the Web Appendix F.3.3). The DID results of event adopters and organic adopters from this analysis are shown in Table A44, A45 and A46 for covid, Black Friday, and Loyalty Program analysis, respectively in Web Appendix F.3.3. Again, we find that the results and conclusions are quite similar to those in the main section.
F.3.1 Propensity Score Model Estimation
Table A39: Propensity Score Logit Model – Event Adopters vs. Organic Adopters
DVs:
Covid_Adopter BlackFriday_Adopter LoyaltyProgram_Adopter
Gender_NoInformation -0.044 (0.027) 0.095 (0.088) -0.031 (0.042)
[0.103] [0.278] [0.459]
Gender_Female 0.052 (0.018) 0.098 (0.062) -0.050 (0.035)
[0.004] [0.114] [0.152]
LogTenureDays 0.243 (0.009) 0.175 (0.031) -0.087 (0.026)
[0.000] [0.000] [0.001]
AvgSpendPerMonth -0.0001 (0.0001) 0.0003 (0.0002) 0.002 (0.0001)
[0.317] [0.133] [0.000]
AvgQuantityPerMonth 0.010 (0.002) 0.008 (0.006) 0.0001 (0.003)
[0.000] [0.183] [0.974]
AvgUniqueOrdersPerMonth -0.147 (0.018) -0.121 (0.063) 0.174 (0.028)
[0.000] [0.055] [0.000]
AvgUniqueItemsPerMonth 0.036 (0.013) 0.022 (0.043) -0.066 (0.021)
[0.006] [0.610] [0.002]
AvgUniqueBrandsPerMonth -0.086 (0.022) -0.031 (0.070) -0.0001 (0.042)
[0.000] [0.659] [0.998]
AvgUniqueSubcategoriesPerMonth 0.003 (0.025) 0.081 (0.080) 0.315 (0.055)
[0.905] [0.311] [0.000]
AvgUniqueCategoriesPerMonth 0.088 (0.026) -0.040 (0.089) -0.096 (0.059)
[0.001] [0.655] [0.103]
AvgUniqueVisitedStoresPerMonth 0.289 (0.041) 0.284 (0.137) -0.074 (0.067)
[0.000] [0.038] [0.272]
AvgMarginPerMonth 0.0002 (0.0001) -0.001 (0.0005) -0.006 (0.0003)
[0.045] [0.045] [0.000]
Constant -1.063 (0.410) -1.191 (0.203) -0.079 (0.189)
[0.010] [0.000] [0.672]
Age Yes Yes Yes
Household Income Yes Yes Yes
Observations 46,045 4,441 18,313
Log Likelihood -48,790.450 -3,983.461 -19,285.020
Akaike Inf. Crit. 97,622.890 8,006.922 38,610.040
Standard‐errors in standard brackets; p‑values in square brackets.
  • Note: AvgUniqueVisitedStoresPerMonth measures the average number of distinct stores a customer visits per month during the pre-adoption period under each study.
Table A39 presents the propensity score model estimates for covid, black friday, and loyalty program adopters versus organic adopters. Figure A20 presents the distribution of propensity scores of event adopters and organic adopters under three studies. After performing nearest-neighbor matching, A21 shows that the distribution of propensity scores between event adopters and organic adopters become more overlapped.
Figure A20: Propensity Score Model Predictions - Before Matching
(a) Covid
(b) Black Friday
(c) Loyalty Program
Figure A21: Propensity Score Model Predictions - After Matching
(a) Covid
(b) Black Friday
(c) Loyalty Program
Table A40: Customer-level Summary Statistics of Pre-Adoption Variables for Event Adopters and Organic Adopters: Re-weighted Sample Based on Inverse Propensity Score. Columns 2 and 3 show the weighted mean of variables of event adopters and organic adopters, with weights being the inverse of propensity.
Variable Event Adopters mean Organic Adopters mean Diff T-stats P-value
Panel A: Covid Analysis
LogTenure(Days) 5.12 5.12 -0.00 -0.40 0.69
AvgSpendPerMonth 459.22 461.68 -2.47 -0.42 0.67
AvgQuantitiesPerMonth 4.72 4.91 -0.19 -1.67 0.09
AvgOrdersPerMonth 0.95 0.95 0.00 0.04 0.97
AvgUniqueItemsPerMonth 2.90 2.90 0.00 0.01 1.00
AvgUniqueBrandsPerMonth 2.32 2.33 -0.00 -0.04 0.97
AvgUniqueSubcategoriesPerMonth 2.03 2.03 -0.00 -0.10 0.92
AvgUniqueCategoriesPerMonth 1.65 1.65 -0.00 -0.09 0.93
AvgUniqueVisitedStoresPerMonth 0.65 0.65 0.00 0.04 0.97
AvgMarginPerMonth 162.25 162.72 -0.47 -0.22 0.83
Panel B: Black Friday Analysis
LogTenure(Days) 5.01 5.01 0.00 0.02 0.99
AvgSpendPerMonth 477.58 477.41 0.17 0.01 0.99
AvgQuantitiesPerMonth 4.75 4.70 0.05 0.21 0.83
AvgOrdersPerMonth 1.00 1.00 0.00 0.01 1.00
AvgUniqueItemsPerMonth 3.00 3.00 0.00 0.01 0.99
AvgUniqueBrandsPerMonth 2.42 2.42 -0.00 -0.00 1.00
AvgUniqueSubcategoriesPerMonth 2.10 2.10 -0.00 -0.01 0.99
AvgUniqueCategoriesPerMonth 1.70 1.70 -0.00 -0.00 1.00
AvgUniqueVisitedStoresPerMonth 0.68 0.68 -0.00 -0.01 0.99
AvgMarginPerMonth 166.97 167.08 -0.11 -0.02 0.99
Panel C: Loyalty Program Analysis
LogTenure(Days) 6.45 6.45 0.00 0.14 0.89
AvgSpendPerMonth 514.74 535.10 -20.36 -1.38 0.17
AvgQuantitiesPerMonth 3.73 3.85 -0.12 -0.70 0.48
AvgOrdersPerMonth 0.93 0.94 -0.01 -0.42 0.67
AvgUniqueItemsPerMonth 2.28 2.29 -0.02 -0.27 0.78
AvgUniqueBrandsPerMonth 1.80 1.81 -0.01 -0.37 0.71
AvgUniqueSubcategoriesPerMonth 1.54 1.55 -0.01 -0.39 0.70
AvgUniqueCategoriesPerMonth 1.30 1.31 -0.01 -0.38 0.70
AvgUniqueVisitedStoresPerMonth 0.61 0.61 -0.00 -0.44 0.66
AvgMarginPerMonth 180.94 184.54 -3.60 -0.70 0.49
F.3.2 DID Results with PSM
Table A41, A42 and A43 present the DID estimates of Equation (3) of key purchase metrics based on the matched sample under covid, Black Friday, and loyalty program analysis. All results are consistent with the main findings.
Table A41: Covid Adopters vs. Organic Adopters – PSM + DID
DVs: Spend Offline Spend Share of Offline Spend Profit
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 438.1 (5.946) 438.1 (5.946) 1.000 (1.94e-15) 147.4 (1.901)
[0.000] [0.000] [0.000] [0.000]
Covid_Adopter 27.61 (6.816) 27.61 (6.816) 0.000 (0.000) 9.298 (2.181)
[0.000] [0.000] [0.000] [0.000]
Post 84.05 (5.916) -202.1 (4.983) -0.5091 (0.0038) 20.28 (1.994)
[0.000] [0.000] [0.000] [0.000]
Covid_Adopter * Post -8.837 (6.768) -1.842 (6.910) 34.42 (5.729) 38.08 (5.813) 0.0793 (0.0044) 0.0693 (0.0046) 7.237 (2.294) 9.258 (2.423)
[0.192] [0.790] [0.000] [0.000] [0.000] [0.000] [0.002] [0.000]
Fixed‐effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 1,689,232 1,689,232 1,689,232 1,689,232 901,177 901,177 1,689,232 1,689,232
R2 0.00125 0.39081 0.01127 0.36431 0.17947 0.50579 0.00122 0.29503
Within R2 1.72e-07 0.00012 0.00125 2.59e-05
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A42: Black Friday Adopters vs. Organic Adopters – PSM + DID
DVs: Spend Offline Spend Share of Offline Spend Profit
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 455.4 (17.04) 455.4 (17.04) 1.000 (8.42e-17) 149.7 (5.392)
[0.000] [0.000] [0.000] [0.000]
BlackFriday_Adopter 35.72 (24.69) 35.72 (24.69) 0.000 (0.000) 12.36 (7.925)
[0.148] [0.148] [0.000] [0.124]
Post 11.79 (14.44) -201.8 (13.40) -0.4275 (0.0127) 7.025 (4.835)
[0.414] [0.000] [0.000] [0.147]
BlackFriday_Adopter * Post -83.50 (18.99) -78.26 (17.94) -27.88 (18.31) -16.68 (17.37) 0.0921 (0.0162) 0.0959 (0.0159) -24.23 (6.380) -22.46 (6.431)
[0.000] [0.000] [0.128] [0.337] [0.000] [0.000] [0.000] [0.001]
Fixed‐effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 42,329 42,329 42,329 42,329 23,221 23,221 42,329 42,329
R2 0.00137 0.41012 0.02553 0.41414 0.23976 0.52090 0.00065 0.34764
Within R2 0.00099 5.74e-05 0.00671 0.00058
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A43: Loyalty Program Adopters vs. Organic Adopters – PSM + DID
DVs: Spend Offline Spend Share of Offline Spend Profit
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 464.3 (5.119) 464.3 (5.119) 1.000 (2.44e-18) 170.0 (1.864)
[0.000] [0.000] [0.000] [0.000]
Loyalty_Adopter 112.1 (19.25) 112.1 (19.25) 0.000 (0.000) 28.59 (6.478)
[0.000] [0.000] [0.000] [0.000]
Post -41.09 (5.787) -245.4 (5.191) -0.4546 (0.0044) -24.27 (2.034)
[0.000] [0.000] [0.000] [0.000]
Loyalty_Adopter * Post -244.3 (22.01) -52.62 (20.60) -198.8 (19.86) -8.541 (19.41) -0.0009 (0.0119) 0.0013 (0.0117) -83.07 (7.797) -16.32 (7.228)
[0.000] [0.011] [0.000] [0.660] [0.940] [0.912] [0.000] [0.024]
Loyalty program controls Yes Yes Yes Yes Yes Yes Yes Yes
Fixed‐effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 159,333 159,333 159,333 159,333 90,778 90,778 159,333 159,333
R2 0.05562 0.47539 0.05221 0.44214 0.28560 0.56883 0.04528 0.44853
Within R2 0.02722 0.01915 0.00922 0.02278
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
F.3.3 DID Results with IPTW
Table A44, A45 and A46 present the DID estimates of Equation (3) of key purchase metrics with inverse propensity score weighting under covid, Black Friday, and loyalty program analysis. All results are consistent with the main findings.
Table A44: Covid Adopters vs. Organic Adopters - IPTW + DID
DVs Spend Offline Spend Share of Offline Spend Profit
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 436.7 (5.445) 436.7 (5.445) 1.000 ($1.68\times 10^{-15}$) 146.7 (1.731)
[0.000] [0.000] [0.000] [0.000]
Covid_Adopter 15.72 (6.310) 15.72 (6.310) 0.000 (0.000) 5.898 (2.016)
[0.014] [0.014] [0.000] [0.004]
Post 89.00 (5.416) -198.3 (4.544) -0.5071 (0.0035) 22.29 (1.833)
[0.000] [0.000] [0.000] [0.000]
Covid_Adopter * Post -13.12 (6.310) -6.305 (6.550) 34.22 (5.316) 37.96 (5.465) 0.0731 (0.0041) 0.0651 (0.0043) 4.667 (2.144) 6.651 (2.307)
[0.040] [0.315] [0.000] [0.000] [0.000] [0.000] [0.030] [0.008]
Fixed-effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 1,783,949 1,783,949 949,448 949,448 1,783,949 1,783,949
R2 0.00126 0.38696 0.01100 0.36110 0.18052 0.50768 0.00105 0.29041
Within R2 $2.31\times 10^{-6}$ 0.00014 0.00125 $1.53\times 10^{-5}$
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A45: Black Friday Adopters vs. Organic Adopters - IPTW + DID
DVs: Spend Offline Spend Share of Offline Spend Profit
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 471.5 (11.07) 471.5 (11.07) 1.000 ($1.37\times 10^{-15}$) 156.7 (3.548)
[0.000] [0.000] [0.000] [0.000]
BlackFriday_Adopter 5.388 (19.85) 5.388 (19.85) 0.000 (0.000) 3.074 (6.679)
[0.784] [0.784] [0.000] [0.646]
Post 5.951 (9.359) -210.2 (8.970) -0.4277 (0.0074) 4.135 (3.139)
[0.532] [0.000] [0.000] [0.186]
BlackFriday_Adopter * Post -74.02 (15.36) -68.89 (14.30) -16.17 (15.14) -4.429 (14.12) 0.0845 (0.0127) 0.0848 (0.0122) -22.85 (5.303) -20.64 (5.231)
[0.000] [0.000] [0.288] [0.765] [0.000] [0.000] [0.000] [0.000]
Fixed-effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 76,842 76,842 76,842 76,842 42,299 42,299 76,842 76,842
R2 0.00123 0.40586 0.02482 0.39667 0.25280 0.53426 0.00074 0.34948
Within R2 0.00067 $3.58\times 10^{-6}$ 0.00455 0.00042
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A46: Loyalty Program Adopters vs. Organic Adopters - IPTW + DID
DVs Spend Offline Spend Share of Offline Spend Margin
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 445.6 (4.422) 445.6 (4.422) 1.000 ($1.64\times 10^{-14}$) 164.7 (1.627)
[0.000] [0.000] [0.000] [0.000]
Loyalty_Adopter 130.8 (19.08) 130.8 (19.08) 0.000 (0.000) 33.90 (6.413)
[0.000] [0.000] [0.000] [0.000]
Post -30.58 (5.047) -230.5 (4.497) -0.4530 (0.0039) -21.23 (1.788)
[0.000] [0.000] [0.000] [0.000]
Loyalty_Adopter * Post -253.8 (20.82) -52.53 (19.43) -210.1 (19.03) -11.43 (18.60) $9.51\times 10^{-5}$ (0.0115) 0.0017 (0.0114) -86.53 (7.352) -15.22 (6.835)
[0.000] [0.009] [0.000] [0.546] [0.994] [0.888] [0.000] [0.024]
Loyalty program controls Yes Yes Yes Yes Yes Yes Yes Yes
Fixed-effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 197,658 197,658 197,658 197,658 110,339 110,339 197,658 197,658
R2 0.05633 0.47155 0.05137 0.43591 0.28420 0.57046 0.04567 0.44366
Within R2 0.02694 0.01910 0.00903 0.02249
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
F.3.4 Doubly Robust DID for Covid Analysis
Given the potential concern of parallel trend assumption under COVID-19 analysis, we seek an alternative identification assumption and conduct a doubly robust DID for the covid analysis. We now show that the main conclusions remain consistent by using the doubly robust DID (Sant’Anna and Zhao 2020). We provide the ATT estimates for each metric based on the package did in R provided by Callaway and Sant’Anna (2021). Table A47 presents the doubly robust DID estimates for all metrics. Figure A22 and A23 below present the event study figure under doubly robust DID results for main metrics (i.e., spend and profit) and other metrics, respectively. We can see that the parallel pre-trends are also well satisfied, conditional on customer demographics and pre-adoption purchase variables. The estimation results are consistent with the main findings.
Table A47: Doubly Robust DID – Covid Adopters vs. Organic Adopters
DVs ATT SE T-stat P-value CI- CI+
Spend -5.963 7.814 -0.763 0.445 -21.279 9.353
Offline Spend 26.893 8.339 3.225 0.001 10.549 43.238
Share of Offline Spend 0.083 0.004 20.550 0.000 0.075 0.091
Profit 11.623 2.942 3.951 0.000 5.856 17.390
Quantity 0.246 0.324 0.760 0.447 -0.389 0.881
Orders -0.034 0.013 -2.579 0.010 -0.059 -0.008
Unique Items 0.079 0.042 1.875 0.061 -0.004 0.162
Unique Brands 0.054 0.028 1.907 0.057 -0.002 0.110
Unique Subcategories 0.047 0.022 2.141 0.032 0.004 0.091
Unique Categories 0.028 0.017 1.719 0.086 -0.004 0.061
Offline Spend 26.893 8.339 3.225 0.001 10.549 43.238
Offline Quantity 0.367 0.234 1.570 0.116 -0.091 0.824
Offline Orders 0.039 0.010 3.783 0.000 0.019 0.060
Offline Unique Items 0.178 0.037 4.778 0.000 0.105 0.251
Offline Unique Brands 0.134 0.026 5.108 0.000 0.082 0.185
Offline Unique Subcategories 0.106 0.021 5.051 0.000 0.065 0.147
Offline Unique Categories 0.074 0.016 4.617 0.000 0.042 0.105
Figure A22: Doubly Robust Event Study for the Main Metrics – Covid Adopters vs. Organic Adopters
(a) Spend
(b) Profit
Figure A23: Doubly Robust Event Study for the Other Total Metrics – Covid Adopters vs. Organic Adopters. The pretreatment outcomes of all offline metrics are the same as the total metrics, since all customers only shop offline.
(a) Quantity
(b) Orders
(c) Unique Items
(d) Unique Brands
(e) Unique Subcategories
(f) Unique Categories
F.4 Appendix to Alternate definitions for the organic adopter and covid adopter cohorts
F.4.1 Appendix to Covid Analysis - Re-defining the Organic Adopter Cohort
In this section, we provide a robustness check by redefining organic adopters as those customers who were offline-only before 2019-10-31 and made their first online purchase between 2019-11-01 to 2020-02-29. The definition of covid adopters remains the same. Under this new definition, we continue to have 33,246 covid adopters and a larger number of 25,958 organic adopters. As before, we exclude the periods used for defining these two cohorts (2019-11-01 to 2020-04-30) and re-run the same set of DID models as in the main analysis. Table A48 presents the DID results between covid adopters and organic adopters. Table A49 and A50 presents DID results for other total and offline metrics respectively. We find that all the results are consistent with the main analysis.
Table A48: Covid Adopters vs. Organic Adopters – Redefining Organic Adopter Group
DVs: Spend Offline Spend Share of Offline Spend Profit
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 450.4 (3.916) 450.4 (3.916) 1.000 (3.24e-16) 151.2 (1.288)
[0.000] [0.000] [0.000] [0.000]
Covid_Adopter 27.56 (5.291) 27.56 (5.291) 0.000 (0.000) 9.284 (1.727)
[0.000] [0.000] [0.000] [0.000]
Post 67.35 (4.220) -214.8 (3.423) -0.4980 (0.0024) 14.89 (1.397)
[0.000] [0.000] [0.000] [0.000]
Covid_Adopter * Post -4.580 (5.459) 10.03 (5.944) 34.76 (4.587) 44.76 (4.773) 0.0682 (0.0032) 0.0580 (0.0034) 8.835 (1.839) 13.20 (2.082)
[0.402] [0.091] [0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Fixed-effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 2,178,690 2,178,690 2,178,690 2,178,690 1,153,061 1,153,061 2,178,690 2,178,690
R2 0.00079 0.39039 0.01289 0.36518 0.16191 0.50935 0.00099 0.30924
Within R2 5.66e-06 0.00021 0.00108 6.31e-05
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A49: Covid Adopters vs. Organic Adopters (Other Total Metrics) – Redefining Organic Adopter Group
DVs: Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Covid_Adopter * Post 0.211 (0.090) -0.007 (0.008) 0.090 (0.024) 0.077 (0.017) 0.073 (0.013) 0.059 (0.010)
[0.019] [0.382] [0.000] [0.000] [0.000] [0.000]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 2,178,690 2,178,690 2,178,690 2,178,690 2,178,690 2,178,690
R2 0.46518 0.37879 0.37581 0.34501 0.32459 0.31320
Within R2 1.37e-05 1.45e-06 2.86e-05 3.97e-05 5.59e-05 6.20e-05
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A50: Covid Adopters vs. Organic Adopters (Other Offline Metrics) – Redefining Organic Adopter Group
DVs: Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Covid_Adopter * Post 0.113 (0.076) 0.036 (0.007) 0.115 (0.024) 0.106 (0.017) 0.091 (0.013) 0.073 (0.010)
[0.137] [0.000] [0.000] [0.000] [0.000] [0.000]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 2,178,690 2,178,690 2,178,690 2,178,690 2,178,690 2,178,690
R2 0.44370 0.36979 0.35206 0.33015 0.31714 0.31261
Within R2 7.65e-06 6.34e-05 5.52e-05 8.80e-05 9.95e-05 1.10e-04
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
F.4.2 Appendix to Covid Analysis - Re-defining the Covid Adopter Cohort
In this section, we provide an alternative definition of covid adopters cohort for the covid analysis. We provide a robustness check by redefining covid adopters as those customers who were offline-only before 2019-12-31 and made their first online purchase between 2020-05-01 and 2020-06-30 (i.e., two months later than the exact onset of COVID-19). The definition of organic adopters remains the same as the definition in §5.1.1. Under this definition, we continue to have 12,799 organic adopters but now have 27,448 covid adopters. As before, we exclude the periods used for defining these two cohorts (i.e., 2020-01-01 to 2020-06-30), re-run the same set of DID models as in the main analysis, and present the results in Table A51. Table A52 and A53 present DID results for other total and offline metrics. We find that all results are consistent with the main analysis.
Table A51: Covid Adopters vs. Organic Adopters – Re‐defining Covid Adopters Group
DVs: Spend Offline Spend Share of Offline Spend Profit
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 432.8 (2.87) 432.8 (2.17) 1.00 (0.002) 145.7 (1.10)
[0.000] [0.000] [0.000] [0.000]
Covid_Adopter -30.9 (3.42) -30.9 (2.59) 0.000 (0.000) -9.07 (1.31)
[0.000] [0.000] [0.000] [0.000]
Post 88.8 (3.23) -199.3 (2.45) -0.510 (0.002) 20.5 (1.23)
[0.000] [0.000] [0.000] [0.000]
Covid_Adopter * Post 3.90 (3.86) 7.16 (6.69) 70.1 (2.93) 70.2 (5.50) 0.086 (0.002) 0.078 (0.004) 9.26 (1.48) 9.36 (2.31)
[0.323] [0.286] [0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Fixed‐effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 1,461,440 1,461,440 1,461,440 1,461,440 744,927 744,927 1,461,440 1,461,440
R2 0.00205 0.37338 0.00919 0.36248 0.18744 0.51561 0.00112 0.27316
Within R2 3.38e-06 0.00055 0.00202 3.41e-05
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A52: Covid Adopters vs. Organic Adopters – Other Total Metrics (Re‐defining Covid Adopters Group)
DVs: Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Covid_Adopter * Post 0.085 (0.103) 0.001 (0.011) 0.154 (0.029) 0.125 (0.021) 0.111 (0.016) 0.083 (0.013)
[0.411] [0.927] [0.000] [0.000] [0.000] [0.000]
Fixed‐effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 1,461,440 1,461,440 1,461,440 1,461,440 1,461,440 1,461,440
R2 0.46377 0.36292 0.35646 0.32818 0.30849 0.29795
Within R2 3.11e-06 5.72e-08 9.05e-05 0.00011 0.00013 0.00012
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A53: Covid Adopters vs. Organic Adopters – Other Offline Metrics (Re‐defining Covid Adopters Group)
DVs: Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Covid_Adopter * Post 0.473 (0.086) 0.092 (0.009) 0.299 (0.028) 0.240 (0.020) 0.201 (0.016) 0.154 (0.012)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Fixed‐effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 1,461,440 1,461,440 1,461,440 1,461,440 1,461,440 1,461,440
R2 0.44248 0.35480 0.33845 0.31801 0.30476 0.30009
Within R2 0.00016 0.00044 0.00039 0.00047 0.00050 0.00049
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
F.4.3 Appendix to Covid Analysis - Comparing Covid Adopters (March/April 2020) vs. Later Covid Adopters (May/June 2020)
Table A54 compares covid adopters (i.e., adopt during March/April 2020) and later covid adopters (i.e., adopt during May/June 2020). Although the later covid adopters spend more than the covid adopters by 9.96 MCU, it is not large in magnitude and with a p-value of 0.048. So we conclude there is no major difference in post-adoption spend, share of offline spend (i.e., channel utilization), and profit.
Table A54: Covid Adopters (March/April) vs. Later Covid Adopters (May/June)
DVs: Spend Offline Spend Share of Offline Spend Profit
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 465.7 (1.690) 465.7 (1.340) 1.00 (0.0010) 156.7 (0.594)
[0.000] [0.000] [0.000] [0.000]
Later_Covid_Adopter -63.7 (2.530) -63.7 (2.010) 0.000 (0.000) -20.0 (0.890)
[0.000] [0.000] [0.000] [0.000]
Post 78.3 (1.930) -164.0 (1.530) -0.423 (0.001) 27.9 (0.676)
[0.000] [0.000] [0.000] [0.000]
Later_Covid_Adopter * Post 14.4 (2.880) 9.96 (5.040) 34.9 (2.290) 30.4 (4.270) -0.001 (0.002) -0.0002 (0.003) 1.82 (1.010) 0.070 (1.770)
[0.000] [0.048] [0.000] [0.000] [0.617] [0.947] [0.072] [0.968]
Fixed-effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 2,240,950 2,240,950 2,240,950 2,240,950 1,170,421 1,170,421 2,240,950 2,240,950
R2 0.00241 0.39921 0.00832 0.36673 0.16862 0.50555 0.00231 0.35185
Within R2 8.12e-06 0.00011 9.28e-09 3.00e-09
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
F.5 Appendix to Using Natural Log Scale Dependent Variables
In this section, we provide a robustness check for using natural log transformed scale of the dependent variables (i.e., ln(dependent variable + 1)), and conclude that the direction of effects are consistent with the main finding. Note that the profit variable can take both positive and negative values, so we do not consider natural log transformation of profit. First, we present the result for spend between adopters and offline-only customers in Table A55, and show that the results are consistent with Table 6.
Table A55: DID Analysis - Spend (in Natural Log Scale) (Adopters vs. Offline-only)
Study: Covid Black Friday Loyalty Program
DV ln(Spend + 1)
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 2.824 (0.0026) 2.946 (0.0027) 2.898 (0.0029)
[0.000] [0.000] [0.000]
Adopter 0.7498 (0.0104) 0.7398 (0.0320) 0.4664 (0.0146)
[0.000] [0.000] [0.000]
Post -0.9089 (0.0025) -1.048 (0.0023) -0.4705 (0.0030)
[0.000] [0.000] [0.000]
Adopter * Post 0.7266 (0.0108) 0.6883 (0.0104) 0.6734 (0.0316) 0.5742 (0.0302) 0.3381 (0.0197) 0.6099 (0.0190)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 22,204,608 22,204,608 9,505,999 9,505,999 4,403,449 4,403,449
R2 0.02929 0.30744 0.03104 0.36627 0.01150 0.29925
Within R2 0.00087 0.00010 0.00208
Clustered (Customer) standard-errors are presented in standard brackets, while p-values are shown in square brackets.
Next, we present DID results for all metrics of comparing event adopters and organic adopters by three events.
Covid Analysis
Table A56: DID Analysis of Total Metrics (Natural Log Scale) – Covid Adopters vs. Organic Adopters
DVs: Spend Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6) (7)
Variables
Covid_Adopter * Post 0.008 (0.023) 0.010 (0.008) -0.007 (0.004) 0.010 (0.006) 0.010 (0.006) 0.011 (0.005) 0.010 (0.005)
[0.728] [0.211] [0.080] [0.096] [0.096] [0.028] [0.046]
Fixed‐effects
Customer Yes Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949
R2 0.28677 0.38486 0.34049 0.33495 0.32103 0.31157 0.30123
Within R2 $2.16\times 10^{-7}$ $4.09\times 10^{-6}$ $6.85\times 10^{-6}$ $5.62\times 10^{-6}$ $7.22\times 10^{-6}$ $9.86\times 10^{-6}$ $9.94\times 10^{-6}$
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A57: DID Analysis of Offline Metrics (Natural Log Scale) – Covid Adopters vs. Organic Adopters
DVs: Spend Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6) (7)
Variables
Covid_Adopter * Post 0.148 (0.021) 0.025 (0.008) 0.017 (0.004) 0.029 (0.006) 0.029 (0.005) 0.027 (0.005) 0.025 (0.005)
[0.000] [0.002] [0.000] [0.000] [0.000] [0.000] [0.000]
Fixed‐effects
Customer Yes Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949
R2 0.30476 0.36449 0.34627 0.33410 0.32530 0.31930 0.31278
Within R2 $9.57\times 10^{-5}$ $2.79\times 10^{-5}$ $5.60\times 10^{-5}$ $5.29\times 10^{-5}$ $6.33\times 10^{-5}$ $6.46\times 10^{-5}$ $6.52\times 10^{-5}$
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A56 and A57 provide DID results between covid adopters and organic adopters with natural log scale dependent variables for total and offline metrics, respectively.
Black Friday Analysis
Table A58 and A59 provide DID results between Black Friday and organic adopters with natural log scale dependent variables for total and offline metrics, respectively.
Table A58: DID Analysis of Total Metrics (Natural Log Scale) – Black Friday Adopters vs. Organic Adopters
DVs: Spend Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6) (7)
Variables
BlackFriday_Adopter * Post -0.268 (0.062) -0.084 (0.021) -0.050 (0.011) -0.060 (0.017) -0.050 (0.015) -0.046 (0.014) -0.043 (0.013)
[0.000] [0.000] [0.000] [0.001] [0.001] [0.001] [0.001]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 76,842 76,842 76,842 76,842 76,842 76,842 76,842
R2 0.32002 0.41568 0.38717 0.38110 0.36995 0.35915 0.34685
Within R2 0.00049 0.00050 0.00069 0.00036 0.00030 0.00029 0.00031
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A59: DID Analysis of Offline Metrics (Natural Log Scale) – Black Friday Adopters vs. Organic Adopters
DVs: Spend Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6) (7)
Variables
BlackFriday_Adopter * Post 0.082 (0.059) 0.003 (0.020) 0.005 (0.010) 0.009 (0.017) 0.013 (0.015) 0.014 (0.014) 0.013 (0.013)
[0.170] [0.882] [0.639] [0.582] [0.383] [0.308] [0.325]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 76,842 76,842 76,842 76,842 76,842 76,842 76,842
R2 0.35766 0.42251 0.41348 0.39672 0.38879 0.38097 0.37182
Within R2 $5.17\times 10^{-5}$ $5.66\times 10^{-7}$ $6.82\times 10^{-6}$ $9.35\times 10^{-6}$ $2.27\times 10^{-5}$ $2.87\times 10^{-5}$ $3.05\times 10^{-5}$
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Loyalty Program Analysis
Table A60 and A61 provide DID results between loyalty program and organic adopters with natural log scale dependent variables for total and offline metrics, respectively.
Table A60: DID Analysis of Total Metrics (Natural Log Scale) – Loyalty Program Adopters vs. Organic Adopters
DVs: Spend Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6) (7)
Variables
Loyalty_Adopter * Post -0.151 (0.064) -0.029 (0.021) -0.022 (0.012) -0.025 (0.017) -0.025 (0.015) -0.021 (0.014) -0.017 (0.012)
[0.017] [0.158] [0.077] [0.144] [0.101] [0.146] [0.173]
Loyalty program controls Yes Yes Yes Yes Yes Yes Yes
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 197,658 197,658 197,658 197,658 197,658 197,658 197,658
R2 0.31155 0.42241 0.39069 0.37743 0.35922 0.34595 0.33365
Within R2 0.02144 0.02661 0.02799 0.02767 0.02723 0.02660 0.02525
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A61: DID Analysis of Offline Metrics (Natural Log Scale) – Loyalty Program Adopters vs. Organic Adopters
DVs: Spend Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6) (7)
Variables
Loyalty_Adopter * Post -0.029 (0.065) 0.011 (0.021) 0.005 (0.011) 0.003 (0.017) 0.002 (0.015) 0.003 (0.014) 0.005 (0.012)
[0.662] [0.606] [0.665] [0.866] [0.890] [0.826] [0.696]
Loyalty program controls Yes Yes Yes Yes Yes Yes Yes
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 197,658 197,658 197,658 197,658 197,658 197,658 197,658
R2 0.32577 0.41498 0.38925 0.38051 0.36537 0.35334 0.34367
Within R2 0.02343 0.02335 0.02347 0.02427 0.02436 0.02433 0.02379
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
F.6 Appendix to Handling Zero-Inflated Dependent Variables
In this section, we outline the procedure and results of the two-part hurdle model and the estimation results. The two-part hurdle model includes two separate estimation steps:
  • $\bullet$
    Step 1: Estimate a linear probability model with outcome variable $\mathbbm{I}\{Spend_{it}\}$ with all independent variables specified in Equation (1) and (2) when comparing adopters vs. offline-only, and independent variables specified in Equation (3) and (4) when comparing event adopters vs. organic adopters;
  • $\bullet$
    Step 2: Estimate Equation (1) and (2) conditional on customers making non-zero spend when comparing adopters vs. offline-only, and independent variables specified in Equation (3) and (4) when comparing event adopters vs. organic adopters.
Next, we present the estimation results for adopters vs. offline-only customers and for each event Adopters vs. organic adopters.
Adopters vs. Offline-only
Table A62 presents the estimation result of linear probability model of $\mathbbm{I}\{Spend_{it}\}$ between adopters and offline-only customers. Table A63 presents the DID estimation results conditional on customer-month with non-zero spend between adopters and offline-only customers.
Table A62: Linear Probability Model with DID Specifications – Adopters vs. Offline-only
Study: Covid Black Friday Loyalty Program
DV $\mathbbm{I}\{Spend_{it}\}$
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 0.4792 (0.0004) 0.5002 (0.0004) 0.4766 (0.0004)
[0.000] [0.000] [0.000]
Adopter 0.0947 (0.0015) 0.0942 (0.0047) 0.0586 (0.0021)
[0.000] [0.000] [0.000]
Post -0.1627 (0.0004) -0.1810 (0.0004) -0.0786 (0.0005)
[0.000] [0.000] [0.000]
Adopter * Post 0.1098 (0.0016) 0.1033 (0.0016) 0.1085 (0.0048) 0.0910 (0.0046) 0.0570 (0.0030) 0.0933 (0.0029)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 22,204,608 22,204,608 9,505,999 9,505,999 4,403,449 4,403,449
R2 0.02892 0.28000 0.03368 0.34101 0.01031 0.26334
Within R2 0.00073 $9.3\times 10^{-5}$ 0.00156
Clustered (Customer) standard-errors are presented in standard brackets, while p-values are shown in square brackets.
Table A63: DID Estimates for Spend (Conditional on Purchase) – Adopters vs. Offline-only
Study: Covid Black Friday Loyalty Program
DV Spend
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 584.6 (0.9550) 583.1 (0.9740) 697.9 (1.448)
[0.000] [0.000] [0.000]
Adopter 212.2 (4.176) 206.7 (12.83) 154.2 (7.012)
[0.000] [0.000] [0.000]
Post 101.9 (1.061) 26.81 (0.8289) 8.823 (1.252)
[0.000] [0.000] [0.000]
Adopter * Post 127.7 (4.204) 135.4 (3.953) 52.78 (10.73) 65.40 (9.385) -52.10 (8.243) 50.28 (6.334)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 8,187,536 8,187,536 3,744,951 3,744,951 1,939,763 1,939,763
R2 0.02034 0.47201 0.00178 0.50834 0.00501 0.62202
Within R2 0.00084 $4.37\times 10^{-5}$ 0.00150
Clustered (Customer) standard-errors are presented in standard brackets, while p-values are shown in square brackets.
Event Adopters vs. Organic Adopters
Table A64 presents the estimation result of linear probability model of $\mathbbm{I}\{Spend_{it}\}$ between event adopters and organic adopters. Table A65 presents the DID estimation results conditional on customer-month with non-zero spend between event adopters and organic adopters.
Table A64: Linear Probability Model with DID Specifications – Event Adopters vs. Organic Adopters
Study: Covid Black Friday Loyalty Program
DVs: $\mathbbm{I}\{Spend_{it}\}$
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 0.5557 (0.0029) 0.5889 (0.0057) 0.5297 (0.0022)
[0.000] [0.000] [0.000]
Covid_Adopter 0.0244 (0.0033)
[0.000]
Post -0.0451 (0.0031) -0.0569 (0.0060) -0.0162 (0.0030)
[0.000] [0.000] [0.000]
Covid_Adopter * Post -0.0100 (0.0036) 0.0039 (0.0034)
[0.005] [0.251]
BlackFriday_Adopter 0.0171 (0.0098)
[0.073]
BlackFriday_Adopter * Post -0.0488 (0.0100) -0.0361 (0.0095)
[0.000] [0.000]
Loyalty_Adopter 0.0634 (0.0076)
[0.000]
Loyalty_Adopter * Post -0.1187 (0.0093) -0.0202 (0.0095)
[0.000] [0.032]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 76,842 76,842 197,658 197,658
R2 0.00210 0.25272 0.00569 0.28908 0.03886 0.27038
Within R2 $2.36\times 10^{-6}$ 0.00036 0.01574
Clustered (Customer) standard-errors are presented in standard brackets, while p-values are shown in square brackets.
Table A65: DID Estimates for Spend (Conditional on Purchase) – Event Adopters vs. Organic Adopters
Study: Covid Black Friday Loyalty Program
DVs: Spend
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 778.8 (7.845) 779.6 (14.49) 839.6 (7.031)
[0.000] [0.000] [0.000]
Covid_Adopter 23.90 (9.168)
[0.009]
Post 237.0 (7.804) 103.2 (13.31) -30.62 (8.003)
[0.000] [0.000] [0.000]
Covid_Adopter * Post -9.354 (9.147) -3.982 (8.373)
[0.316] [0.623]
BlackFriday_Adopter 30.84 (29.04)
[0.288]
BlackFriday_Adopter * Post -75.22 (21.45) -64.19 (19.68)
[0.000] [0.001]
Loyalty_Adopter 132.3 (27.63)
[0.000]
Loyalty_Adopter * Post -232.3 (27.67) -36.03 (23.92)
[0.000] [0.126]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Standard-Errors Customer
Observations 949,448 949,448 42,299 42,299 110,339 110,339
R2 0.00803 0.44326 0.00248 0.45643 0.02928 0.56886
Within R2 $7.04\times 10^{-7}$ 0.00047 0.01280
Clustered (Customer) standard-errors are presented in standard brackets, while p-values are shown in square brackets.