### v2

# Channel Adoption Pathways and Post-Adoption Behavior

Abstract
The rapid growth of digital shopping channels has prompted many traditional retailers to invest in e-commerce websites and mobile apps. While prior literature shows that multichannel customers are more valuable, it overlooks how the motive for adopting a new channel shapes post-adoption behavior. Using transaction-level data from a major Brazilian pet supplies retailer, we study offline-only consumers who adopt online shopping via four distinct pathways: organic adoption, the COVID-19 pandemic, Black Friday promotions, and a loyalty program. We examine how these pathways affect post-adoption spend, profitability, and channel usage using consumer-level panel data and difference-in-differences estimates. We find that all adopters increase spending relative to offline-only consumers, but their post-adoption behaviors differ systematically by adoption motive. Promotion-driven adopters engage in forward buying and exhibit lower subsequent profitability, whereas COVID-19 adopters display stronger offline persistence consistent with consumer inertia and habit theory. Our findings have important managerial implications: firms should design promotions that discourage stockpiling, reinforce habits among customers pushed online by external shocks, and explicitly account for heterogeneity in channel adoption motives when forecasting customer lifetime value and assessing the breakeven and ROI of promotions designed to induce the adoption of new channels.
Keywords: Retailing, Multichannel, Online Shopping, Customer value
1 Introduction
Background and Motivation
Online shopping has grown rapidly over the last two decades with the proliferation of computers and smartphones. In response, many grocery and retail chains (e.g., Walmart, Target, Costco), historically brick-and-mortar, now also sell through websites and mobile apps. Given the size of retail and the rise of digital channels, it is important to quantify how adopting digital shopping channels affects consumers’ behavior (Goyal 2023; Bloomberg 2020).
Prior research has found that multichannel consumers (those who buy using multiple channels of a retailer) tend to spend more than single-channel consumers; see $\S$2. However, this work largely ignores why consumers adopt new channels, implicitly assuming that the benefits from multichannel adoption are the same regardless of the reasons for adoption of new channels. Yet consumers may adopt a retailer’s online shopping channels (from previously being offline-only) for many different reasons, and it is unlikely that channel adoption pathways have no impact on post-adoption behavior. We classify potential online adoption pathways into three categories: (1) macro-environmental shocks such as COVID-19, nearby store closures, or severe weather events that push consumers online; (2) firm-driven promotions that require engagement with online channels to obtain benefits – this can include both short-term promotional events and holiday sales as well as long-term offers such as loyalty programs; and (3) an organic pathway, in which consumers adopt online shopping of their own volition due to the convenience of an additional channel.
It is important for managers to understand heterogeneity in post-adoption behavior across these pathways. If changes in purchase behavior after online adoption depend on the motivations for adoption, then it is difficult to extrapolate the behavior of consumers who adopt organically (without external incentives) to those induced to adopt through marketing campaigns or external factors. Moreover, if differences relative to organic adopters vary by the type of external factor (e.g., macro shock vs. promotion), forecasting post-adoption behavior becomes even more complicated. Conversely, if consumers who adopt online shopping for different reasons behave similarly post-adoption, managers can extrapolate from organic adopters and worry less about the underlying reason for adoption. These considerations motivate our research questions.
Research Questions and Theoretical Predictions
We consider two research questions. First, how does the adoption of a retailer’s digital shopping channels (e-commerce website and/or mobile app) affect the purchase behavior of customers who were previously purely offline-only (brick-and-mortar) shoppers? We focus on three managerially salient outcomes: (1) total spend, (2) profitability, and (3) stickiness to offline channels post-adoption. Second, conditional on adopting online shopping, how does post-adoption behavior differ across the three adoption pathways (organic, macro-shock-driven, and promotion-driven)?
Building on the multichannel literature and behavioral frameworks (discussed in detail in $\S$2 and $\S$3), we hypothesize that customers who adopt a firm’s online channels and become multichannel shoppers will, on average, increase their total spending relative to offline-only customers. Beyond this average effect, we anticipate substantial heterogeneity across adoption pathways. Consumer inertia and habit theory (Verplanken and Wood 2006; Wood and Neal 2009) suggest that when adoption is induced by external shocks or incentives rather than intrinsic preference, consumers are more likely to revert toward prior routines once the stimulus subsides. Thus, relative to externally driven adopters, organically adopting customers should exhibit a higher post-adoption share of spending on online channels and faster growth in that share. We further anticipate that such differences will be more pronounced for older, habit-driven customers.
For adoption spurred by time-limited promotions, theories of forward buying and demand distortion (Neslin, Henderson, and Quelch 1985; Hendel and Nevo 2006; Van Heerde and Neslin 2017) imply that promotion-driven adopters may temporarily inflate their spending during the promotional period by stocking up, but then reduce spending afterwards. This mechanism can dampen or even offset the long-run spending uplift typically associated with multichannel adoption. Finally, because profit margins generally differ across online and offline channels, heterogeneity in the post-adoption mix of spending across channels is likely to translate into systematic differences in profitability across adopter types, even when their total spend is similar.
In summary, our theoretical framework predicts that (i) adopters will increase spending relative to offline-only customers, but (ii) the magnitude and persistence of this effect, as well as the resulting channel mix and profitability, will differ systematically across organic, macro-shock-driven, and promotion-driven adoption pathways, driven by consumer inertia and habit-formation, forward buying around promotions, and the relative profitability of the two channels.
Empirical Setting and Methodology
To test these predictions, we use data from a large pet supplies retailer in Brazil with an extensive network of brick-and-mortar stores and an online presence (website and mobile app). Our data comprise all consumer purchases across all channels from 2019 to 2024. During this period, there were three key events: a macro-environmental shock (COVID-19) and two firm-driven promotions designed to incentivize online shopping. This allows us to observe a variety of adoption pathways.
We construct four groups of adopters of online shopping who previously only shopped at the firm’s physical stores: (1) covid adopters, who started shopping online during the onset of COVID-19; (2) Black Friday adopters, who adopted online shopping due to deep discounts available online during Black Friday; (3) Loyalty Program adopters, who adopted online shopping during the launch of a loyalty program that had to be activated on the retailer’s mobile app; and (4) organic adopters, who started shopping online without any identifiable external stimulus. In addition, we consider a control group of non-adopters or offline-only customers (consumers who did not adopt online shopping during our observation period).
Using these data, we construct consumer-level monthly panel data to quantify revenue, profitability, and the share of spend in the offline and online channels, and use this panel to answer our research questions. Specifically, we measure the Average Treatment Effect on the Treated (ATT) of online adoption on post-adoption behavior for different adoption pathways using a difference-in-differences framework with individual and time fixed effects. Our results should be interpreted as descriptive with appropriate controls; we are not focused on estimating the causal Average Treatment Effect (ATE) of online adoption. Each adoption pathway represents a different type of consumer self-selection, i.e., different types of consumers self-select into adopting online shopping as a result of each stimulus. Our goal is not to rule out selection effects or separate out causal effects from selection effects; rather, we seek to document whether and how different reasons for adopting online shopping shape post-adoption behavior. From a managerial perspective, firms cannot randomize online adoption across consumers; they can only “encourage” adoption through promotional activities. Thus, understanding how different adoption pathways affect post-adoption purchase behavior is of primary interest to managers as opposed to a causal ATE.
Main Findings, Managerial Implications, and Contributions
We now summarize our main findings and their managerial and substantive implications. First, relative to offline-only customers, adopters of online shopping have higher spend and profitability in the post-adoption period, consistent with our predictions and prior evidence on multichannel shoppers. However, the post-adoption behaviors of different adopter types diverge sharply across three key dimensions: total spend, profitability, and the share of offline/online spending. While covid adopters spend similarly to organic adopters post-adoption, promotion-driven adopters (Black Friday adopters and Loyalty Program adopters) spend less. Consistent with our theoretical predictions, this reduced spend is driven largely by forward buying during the promotional period: these customers stock up and then reduce future purchases, which does not occur for covid adopters.
In terms of the share of offline/online spend post-adoption, and consistent with consumer inertia and habit theory, all three adopter groups motivated by external stimuli (covid adopters, Black Friday adopters, and Loyalty Program adopters) have a lower online spend share post-adoption relative to organic adopters. Moreover, covid adopters show slower growth in online spend share over time, consistent with their being older and more set in their habits. Black Friday adopters, by contrast, are younger and show a faster growth rate in online spend share relative to organic adopters, although they still take time to catch up. There is no difference in trends for online spend share between Loyalty Program adopters and organic adopters, likely because these two groups are similar demographically and because the loyalty program involves monthly online activation, reinforcing online shopping habits. Post-adoption profits for the different adopter groups depend on both total spend and the offline/online spend share. We find that despite similar spend, covid adopters are more profitable post-adoption than organic adopters, due to their higher spend share in the offline channel, where margins are higher. In contrast, Black Friday adopters and Loyalty Program adopters are less profitable due to their lower overall spending in the post-adoption period.
Taken together, these findings show that the reason for a consumer’s decision to adopt online shopping significantly shapes long-term multichannel behavior and that the observed patterns are broadly consistent with theories of consumer inertia, age-related habit strength, and forward buying around promotions. Because our mechanisms are grounded in these behavioral forces, the insights are likely to apply more broadly in settings where firms use promotions or external triggers to induce adoption of new channels or digital services.
Our results have important implications for two core managerial tasks: forecasting Customer Lifetime Value (CLV) and evaluating the ROI of promotions. We show that ignoring adoption pathways can lead to substantial bias in incremental CLV forecasts. In our setting, using organic adopters as proxies for promotion-induced adopters overstates the incremental CLV of promotion-induced adopters by roughly 70–80%, and pooling all adopters into a single group similarly misstates the CLV of both promotion-induced and covid adopters. Likewise, breakeven and ROI calculations for promotions can be severely distorted if managers implicitly assume that promotion-induced adopters will behave like organic adopters: in our data, naive breakeven times for online-adoption promotions can be off by factors of three to six. These results imply that firms should (a) explicitly account for channel adoption pathways when estimating incremental CLV from online adoption and (b) set pathway-specific caps on acceptable acquisition and promotional costs.
Overall, our analysis makes both substantive and managerial contributions. Substantively, we extend the multichannel shopping literature by introducing channel adoption pathways as a key dimension of heterogeneity and by documenting how organic, macro-shock-driven, and promotion-driven adopters follow systematically different post-adoption trajectories in spend, channel mix, and profitability. Managerially, we provide a cautionary tale against naively extrapolating the behavior and profitability of one adopter group (e.g., organic adopters) to others and offer a framework for incorporating adoption pathways into CLV forecasting, promotion design, and ROI/breakeven analysis. Firms that explicitly condition on adoption pathways when designing and evaluating interventions to move customers online can better avoid overpaying for low-value adopters and more effectively nurture high-value ones.
2 Related Literature
Paper Setting Channels Methodology Adoption Pathways
Our Paper Pet Supplies Physical stores only to Diff-in-diff Organic +
adding website and/or multiple external
mobile app stimuli
Narang and Shankar (2019) Video games Physical stores + website Diff-in-diff Organic only
and Electronics to adding mobile app
Montaguti, Neslin, and Valentini (2016) Books Mail order, fax, phone, Cross-sectional Organic +
(subscribers only) SMS, Internet, mailing campaign
and bookstores
Wang, Malthouse, and Krishnamurthi (2015) Online grocery Website to adding Diff-in-diff Organic only
mobile app
Kushwaha and Shankar (2013) Multiple categories Website and catalog Cross-sectional Organic only
Ansari, Mela, and Neslin (2008) Apparel Website and catalog Within-customer Organic only
Venkatesan and Ravishanker (2007) Apparel Full price stores, Within-customer Organic only
Discount Stores,
and website
Kumar and Venkatesan (2005) Technology Salespersons, direct mail, Cross-sectional Organic only
telephone and website
Thomas and Sullivan (2005) Retail Catalog, physical stores, Cross-sectional Organic only
website
Table 1: Prior Empirical Literature on Multichannel Shopping in Retail Settings
Our work relates to the empirical literature on how multichannel versus single-channel shopping affects consumers’ purchase behavior. Table 1 summarizes the most closely related papers, and below we highlight how our study differs from and contributes to this literature. For broader reviews, see Neslin et al. (2006); Neslin and Shankar (2009); Verhoef (2012).
From a substantive perspective, our paper contributes in two main ways (see Table 1).
  • $\bullet$
    First, the existing literature overlooks heterogeneity in channel adoption pathways, implicitly assuming that the value of multichannel customers is independent of the motivations underlying their adoption decisions. As Table 1 shows, almost all prior work focuses on organic adopters. Montaguti, Neslin, and Valentini (2016) is an exception, using a mailing campaign to nudge consumers into adopting multiple channels. However, they do not compare organic multichannel customers to those nudged by the campaign, whereas this comparison is central in our paper.
  • $\bullet$
    Second, we study a specific single-channel to multichannel transition that has not been examined before and is particularly relevant for traditional brick-and-mortar retailers such as Walmart, Target, and Costco: moving from shopping only in physical stores to using both physical stores and the retailer’s website/mobile app. Earlier work largely considers transitions from any single channel to some combination of multiple channels (e.g., physical stores, catalogs, phone, mail order, salespeople, mobile apps, websites). These findings may not generalize to the physical-store-to-e-commerce transition because the channels differ in important ways. [^1] Websites and apps provide easy search and comparison between products and between competing retailers, which is not true for catalogs, salespersons, or phone shopping. In many cases, prices online are also lower than in physical stores, which again differs from other channels.
    Moreover, our setting is a fast-moving consumer packaged goods retailer (pet supplies), whereas much of the earlier literature examines books, apparel, video games, electronics, and technology. These differ from our context along several dimensions, including purchase frequency, competitive intensity, and product type (often more hedonic than utilitarian), which has been shown to matter for multichannel effects (Kushwaha and Shankar 2013).
Thus, we extend the multichannel literature by (i) examining how post-adoption behavior varies with channel adoption pathways and providing theoretical explanations for these differences, and (ii) focusing on a practically important channel transition—offline-only to combined offline and online usage—that has not been separately studied.
From a methodological perspective, prior research has used three broad approaches to documenting the value of adopting new channels: cross-sectional, within-customer, and difference-in-differences (DID). Early work (Thomas and Sullivan 2005; Kumar and Venkatesan 2005; Kushwaha and Shankar 2013; Montaguti, Neslin, and Valentini 2016) uses cross-sectional comparisons of multichannel versus single-channel customers and concludes that multichannel customers tend to be more valuable. However, cross-sectional comparisons lack individual-specific controls and cannot isolate how behavior changes after multichannel adoption. A second stream (Venkatesan and Ravishanker 2007; Ansari, Mela, and Neslin 2008) employs within-customer analysis, comparing a customer’s purchase behavior before versus after adopting multiple channels. This work also finds higher spending after multichannel adoption, but without an appropriate control group of non-adopters, it requires strong assumptions that no other time-varying factors affect behavior. More recent work (Wang, Malthouse, and Krishnamurthi 2015; Narang and Shankar 2019) uses DID with non-adopters as a control group, addressing many of these concerns. Our paper also employs a DID framework with non-adopters as a control group, but brings this methodology to a new setting and focuses on heterogeneity across adoption pathways, which has not been explored in prior empirical work.
3 Background, Goals, and Theory
As discussed in $\S$2, the existing multichannel literature has not considered heterogeneity in channel adoption pathways and has implicitly assumed that all multichannel customers behave similarly, regardless of why they adopted new channels to shop with a given retailer. However, consumers may adopt a retailer’s online shopping channels for many different reasons, and it is plausible that these different pathways affect post-adoption behavior. At a high level, there are three broad classes of online adoption pathways: (1) organic adoption, where customers voluntarily adopt a retailer’s online shopping channels due to the added convenience of an additional channel; (2) macro-environmental shocks, such as public health crises (e.g., COVID-19), nearby store closures, or severe weather events that prevent store visits; and (3) firm-driven promotions, where retailers explicitly incentivize adoption of their online channels.
Each pathway represents a different type of consumer self-selection, i.e., different types of consumers choose to adopt a retailer’s online shopping channels in response to each stimulus. Our goal is not to rule out selection effects; rather, we seek to document whether and how different reasons for adopting online shopping with a retailer affect post-adoption behavior. From a managerial perspective, firms cannot randomize online adoption across consumers; they can only “encourage” adoption through promotional activities and observe how different encouragements translate into different adoption pathways and subsequent purchases. [^2] We refer readers to the broader literature on “encouragement design” in marketing and the social sciences. Broadly, this literature distinguishes between interventions where firms can assign consumers to treatment arms (e.g., pricing, advertising, web page design) and those where consumers must self-select into the focal action (e.g., signing up for a newsletter, adopting online shopping) (Bradlow 1998; Syrgkanis et al. 2019; Mummalaneni, Yoganarasimhan, and Pathak 2023). In the latter, firms can only encourage the action (e.g., via promotions), and even a fully randomized experiment cannot eliminate selection because the outcome of interest (online adoption) is itself chosen by the consumer. In such settings, the natural estimand is the Average Treatment Effect on the Treated (ATT), rather than the Average Treatment Effect (ATE). Accordingly, we aim to document the ATT of online adoption on post-adoption behavior for different pathways and to relate differences in these ATTs to theories of consumer behavior.
We next discuss behavioral mechanisms that link channel adoption pathways (organic, macro shocks, and promotions) to post-adoption outcomes. First, customers who adopt online channels organically may behave differently from those who adopt due to external factors. Organic adoption most plausibly signals genuine preference or readiness to use online channels, rather than being forced or incentivized by external shocks or promotions. Consumer inertia and habit theory (Verplanken and Wood 2006; Wood and Neal 2009) suggest that adoption driven by external factors may temporarily disrupt existing routines, but absent reinforcement, consumers may revert toward prior habits and use the newly adopted channels less intensively. Organic adopters, by contrast, are more likely to invest in learning and to enjoy the convenience of online channels, leading to more sustained use.
If so, we can make several predictions about differences between organic adopters and externally driven adopters. (1) The post-adoption share of spending on online channels should be higher for organic adopters than for externally driven adopters. (2) Because profit margins typically differ across channels, this difference in online share may translate into lower (higher) profitability for organic adopters if online channels have lower (higher) margins than offline channels. Prior work also shows that older consumers tend to be more entrenched in their habits and exhibit higher inertia than younger consumers (Yang and Ching 2014). Thus, we expect divergence in online usage between organic adopters and externally driven adopters to be more pronounced when external factors disproportionately trigger adoption among older consumers (e.g., public health concerns like COVID-19 or promotions targeted at older segments). This leads to a third prediction: (3) the gap in online share between organic adopters and externally driven adopters should be larger when the external stimulus induces relatively older consumers to adopt.
Promotions offering discounts or rewards in return for increased spending during a limited period may also lead to distinctive post-adoption behavior relative to other external shocks. Prior research documents forward buying and demand distortion (Neslin, Henderson, and Quelch 1985; Hendel and Nevo 2006; Van Heerde and Neslin 2017), whereby time-limited promotions induce stockpiling and reduce demand in the post-promotion period. This mechanism is particularly relevant for promotions designed to induce customers to start shopping online and motivates a fourth prediction: (4) consumers who adopt in response to promotions may increase spending during the promotional period but reduce spending afterwards, thereby attenuating (or even reversing) the longer-run spending boost normally associated with multichannel adoption.
Based on these mechanisms, we seek to empirically test the following predictions:
  • $\bullet$
    H1: Customers who adopt a firm’s online channels and become multichannel shoppers will increase their spending post-adoption relative to offline-only customers (consistent with prior findings on multichannel shoppers; see $\S$2).
  • $\bullet$
    H2: Customers with different channel adoption pathways will exhibit different post-adoption behavior in terms of spend, profitability, and channel mix. These differences will arise from a combination of customer demographics, consumer habit theory, and forward buying. Specifically,
    • $\bullet$
      H2a: Consumers who adopt in response to time-limited promotions may increase spending during the promotional period but reduce spending thereafter, dampening (or reversing) the spending boost from multichannel adoption in the post-promotion period.
    • $\bullet$
      H2b: Externally driven adopters (vs. organic adopters) may be less likely to show the same level and growth in the share of spending on the newly adopted online channel, as predicted by consumer inertia and habit theory.
    • $\bullet$
      H2c: The difference in online share between organic adopters and externally driven adopters may be larger when external factors cause relatively older consumers to adopt.
    • $\bullet$
      H2d: Differences in online share may translate into lower (higher) profitability for organic adopters if online channels have lower (higher) margins relative to offline channels.
4 Setting and Data
4.1 Setting
Recall that the key goals of the paper are: (1) to empirically document the effect of adopting online shopping on post-adoption consumption behavior, (2) to test whether different channel adoption pathways lead to different post-adoption behaviors, and (3) to verify that the differences in post-adoption behavior can be predicted using theoretical underpinnings of purchase behavior in retail settings, so as to provide generalizable insights for managers and researchers. To accomplish these goals, we need data from a setting where we can observe not only when an offline-only consumer adopts online shopping but also the driver of this decision, i.e., the adoption pathway. Further, we need sufficient variants in channel adoption pathways to meaningfully connect the post-adoption behavior to theory. Therefore, we focus on a retail data source that satisfies these requirements.
Our data comes from one of the largest Brazilian pet supplies retailers, which sells a wide variety of products for pets, such as pet food and snacks, medicines/pharmacy items, hygiene and grooming products, and pet toys and accessories. Pet supplies is a large and competitive industry with about \$250 billion in annual revenue globally (Fortune 2024). The firm has both offline and online channels. The retailer operates more than 200 physical stores in Brazil, which serve as the offline channel, and has a website as well as a mobile app, which serve as the online channel. Moreover, our data span 2019-2024 during which there were three different events including a macro/environmental shock (COVID-19) and two different types of promotions designed to incentivize online shopping (discussed in more detail in $\S$5) which led to many consumers adopting online shopping with the firm. Further, our data are quite rich, and we are able to observe when a consumer adopts online shopping (and thus can connect the timing of adoption to the channel adoption pathway) and both the pre- and post-adoption consumption behavior. For these reasons, it is an ideal setting for our study.
4.2 Data
Our data includes transaction-level records of all individual customers from both online and offline channels during the 63-month period from January 1st, 2019, to March 31st, 2024. Each transaction record is uniquely identified by an order number with one or more purchased items. For each purchased item, we have information on the unit price, quantity purchased, brand, purchased channel (online/offline), subcategory, and category (e.g., food, hygiene).
The firm tracks customer purchases using a National Identification Number (similar to a Social Security Number), which is unique to each individual in Brazil. Customers who purchase online (either through the website or the app) must be logged in to their profile, and therefore, all online purchases can be traced back to the National ID. Customers are also asked for their National ID when making an in-store purchase; however, this is not mandatory. Discussions with the retailer revealed that while most offline transactions are also linked with a customer profile, about a quarter of in-store transactions are not linked to a National ID. This number has remained fairly steady during the entire observation period and was not affected by COVID-19 or other marketing events [^3] Because the number of offline transactions not linked to customer IDs remained similar through our observation period and exhibited no significant shifts in response to any of the events that spurred online adoption (including COVID-19), it is unlikely that this issue systematically biases our findings on the adoption of online shopping.
[…]
5 Estimation Sample and Summary Statistics
We first discuss the sample construction process for the three focal events in $\S$5.1, and present the details of the samples for each of the three focal events in $\S$5.2, and then present some summary statistics of the three samples of interest in $\S$5.3.
5.1 Sample Construction
We now discuss how we construct the samples used in our empirical analyses and the main outcome variables of interest.
Focal events that triggered different pathways of adoption: Since we seek to connect channel adoption pathway to post-adoption behavior, we consider three events that each led to a surge in online adoption: (1) a macro-environmental shock (COVID-19), (2) a short-term promotion (Black Friday sales), and (3) a long-term loyalty program. The first event was a macroeconomic shock outside the firm’s control, whereas the latter two were firm-driven marketing activities.
Customer cohorts: For each event, we first define three customer cohorts – (a) event adopters, consumers who adopted online shopping due to the event, (b) organic adopters, consumers who adopted organically (without external incentives) around the same time frame as the event, and (c) offline-only customers, consumers, who did not adopt online shopping at all.
Time periods: Next, we establish the pre-adoption, adoption, and post-adoption periods for each focal event as follows:
  • $\bullet$
    The pre-adoption period is defined as the common period during which all three groups of customers purchase only through physical stores (i.e., before the onset of the focal event).
  • $\bullet$
    The adoption period is defined as the time during which adopters started purchasing online. Because each event has two sets of adopters, we have two adoption periods:
    • $\bullet$
      Adoption period for event adopters: This is the period during which the focal event took place, and induced customers to adopt online shopping. For instance, in the case of COVID-19, this is the first month and a half since the onset of COVID-19. Similarly, for the Black Friday event, the adoption period for event adopters is simply the day of the Black Friday sale.
    • $\bullet$
      Adoption period for organic adopters: Since organic adopters serve as a control group for event adopters, for each event, we choose consumers who organically adopted online shopping just before the onset of the focal event, and this period varies based on the event.
  • $\bullet$
    The post-adoption period refers to the months after the adoption periods of both event adopters and organic adopters.
See Table 2 for an overview of these periods for each of the three events. In our estimation sample for each event, we only include data from the pre- and post-adoption periods for that event. We exclude the adoption periods for both event and organic adopters from the estimation sample because consumers are in flux during that time (where some adopters have adopted while others have not). Furthermore, by excluding these observations, we avoid capturing any abnormal purchase behaviors associated with the initial adoption of online shopping and the onset of specific events (i.e., COVID-19, Black Friday, Loyalty Program). As such, it allows us to cleanly identify the post-online-adoption difference in customers’ purchase behaviors; see Gu and Kannan (2021) for a similar specification. We now summarize the choice of these three periods for each event.
  • $\bullet$
    COVID-19 event: We use one year of data before the adoption period as the pre-adoption period (the calendar year of 2019). For the post-adoption period, we use all the data from May 2020 till April 2023444Pilot studies for the firm’s Loyalty program started running from May 2023; so we cut off the post-adoption period with April 2023..
  • $\bullet$
    Black Friday event: We use eight months of data, starting from January 2019 to October 2020, as the pre-adoption period. For the post-adoption period, we use data from December 2019 (following the Black Friday event) until the next Black Friday.
  • $\bullet$
    Loyalty Program event: We use the data from five months before the adoption period, and all the data available (from the end of the adoption period till March 2024) as the post-adoption period.
Table 2: Pre-adoption, Adoption, and Post-adoption Periods Under Each Event
COVID-19 Black Friday Loyalty Program
Pre-adoption 2019-01-01 – 2019-12-31 2019-01-01 – 2019-10-31 2023-01-01 – 2023-05-31
Adoption
     Organic 2020-01-01 – 2020-02-29 2019-11-14 – 2019-11-28 2023-06-01 – 2023-07-31
     Event-Induced 2020-03-16 – 2020-04-30 2019-11-29 2023-08-01 – 2023-09-30
Post-adoption 2020-05-01 – 2023-04-30 2019-12-01 – 2020-10-31 2023-10-01 – 2024-03-31
We discuss each event and the definition of the three cohorts of consumers in detail below.
Outcome metrics: For all our analyses, we consider three metrics that quantify a consumer’s shopping behavior in a given month: (1) Total Spend, (2) Profitability, and (3) Stickiness to the offline channels post-adoption. Formally, we are interested in the Average Treatment Effect on Treated (ATT) of adopting online shopping on these three metrics for each of the three focal events. Each of these metrics is defined at the customer-month level.
5.2 Estimation Samples for the Focal Events
We now describe the estimation sample used in the analysis of each of the three focal events.
5.2.1 Event 1: COVID-19
Figure 1: Number of Offline-only Customers First Adopting Online Shopping due to COVID-19 (by month, Black Friday (by day), and Loyalty Program (by month)
(a) COVID-19
(b) Black Friday
(c) Loyalty Program
The COVID-19 pandemic was an exogenous macro-environmental shock that prompted many consumers to adopt online shopping. Figure 1(a) shows the number of customers who adopted online shopping through the retailer’s website and/or mobile app (from previously only shopping at physical stores) for each month for our observation period. As we can see, during March and April 2020, there was a large peak in the number of offline customers who first switched to online shopping; the number of adopters sharply increased from approx. 7,000 in January and February 2020 to approximately 20,000 in March and April 2020555In Figure 1(a), we also see smaller peaks associated with the second and third waves of COVID-19.. Customers who adopted online shopping during the onset of COVID-19 may have done so for reasons like health concerns, social distancing norms, etc., whereas the customers who adopted online shopping before that did so of their own volition, for other reasons (e.g., the convenience of online shopping), unrelated to COVID-19.
We consider the set of customers who made their first offline purchase before 2019-12-31 and were offline-only through the end of 2019 and continued shopping with the firm after the onset of COVID-19. Then, we categorize each consumer from this set into one of the following three groups (see second column in Table 2 for exact dates):
  • $\bullet$
    Covid adopters: Customers who made their first online purchase during the onset of COVID-19, i.e., from 2020-03-16 6662020-03-16 was the date when the Brazilian government started implementing physical distancing and confinement measures in response to COVID-19 (Faria de Moura Villela et al. 2021). to 2020-04-30.  [^7] The set of people who adopted during this period could also include consumers who would have organically adopted even without the COVID-19 shock. We have no clear way of distinguishing such potential organic adopters from those who adopted online shopping as a result of COVID-19. Nevertheless, based on Figure 1(a), it seems like approximately two-thirds of these consumers switched due to COVID-19. That is, a significant majority of these consumers switched due to COVID-19, and we refer to this cohort as covid adopters for simplicity. Please see $\S$9.1 for a detailed discussion on the implications of this potential mixing on the ATT estimates for COVID-19 and the other two events.
  • $\bullet$
    Organic adopters: Customers who adopted online shopping organically in the period just before the onset of COVID-19. They made their first online purchase between 2020-01-01 and 2020-02-29, i.e., the two months before the onset of COVID-19.
  • $\bullet$
    Offline-only customers: Customers who are active through both the pre- and post-adoption periods (i.e., at least one purchase before 2019-12-31 and after 2020-05-01), and never make online purchases during the estimation period.
In total, we have 33,246 covid adopters, 12,799 organic adopters, and 573,934 offline-only customers for this covid analysis. Note that a large fraction of consumers were able to remain offline-only during COVID-19 because the focal retailer’s stores were allowed to remain open, as pet supplies retailing was deemed an essential service by the Brazilian government.
5.2.2 Event 2: Black Friday Sales
Next, we discuss the retailer’s Black Friday sales event in 2019, characterized by a one-day, time-limited flash sales event offering discounts exclusively for online purchases. [^8] We focus on the Black Friday sales in 2019 and do not consider Black Friday sales in 2020 and afterwards because online adoption during those years may have been driven by COVID-19 or other marketing campaigns (e.g., the loyalty program), rather than purely driven by the sales event itself like 2019. Further, unlike 2019, the firm did not have major Black Friday sales in later years. Because of the exclusive online nature of the event, consumers interested in these deals were motivated to adopt online shopping. Figure 1(b) indicates a peak in the number of adopters on 2019-11-29, which is the date that Black Friday sales took place.
For this estimation sample, we consider the cohort of customers who made their first offline purchase before 2019-10-31 and continued to be the firm’s customers after the Black Friday sales date. Then, we categorize each consumer from this cohort into one of the following three groups (see third column in Table 2 for exact dates):
  • $\bullet$
    Black Friday adopters: Customers who made their first online purchase during the 2019 Black Friday sales.
  • $\bullet$
    Organic adopters: Customers who adopted online shopping organically in the two weeks preceding the 2019 Black Friday sales.
  • $\bullet$
    Offline-only customers: Customers who are active through both the pre- and post-adoption periods, and never make online purchases during the estimation period.
In total, we have 1380 Black Friday adopters, 3061 organic adopters, and 534,871 offline-only customers for the Black Friday analysis.
5.2.3 Event 3: Loyalty Program
In August and September 2023999Pilot studies for the loyalty program ran in two states (Rio Grande do Sul and Paraná) as early as May. Therefore, we exclude these two states from the estimation sample for this event., the retailer launched a loyalty program and extensively advertised it to customers through email and store marketing. Customers who wanted to join the LP had to use the mobile app (no offline options were available for activating the loyalty program). After completing the online activation, customers could access and activate monthly special offers and receive discounts on both online and offline purchases [^10] Note that customers had to activate each month for participation in each month’s special offers.. Figure 1(c) presents the number of adopters by month since January 2023 and shows clear peaks in the number of adopters during August and September 2023, coinciding with the launch of the loyalty program. This suggests that online-only activation of the newly available loyalty program prompted a significant number of customers to adopt online shopping during the roll-out period. For this event, we consider the cohort of customers who made their first offline purchase before 2022-12-31 and continued to be the firm’s customers during the post-online-adoption period (i.e., after 2023-10-01). Then, we categorize each consumer from this cohort into one of the following groups (see last column of Table 2 for details):
  • $\bullet$
    Loyalty Program adopters: Customers who made their first online purchase in August or September 2023, with their first online purchase occurring after they activated the loyalty program online. In such cases, we can attribute the online loyalty activation as a reason for online adoption, as the activation precedes the first online purchase and both events occur within a similar time frame.
  • $\bullet$
    Organic adopters: Customers who made their first online purchase in June or July 2023, prior to the launch of loyalty program. [^11] Note that these consumers adopt online shopping organically, but may still activate loyalty offers after their online adoption. We will take this into account in our empirical analysis.
  • $\bullet$
    Offline-only customers: Customers who are active through both the pre- and post-adoption periods (i.e., at least one purchase before 2022-12-31 and after 2023-10-01), and never make online purchases during the estimation period.
In total, we have 1567 loyalty program adopters, 16,746 organic adopters, and 396,396 offline-only customers for this analysis.
5.3 Summary Statistics
For each consumer in the estimation sample of any focal event, we aggregate their transaction-level data into customer-month-level panel data over the customer’s entire purchase history within the relevant period (January 2019 to March 2024), starting from their initial purchase month and ending with the last observed purchase month. For example, if a customer purchased for the first time in our data in August 2019 and the last observed purchase is in July 2020, then the panel for this customer spans 12 months, from August 2019 to July 2020. If there are intermediate months in which a customer did not make any purchase, the values for the monthly purchase-related variables are set to zero.
[…]
Across our three samples for COVID-19, Black Friday, and the Loyalty Program, we find consistent differences between the two groups of adopters (organic adopters and event adopters) and offline-only customers. For instance, the adopter groups are consistently younger, higher income, and have higher spend with the retailer pre-adoption, relative to offline-only customers.
However, there are some important differences between the two groups of adopters (event adopters and organic adopters) across all three events. For instance, in the case of COVID-19, event/covid adopters tend to be older, have higher incomes, and have been shopping with the retailer for a longer time compared to organic adopters. However, these two groups have similar pre-adoption monthly spending. In the case of Black Friday, event/Black Fridayadopters are younger, have been shopping with the retailer for a longer time, and have lower monthly spend with the retailer in the pre-adoption period relative to organic adopters. In contrast, in the Loyalty program case, event adopters tend to be slightly older, have been shopping with the retailer for a shorter period, and have a higher monthly spend with the retailer in the pre-adoption period compared to organic adopters. However, the differences in characteristics between the two groups of adopters for each event are much smaller in magnitude relative to the differences between adopters and non-adopters (i.e. offline-only customers).
6 Empirical Strategy
As discussed earlier, our goal is to estimate the ATT (Average Treatment Effect on the Treated) of online adoption on post-adoption purchase behavior for different types of events that trigger the adoption of online shopping, and then connect these ATTs estimates with the different adoption pathways. To estimate our ATTs of interest, we use the difference-in-differences approach. This approach allows us to account for unobserved heterogeneity across customers and time-specific shocks. For each event, we use customer-month-level panel data for the estimation sample corresponding to that event, and we utilize the customer-month observations from the pre- and post-adoption periods. We exclude the adoption periods from the estimation for the reasons discussed in $\S$5.1.
Specifically, we use the following DID specification to measure the ATT of online adoption on an outcome variable of interest, $y_{it}$, for customer $i$ in month $t$:
$y_{it}$ $=\alpha_{0}+\alpha_{1}\mathbbm{I}\{Adopter_{i}\}+\alpha_{2}\mathbbm{I}\{Post_{t}\}+\alpha_{3}\mathbbm{I}\{Adopter_{i}\}\times\mathbbm{I}\{Post_{t}\}+\boldsymbol{{\alpha}}_{4}^{\prime}\boldsymbol{{x}}_{it}+\epsilon_{it}.$ (1)
As discussed earlier, we have three main outcome variables of interest: spend, profitability, and share of offline spend. We now discuss the interpretation of the coefficients and variables in the above equation for the case when spend is the outcome variable of interest. The interpretations are similar for the other outcome variables.
[…]
We also use a second specification including customer fixed effects ($\gamma_{i}$) and time (month) fixed effects ($\eta_{t}$) as follows:
$y_{it}$ $=\gamma_{i}+\eta_{t}+\beta_{0}\mathbbm{I}\{Post_{t}\}\times\mathbbm{I}\{Adopter_{i}\}+\boldsymbol{{\beta}}_{1}^{\prime}\boldsymbol{{x}}_{it}+\epsilon_{it}$ (2)
Here, $\beta_{0}$ is the coefficient of interest, representing the ATT estimates.
Finally, we have a specification where we focus on the different types of adopters and use a model that compares organic adopters and event adopters, as follows:
$y_{it}=\delta_{0}+\delta_{1}\mathbbm{1}\{Event\_Adopter_{i}\}+\delta_{2}\mathbbm{1}\{Post_{t}\}+\delta_{3}\mathbbm{1}\{Event\_Adopter_{i}\}\times\mathbbm{1}\{Post_{t}\}+\boldsymbol{{\delta_{4}}}^{\prime}\boldsymbol{{x}}_{it}+\epsilon_{it}$ (3)
$(\delta_{3})$, the coefficient of interest, represents the difference in the change in spend between event adopters and organic adopters in the post-adoption period relative to pre-adoption. We also include a specification, Equation (4) with customer fixed effects ($\gamma_{i}$) and time (month) fixed effects ($\eta_{t}$).
$y_{it}=\gamma_{i}+\eta_{t}+\beta_{2}\mathbbm{1}\{Event\_Adopter_{i}\}\times\mathbbm{1}\{Post_{t}\}+\boldsymbol{{\beta_{3}}}^{\prime}\boldsymbol{{x}}_{it}+\epsilon_{it}.$ (4)
Here $(\beta_{2})$ is the coefficient of interest.
The validity of all the DID models discussed here relies on the parallel trends assumption, which we discuss in detail in $\S$9.2. Further, we discuss in $\S$9.1 how potential contamination between the different types of adopters could affect our results and conclusions. For instance, some adopters who we identify as covid adopters because they adopt during March-April 2020 could actually be organic adopters who would have adopted organically during that time anyway. If this is the case, we show that our measured effect would be an underestimate of the true the difference between event adopters and organic adopters.
7 Results
7.1 Adopters vs. offline-only customers
[…]
The results from the DID models specified in Equations (1) and (2) with offline-only customers as the control group and spend as the outcome variable are presented in Table 3. Across all three events (COVID-19, Black Friday, and Loyalty Program), we find that adopters (pooling organic and event adopters) increase their spending in the post-adoption period (relative to pre-adoption) compared to offline-only customers. To measure effect sizes on a relative scale, we report each ATT as a percentage of pre-adoption monthly spend. The post-adoption monthly spend increases by 30.6% for COVID-19 (i.e., 140.1 / (280.1 + 177.2)), 16.7% for Black Friday (i.e., 78.37 / (291.7 + 178.2)), and 16.6% for the Loyalty Program (i.e., 84.89 / (332.6 + 179.2)). It is interesting to note that the macro-shock (COVID-19) and promotions (Black Friday and Loyalty Program) yield differential increases in spending. We explore this more in the following sub-section.
Table 3: DID Main Analysis – Spend (Adopters vs. Offline-only)
Study: Covid Black Friday Loyalty Program
DV: Spend
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 280.1 (0.557) 291.7 (0.586) 332.6 (0.800)
[0.000] [0.000] [0.000]
Adopter 177.2 (2.885) 177.8 (9.177) 123.4 (4.42)
[0.000] [0.000] [0.000]
Post -62.90 (0.496) -96.95 (0.426) -51.36 (0.632)
[0.000] [0.000] [0.000]
Adopter * Post 140.4 (2.843) 140.1 (2.974) 81.29 (7.543) 78.37 (7.078) 10.02 (5.191) 84.89 (4.238)
[0.000] [0.000] [0.000] [0.000] [0.053] [0.000]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 22,204,608 22,204,608 9,505,999 9,505,999 4,403,449 4,403,449
R2 0.01905 0.38214 0.00980 0.41557 0.00918 0.48127
Within R2 0.00102 $6.84\times 10^{-5}$ 0.00294
Clustered (Customer) standard‐errors are presented in standard brackets, while p‐values are shown in square brackets.
[…]
Table 4: DID Main Analysis - Profit (Adopters vs. Offline-only)
Study: Covid Black Friday Loyalty Program
DV: Profit
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 92.57 (0.177) 96.13 (0.187) 125.8 (0.288)
[0.000] [0.000] [0.000]
Adopter 61.35 (0.925) 60.94 (3.041) 41.57 (1.607)
[0.000] [0.000] [0.000]
Post -10.56 (0.174) -22.83 (0.152) -16.81 (0.238)
[0.000] [0.000] [0.000]
Adopter * Post 35.87 (0.974) 37.37 (1.070) 19.61 (2.547) 18.44 (2.506) -7.231 (1.820) 18.41 (1.506)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 22,204,608 22,204,608 9,505,999 9,505,999 4,403,449 4,403,449
R2 0.01228 0.31903 0.00450 0.34750 0.00651 0.44546
Within R2 0.00047 $2.48\times 10^{-5}$ 0.00193
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
In summary, we see that consumers who adopt online shopping both spend more and are more profitable than offline-only users, across all three events, confirming our prediction H1. This finding is consistent with the earlier literature on multichannel shopping (Neslin and Shankar 2009; Montaguti, Neslin, and Valentini 2016). However, as discussed earlier, both prior work and this analysis treat all adopters as a homogeneous unit, i.e., they ignore channel adoption pathways. As such, it masks any heterogeneity in how channel pathways affect post-adoption spend, profitability, and purchase behavior. In the next section, we present empirical findings on the relationship between channel adoption pathways and post-adoption consumer behavior.
7.2 Event adopters vs. organic adopters
We now examine the differences between each of the different event adopters and organic adopters by estimating the DID models specified in Equations (3) and (4). We will first present the results on spend, followed by the share of offline/online spending, and then look at profitability. These results will help us answer our second research question, and verify the second set of predictions (H2a through H2d) presented in $\S$3.
7.2.1 Spend
The results from the estimation of Equation (3) for spend are in columns (1), (3), and (5) of Table 5 for Covid, Black Friday, and Loyalty Program, respectively. We see that there is no significant difference in the post-adoption spend between covid adopters and organic adopters. However, the incremental spending of Black Friday adopters and Loyalty Program adopters is lower than organic adopters, post-adoption. We also control for customer and time fixed effects as shown in Equation (4) (in columns (2), (4), and (6) of Table 5) and confirm the robustness of these findings to these additional controls.
Table 5: DID Main Analysis – Spend (Event Adopters vs. Organic Adopters)
Study: Covid Black Friday Loyalty Program
DV: Spend
Model: (1) (2) (3) (4) (5) (6)
Variables
Organic_Adopter 432.8 (5.325) 459.1 (10.48) 444.7 (4.415)
[0.000] [0.000] [0.000]
Covid_Adopter 32.84 (6.281)
[0.000]
Post 85.85 (5.34) 10.54 (9.334) -29.92 (5.041)
[0.000] [0.257] [0.000]
Covid_Adopter * Post -10.64 (6.271) -3.095 (6.465)
[0.089] [0.632]
BlackFriday_Adopter 32.01 (20.72)
[0.122]
BlackFriday_Adopter * Post -82.25 (15.47) -69.88 (14.26)
[0.000] [0.000]
Loyalty_Adopter 131.7 (19.08)
[0.000]
Loyalty_Adopter * Post -254.7 (20.82) -53.22 (19.42)
[0.000] [0.006]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 76,842 76,842 197,658 197,658
R2 0.00131 0.39193 0.00081 0.41147 0.05635 0.47126
Within R2 $5.37\times 10^{-7}$ 0.00068 0.02692
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
These findings suggest that the reason why a consumer adopts online shopping affects the post-adoption incremental spending. While all adopters of online shopping increase their spending with the retailer in the post-adoption period, relative to offline-only customers, those who adopt due to a macro-environmental shock (COVID-19, in our case) spend similarly post-adoption, relative to organic adopters. In contrast, those who adopt due to a promotional activity (Black Friday and Loyalty Program) by the firm, spend less than organic adopters. This is an important finding for managers and researchers since it suggests that ignoring channel adoption pathways can significantly misestimate the incremental gains from online adopters. Further, it is consistent with our theoretical prediction (H2a) of forward buying and demand distortion (Neslin, Henderson, and Quelch 1985; Hendel and Nevo 2006; Van Heerde and Neslin 2017) from $\S$3. We know that promotions are an important mechanism for firms to induce various kinds of consumer behavior, including the adoption of online shopping. However, time-limited promotions may induce forward buying, leading to demand distortion and reduced post-promotion consumption.
Theory and Mechanism: Black Friday adopters and Loyalty Program adopters could be increasing their spending during the promotion (i.e., the adoption period) and effectively pulling forward their future demand, resulting in a “post-promotion dip” in spending in the post-adoption period.
To formally examine if this is the case, we investigate the differences between the two groups of event adopters relative to organic adopters during each month around adoption. We estimate the following event study regression:
$y_{it}$ $=\gamma_{i}+\eta_{t}+\sum_{j}\theta_{j}(PRE_{it}(j)\cdot\mathbbm{I}\{Event\_Adopter_{i}\})$ (5)
$+\sum_{k}\theta_{k}(POST_{it}(k)\cdot\mathbbm{I}\{Event\_Adopter_{i}\})+\boldsymbol{\lambda}^{\prime}\boldsymbol{x}_{it}+\epsilon_{it}$
In this model, $PRE_{it}(j)$ is an indicator that equals 1 if period $t$ is $j$ months prior to the month of adoption. The coefficient $\theta_{j}$ for $j=-J,-(J-1),\ldots,-1$ measures the difference in pre-adoption trends between a type of event adopter and organic adopters. These coefficients allow us to test for the validity of the parallel trends assumption, which is discussed in detail in §9.2. Next, $POST_{it}(k)$ is an indicator that equals 1 if period $t$ is $k$ months after online adoption. The coefficient $\theta_{k}$ captures the differences in spend between event adopters and organic adopters during each month following adoption. This event study regression allows us to measure: (1) the differences spending between event adopters and organic adopters during the month of adoption, $\theta_{0}$; and (2) the dynamic post-adoption effects $\theta_{k}$ for $k=1,2,\ldots$, which reflect the spend differences in the subsequent months post-adoption. We compare each group’s spending in a time period to their most recent pre-adoption outcome, i.e., spend in the month immediately prior to adoption.
[…]
Figure 2: Event Study for Forward Buying: Both Black Friday adopters and Loyalty Program adopters spend more than organic adopters during the month of adoption. However, they spend less than organic adopters in the following months
(a) Black Friday
(b) Loyalty Program
Thus, based on the theory of forward buying during promotions and our empirical findings, managers should not expect as high a post-adoption spend for promotion-driven adopters of online shopping relative to organic adopters.
7.2.2 Share of offline/online spend
As discussed in $\S$3, consumer inertia and habit theory (Verplanken and Wood 2006; Wood and Neal 2009)would predict that adoption driven by external shocks (e.g., COVID-19, firm driven promotions) may lead to temporary disruptions in consumer habits (adopting a new channel and multichannel shopping), but without reinforcement, consumers may revert to prior routines (not use the newly adopted channels much). We test this prediction (H2b) in our empirical setting, by comparing post-adoption channel usage between consumers who adopted online shopping organically and those who did so in response to the COVID-19 shock, a short-term one-time firm-driven promotion (Black Friday) and a more long-term and recurring firm driven promotion (Loyalty Program). In other words, we examine if there are differences in how the total spend with the retailer is distributed across the offline and online channels for the different groups of event adopters relative to organic adopters in the post-adoption period.
[…]
The results above represent the averages for the entire post-adoption period and mask any trends. So we also investigate whether there are differential trends in the share of offline spend between each type of event adopter relative to organic adopters in the post-adoption period. In Figure 3, we show the trend lines for share of offline spend for each type of event adopter vs. organic adopters (and the differences).
Figure 3: Share of offline spend relative to total spend – Event Adopters vs. Organic Adopters
(a) Covid
(b) Black Friday
(c) Loyalty Program
[…]
7.2.3 Profitability
Now, we examine the differences in profitability between the different groups of event adopters relative to organic adopters. We know that there are differences in spend and in the share of offline/online spending.
There are also differences in prices and profit margins across the two channels in our setting. To highlight the differences in these three variables across channels, we first generate a panel of prices at the item-channel-month level [^21] For instance, in the case of price, an observation in this panel is the average price of an item $j$ sold through a channel $c$ in time period $t$, and is calculated by the total sales of item $j$ in channel $c$ in month $t$ divided by the total quantity of that item sold in that channel-month.. Next, we calculate the offline-online differences at the item-month level to obtain item-specific offline-online differences in price and profit margin. We then average over months and obtain item-level average differences in price and profit margin between offline and online channel. Table 6 provides item-level summary statistics for the 22,803 items that were sold in both offline and online channels in our data (and were purchased at least once in the same month). For confidentiality, we multiply the profit margin by the same undisclosed multiplier used for the price.
Table 6: Item-Level Offline Price and Profit Margin Relative to Online
mean std 25% 50% 75% (min, max) count
Offline - Online Price 6.01 91.07 -0.84 4.61 17.71 (-7,598.46,1,430.91) 22,803
Offline - Online Profit Margin 5.37 78.49 -1.77 3.47 14.05 (-6,586.62, 2,163.31) 22,803
Thus, on average, prices in physical stores are higher than online prices, and profit margins are higher in physical stores vs. online. Therefore, even for similar levels of spend, consumers who have a higher share of offline spend may be more profitable (our prediction H2d). We present the results from estimating the DID models specified in Equation (3) and Equation (4) for profitability are shown in Table 7.
Table 7: DID Main Analysis - Profit (Event Adopters vs. Organic Adopters)
Study: Covid Black Friday Loyalty Program
DV: Profit
Model: (1) (2) (3) (4) (5) (6)
Variables
Organic_Adopter 145.7 (1.698) 154.7 (3.519) 164.5 (1.627)
[0.000] [0.000] [0.000]
Covid_Adopter 10.96 (2.007)
[0.000]
Post 20.27 (1.788) 3.315 (3.163) -21.05 (1.788)
[0.000] [0.295] [0.000]
Covid_Adopter * Post 7.248 (2.117) 9.079 (2.274)
[0.000] [0.000]
BlackFriday_Adopter 7.338 (6.79)
[0.28]
BlackFriday_Adopter * Post -20.52 (5.228) -16.04 (5.094)
[0.000] [0.002]
Loyalty_Adopter 34.13 (6.413)
[0.000]
Loyalty_Adopter * Post -86.83 (7.352) -15.41 (6.835)
[0.000] [0.024]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 76,842 76,842 197,658 197,658
R2 0.00134 0.29792 0.00039 0.35806 0.04568 0.44335
Within R2 $2.79\times 10^{-5}$ 0.00025 0.02248
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
We see that covid adopters are significantly more profitable than organic adopters post-adoption. The differences are economically meaningful and large in magnitude – the post-adoption gain in profitability is over 30% (i.e., 7.248/20.27) higher for covid adopters relative to organic adopters, consistent with our prediction H2d. The fact that covid adopters are more profitable than organic adopters despite there being no significant differences in spend is due to the differences in prices across channels in our setting and the differences in the share of offline/online spending, as documented in $\S$7.2.2. This shows that the differences in share of spend across channels arising due to consumer habit and inertia have significant economic implications for firms. Also, we find that Black Friday adopters and Loyalty Program adopters are less profitable relative to organic adopters in the post-adoption period by 10.2% (i.e., 16.04 / (154.7 + 3.3)) and 10.7% (i.e., 15.41 / (164.5 - 21.05)), respectively, and this is driven by their lower spending post-adoption as shown in $\S$7.2.1.
To summarize, in our empirical setting, we find that adopters due to the macro environmental shock of COVID-19 are more profitable post-adoption, relative to organic adopters despite similar spending. However, those who adopted due to firm-driven promotions like Black Friday and the Loyalty Program are less profitable. An important managerial implication is that the firm should not naively assume that users who adopt online shopping due to external factors will show the same levels of profitability post-adoption relative to organic adopters.
8 Managerial Implications
Our main findings indicate that consumers who adopt online shopping through different channel pathways (organic, macro-environmental shocks, or promotions) behave differently in the post-adoption period. These patterns are well explained by consumer inertia and forward buying, and should generalize to other settings where similar mechanisms operate. In particular, forward buying and stockpiling can reduce the long-run spending increase from promotion-induced online adopters. To mitigate this and maximize long-run value, firms should design incentives that discourage stockpiling (e.g., phased or behavior-contingent rewards) and combine promotions with rapid follow-up nudges to encourage repeat (rather than one-off) behavior. For consumers pushed online by shocks (like COVID-19), habit-reinforcing promotions and personalized nudges in the immediate post-adoption window can help convert trials of online channels into stable habits, thereby improving long-run profitability. Beyond these general lessons, we derive concrete implications for two common marketing tasks: (a) CLV forecasting and (b) ROI/breakeven analysis for promotions.
8.1 Implications for CLV forecasts
Customer Lifetime Value (CLV) is a core metric for customer relationship management, targeting, and budgeting (Fader, Hardie, and Lee 2005; Gupta et al. 2006; Fader, Hardie, and Shang 2010). Our results show that when offline-only customers adopt online shopping for different reasons (e.g., Black Friday, COVID-19, loyalty program, or organically), their spending and profitability trajectories differ markedly by pathway. Ignoring this heterogeneity leads to biased forecasts of incremental CLV from online adoption and, in turn, to suboptimal decisions, for example, overestimating the effectiveness of promotions aimed at inducing online adoption.
We quantify the average incremental (relative to offline-only customers) CLV from online adoption for each adopter cohort $s$. Following standard practice (Gupta et al. 2006), the (average) incremental CLV equals the present value of all future incremental monthly profits attributable to adoption under an infinite horizon:
$Incremental~CLV_{s}=\sum_{t=1}^{\infty}\frac{ATT_{s}\times r_{s}^{t}}{(1+i)^{t}}=ATT_{s}\frac{r_{s}}{1+i-r_{s}},$ (6)
where: (1) $ATT_{s}$ is the monthly average incremental profit (our ATT estimate) for cohort $s$ relative to offline-only customers; (2) $i$ is the firm’s monthly cost of capital (WACC); and (3) $r_{s}$ is the (average) monthly retention rate of cohort $s$ post-adoption, defined as the probability of purchasing next month conditional on purchasing in the focal month. [^22] We follow Gupta, Lehmann, and Stuart (2004) and use an infinite-horizon formulation to avoid arbitrary truncation; our results are qualitatively similar under a finite horizon. All quantities are monthly. The $ATT_{s}$ estimates come from our DID models for each event, and retention rates are obtained non-parametrically. We report monthly retention rates for organic adopters and event adopters for each event in Table 8. Based on discussions with the retailer, we set WACC to 15% annually, implying a monthly cost of capital of $i=(1+15\%)^{1/12}-1=1.17\%$.
Table 8 presents the incremental CLV from online adoption for organic versus event adopters within each event (COVID-19, Black Friday, Loyalty Program). Covid adopters have an incremental CLV of 73.486 MCU, 36.4% higher than organic adopters (i.e., (73.486 - 53.859) / 53.859). By contrast, promotion-induced adopters are far less valuable in the long run: Black Friday adopters’ incremental CLV is 69.2% lower, and Loyalty Program adopters’ is 78.9% lower than that of organic adopters. These differences mirror our monthly profit findings: covid adopters generate higher long-run value, while promotion-induced adopters (Black Friday, Loyalty Program) generate substantially lower value than organic adopters.
Table 8: Incremental CLV (measured in MCU) of Organic Adopters and Event Adopters – By Three Events (COVID-19, Black Friday, and Loyalty Program)
COVID-19 Black Friday Loyalty Program
Organic Adopters
ATT 30.280 23.570 18.550
Monthly Retention Rate 0.648 0.638 0.559
Incremental CLV 53.859 40.319 22.873
Event Adopters
ATT 39.780 7.813 2.986
Monthly Retention Rate 0.656 0.621 0.625
Incremental CLV 73.486 12.417 4.831
Diff in Incremental CLV of Event Adopters vs. organic adopters +36.4% -69.2% -78.9%
[…]
We next show how CLV forecasts go wrong when managers ignore adoption pathways. We consider two common scenarios:
  • $\bullet$
    Using organic adopters as proxies. Suppose the firm launches a new promotion to encourage online adoption and proxies long-term CLV using organic adopters (because those are the only adopters observed pre-promotion). Table 8 shows that doing so overstates the incremental CLV of promotion-induced adopters by 69.2% for Black Friday and 78.9% for the loyalty program. Using organic adopters as proxies can therefore severely overvalue promotion-induced adopters.
  • $\bullet$
    Pooled forecasts. Alternatively, the firm might pool event and organic adopters and estimate a single incremental CLV for all adopters. The results of this exercise are shown in Table 9.232323For the COVID-19 event, for example, the pooled ATT from Table 4 is 37.30 MCU, and the pooled monthly retention rate is 0.654. Using Equation (6), the incremental CLV is $37.30\times\frac{0.654}{1+0.0117-0.654}=68.332$ MCU. We perform similar calculations for Black Friday and the loyalty program events. Pooled estimates overstate the value of promotion-induced adopters by almost 60% (Black Friday) and 80% (Loyalty Program), while understating the value of COVID-induced adopters by 7.5%.
Table 9: Incremental CLV (measured in MCU) Under Pooled Estimation – By Three Events (COVID-19, Black Friday, and Loyalty Program)
COVID-19 Black Friday Loyalty Program
ATT (Pooled) 37.370 18.440 18.410
Monthly Retention Rate (Pooled) 0.654 0.633 0.565
Incremental CLV (Pooled) 68.332 30.822 23.301
Diff in Incremental CLV of Event Adopters vs. Pooled Estimates +7.5% -59.7% -79.3%
[…]
In both scenarios, failing to distinguish adoption pathways biases CLV forecasts—typically overstating the value of promotion-induced adopters and making weak promotional strategies appear attractive, which can lead to overspending and erode long-run profitability. A key takeaway is that retailers should (a) estimate (incremental) CLV separately by adoption pathway and (b) set pathway-specific caps on the cost of inducing online adoption, aligning offer depth and paid-media spend with those caps.
8.2 Implications for ROI and Breakeven Analyses for Promotions
To assess whether to run a promotional campaign (such as Black Friday online discounts or an app-activated loyalty program) to induce online adoption, firms must compare the expected incremental profits from adopters to the cost of the campaign. If managers naïvely use organic adopters’ incremental profit as a proxy for promotion-induced adopters, they will overstate expected benefits and miscalculate breakeven times and ROI. We illustrate this with the Black Friday and Loyalty Program campaigns.
[…]
In sum, managers must calibrate expectations of incremental spending when running promotions to encourage online adoption. Ignoring adoption pathways can lead to overly optimistic ROI assessments and underestimation of the time required to recoup promotional investments.
9 Robustness Checks
We now present a series of robustness checks to provide additional credibility for our findings.
9.1 Effect size measurement
A potential concern with our ATT measurement, especially for our Covid analysis, is that covid adopters may be “contaminated” with some organic adopters – customers who would have adopted the online channel organically during the onset of COVID-19, had COVID-19 not occurred. As a result, the treatment group may include a mix of “true” covid adopters and organic adopters. If such contamination exists, the ATT estimates reported in §7.2 would be a lower bound on the true ATT for covid-driven adoption (relative to organic), and we can back out the ATT for the subset of customers of interest (i.e., the “true” covid adopters). The true ATT for covid adopters relative to organic adopters can be obtained by dividing the estimated ATT (reported in §7.2) by the fraction of “true” covid adopters, given by Equation 7:
$\textrm{True}~ATT_{covid,organic}=\frac{\textrm{Estimated}~ATT_{covid,organic}}{w_{covid}}$ (7)
[…]
Note that we are less concerned about the contamination problem for our other two events for a few reasons, which we briefly discuss below.
  • $\bullet$
[…]
  • $\bullet$
    Next, recall that the Loyalty Program adopters are customers who specifically went online for the first time to activate the loyalty program on their mobile app and only made their first online purchase after activation, so contamination with organic adopters (those adopted online shopping before the loyalty program was launched and then activated the loyalty program after it was launched) is unlikely.
9.2 Parallel pre-trends
[…]
9.3 Doubly robust DID, Propensity Score Matching and IPW approaches
[…]
9.4 Alternate definitions for the organic adopter and covid adopter cohorts
[…]
These analyses also addresses another potential concern with our main analysis, which is that organic adopters adopt for two extra months relative to covid adopters before our common post-adoption period, and this may drive some of the results. The fact that our conclusions don’t change when we change the set of months used to define our cohorts also allays this concern. Finally, our definitions for Black Friday adopters and Loyalty Program adopters are much more precise than for covid adopters, so we don’t repeat this robustness check for those analyses.
9.5 Alternate operationalizations of dependent variables
[…]
10 Conclusions
This paper examines how the adoption of a retailer’s online channels by previously offline-only customers affects their subsequent behavior, and how this effect varies by the reason for adoption. Building on prior work on multichannel shopping, and the theories of forward buying, and consumer inertia, we argue that (i) adopting online channels and becoming multichannel should increase overall spending relative to remaining offline-only, but (ii) the magnitude, persistence, and channel mix of this increase should differ systematically across adoption pathways. Using transaction data from a large Brazilian pet supplies retailer (2019–2024), we study four adopter groups: covid adopters (adoption during the onset of COVID-19), Black Friday adopters (promotion-driven adoption via deep online discounts), Loyalty Program adopters (adoption triggered by a loyalty program activated on the app), and organic adopters (voluntary adoption without identifiable external stimuli).
Consistent with our predictions, all four adopter types increase their spending relative to offline-only customers. However, their post-adoption trajectories differ markedly. Relative to organic adopters, covid adopters exhibit similar total spend but, in line with forward-buying theory, Black Friday adopters and Loyalty Program adopters spend less in the post-adoption period, having stocked up during the promotion. In terms of channel mix, and consistent with consumer inertia and habit theory, all three externally driven adopter groups have a lower online spend share than organic adopters. Moreover, covid adopters show slower growth in online share, reflecting their older and more habit-driven profile, while younger Black Friday adopters catch up more quickly. Loyalty Program adopters do not differ significantly from organic adopters in the growth of online share, likely because of similar demographics and the monthly online activation requirement, which reinforces online use. Finally, consistent with typical cross-channel margin differences, covid adopters are more profitable than organic adopters due to their higher offline spend share, whereas Black Friday adopters and Loyalty Program adopters are less profitable due to lower overall spending.
These results highlight that the motives and circumstances surrounding online adoption play a crucial role in shaping long-term multichannel behavior, through mechanisms linked to demographics, forward buying, and consumer inertia. Managerially, our findings underscore the need to tailor incentives to the adoption context. Promotions designed to induce online adoption should be structured to limit stockpiling (e.g., via phased or behavior-contingent rewards) and followed by targeted nudges to sustain engagement. For customers who move online due to external shocks, habit-reinforcing incentives can help convert temporary adoption into lasting behavior, improving long-run profitability. We also show that forecasts of incremental Customer Lifetime Value (CLV) and breakeven analyses can be substantially biased if adoption pathways are ignored: using organic adopters as proxies or pooling all adopters can overstate the value of promotion-induced adopters and mislead ROI assessments. Firms should therefore explicitly distinguish organic from externally driven adopters when forecasting CLV and evaluating promotions.
Our study has limitations that suggest directions for future research. First, the external stimuli in our setting are limited to one macro shock (COVID-19) and two specific promotions run by a single retailer. Future work could examine other events (e.g., store closures, different promotion designs, or alternative macro shocks) that may attract different demographic profiles and interact differently with forward buying and inertia. Second, we focus on one retail category (pet supplies) in a single country. Studying similar questions in other categories and contexts would help assess the generalizability of our findings. Finally, we estimate descriptive ATTs of online adoption using a difference-in-differences approach with consumer and time fixed effects, without fully separating self-selection from causal effects of adoption. We do so because the overall ATT, which embeds both selection and causal components, is the quantity most relevant for managerial decisions. Nonetheless, future academic work could complement our analysis by estimating a causal ATE of online adoption, for example using instrumental variables or other identification strategies.
Funding and Competing Interests
The authors were not paid by the data provider. A preliminary version of this manuscript was reviewed by the data provider to ensure that data was accurately depicted, but they did not review/comment on the content of the manuscript.
[…]
Appendix A Data cleaning
Table A1: Item categories with substantially different availability time via online and offline channels ($Y$-$M$-$D$)
Item Category First Offline Purchase Date First Online Purchase Date
Large size food 2020-09-12 2020-02-28
Human food 2019-01-02 2021-02-12
Large size pharmacy 2021-02-03 2020-12-09
Pet Cleaning Supplies 2019-01-03 2022-03-13
There are four product categories with significant discrepancies in their initial availability time on the offline and online channels. For example, the product category Large size food (e.g. pack size greater than 10KG) was available online in February 2020 but only became available offline in September 2020. In addition, the product category Pet Cleaning Supplies was only available offline before COVID-19 but became available online in August 2022.
Including these food categories in our analysis can be problematic since it also reflects retailers’ assortment and the channel-specific effects (i.e. consumers can only buy some products exclusively from one channel). Therefore, we exclude these four categories from our analysis. The fraction of sales of these four categories only accounts for 0.63% of the total revenue during the sample period, and as such unlikely to have any meaningful impact on our findings.
Appendix B Appendix to Summary Statistics
In this section, we summarize the demographics, pre-adoption, and post-adoption purchase behaviors of three customer groups for each focal event: event adopters, organic adopters, and offline-only customers.
To characterize the pre-online-adoption purchase behavior of each group of customers, we compute several consumer-level metrics, such as AvgSpendPerMonth, AvgOrdersPerMonth, AvgUniqueCategoriesPerMonth, using data from the pre-adoption period specific to each case study. We summarize these variables for each group – event adopters, organic adopters, and offline-only – by calculating the group mean as follows: $\frac{1}{N_{g}}\sum_{i=1}^{N_{g}}(\frac{1}{T_{i}^{pre}}\sum_{t=1}^{T_{i}^{pre}}X_{it})$, where $i$ denotes a customer belonging to group $g$, $t$ denotes the year-month index, $T_{i}^{pre}$ denotes the length of the consumer $i$’s panel during the pre-adoption period, and $N_{g}$ is the number of customers in group $g$. We next summarize the post-online adoption behaviors in a similar way using the post-online adoption data under each case study. The post-online adoption means for each group ggevent adopters, organic adopters, and offline-only – are calculated as follows: $\frac{1}{N_{g}}\sum_{i=1}^{N_{g}}(\frac{1}{T_{i}^{post}}\sum_{t=1}^{T_{i}^{post}}X_{it})$, where $i$ denotes a customer belonging to group $g$, $t$ denotes the year-month index, $T_{i}^{post}$ denotes the length of the consumer $i$’s panel during the post-adoption period, and $N_{g}$ is the number of customers in group $g$.
B.1 Event 1: COVID-19
In this section, we present the descriptive statistics for customers in the sample for the COVID-19. We first summarize demographics in Table A2. Next, in Tables A3 and A4, we show customer-level summary statistics of average purchase variables over the pre-adoption and post-adoption period, respectively. Finally, in Table A5, we present more detailed summary statistics (e.g., median and quartiles) for all three groups during the post-online-adoption periods (i.e., between May 1st, 2020 and April 30, 2023). Note that the counts of organic adopters and covid adopters from Table A5 slightly differ from the raw count based on our cohorts’ definition because 469 (3.66% of) organic adopters and 933 (2.81% of) covid adopters churned before 2020-05-01, and therefore are not observed in the post-COVID-19 panel. Including these customers does not affect our identification of post-online-adoption ATT, and the estimation results remain the same.
Table A2: Customer Demographics – Summary Statistics of Gender (Panel A) and Geographic State (Panel B).
Categories Covid Adopter (%) Organic Adopter (%) Offline-only (%)
Panel A: Gender
No Information 12.09 13.52 15.63
Female 59.59 57.36 46.51
Male 28.31 29.13 37.86
Total 100.00 100.00 100.00
Panel B: Geographic State
São Paulo 58.78 49.68 63.70
Minas Gerais 8.58 12.13 4.81
Rio de Janeiro 6.69 6.95 4.30
Distrito Federal 6.21 5.48 4.26
Zip_State_Missing 4.07 4.28 10.34
Santa Catarina 3.33 3.46 2.90
Goiás 2.37 3.36 2.13
Bahia 1.74 3.31 0.69
Rio Grande do Sul 3.21 2.77 2.20
Espírito Santo 1.27 2.43 0.91
Mato Grosso do Sul 0.93 1.88 1.05
Paraná 1.68 1.86 2.31
Other States 1.14 2.41 0.39
Total 100.00 100.00 100.00
Table A3: Customer-Level Summary Statistics of Age and Monthly Household Income (Panel A), Pre-online-adoption Purchase Behaviors between January 1st, 2019 and December 31st, 2019 (Panel B). Columns 1 - 3 describe statistics for Covid adopters, and Organic adopters, respectively. Column 4 calculates the difference between Covid adopters and Organic adopters, and columns 5 and 6 show the two-sample t-test and p-value results. The variable, Tenure, is the number of days between a customer’s first offline purchase in the retailer and the last date of the pre-adoption period of the covid study (i.e., 2019-12-31).
Variable Covid Adopters (Mean) Organic Adopters (Mean) Diff t-stats p-value Offline-only (Mean)
Panel A: Numerical Demographic Variables and Other Pre-online-adoption Variable
Age 43.36 41.44 1.92 15.65 0.00 45.62
Monthly Household Income 2432.11 2200.69 231.41 12.30 0.00 2033.01
Tenure(Days) 245.12 215.63 29.49 23.99 0.00 239.70
Panel B: Pre-online-adoption Monthly Purchase Behaviors
AvgSpendPerMonth 463.46 453.26 10.20 1.80 0.07 284.38
AvgQuantitiesPerMonth 4.98 4.43 0.55 6.15 0.00 2.96
AvgOrdersPerMonth 0.95 0.95 -0.00 -0.38 0.71 0.74
AvgUniqueItemsPerMonth 2.92 2.89 0.03 0.97 0.33 1.91
AvgUniqueBrandsPerMonth 2.32 2.34 -0.02 -0.92 0.36 1.55
AvgUniqueSubcategoriesPerMonth 2.02 2.04 -0.02 -1.10 0.27 1.39
AvgUniqueCategoriesPerMonth 1.65 1.66 -0.01 -0.62 0.54 1.17
AvgMarginPerMonth 162.12 162.36 -0.24 -0.12 0.91 97.52
Table A4: Customer-Level Summary Statistics of Post-online-adoption Purchase Behaviors between May 1st 2020 and April 30th 2023 (Panel A). Columns 1 - 2 describe statistics for Covid adopters and Organic adopters, respectively. Column 3 calculates the difference between Covid adopters and Organic adopters, and columns 4 and 5 show the two-sample t-test and p-value results.
Variable Covid Adopters (Mean) Organic Adopters (Mean) Diff t-stats p-value Offline-only (Mean)
Panel A: Post-online-adoption Monthly Purchase Behaviors
AvgSpendPerMonth 514.47 489.47 25.00 3.75 0.00 202.29
AvgQuantitiesPerMonth 4.88 4.11 0.77 7.58 0.00 1.85
AvgOrdersPerMonth 0.89 0.89 -0.00 -0.08 0.93 0.45
AvgUniqueItemsPerMonth 2.28 2.04 0.23 9.46 0.00 1.13
AvgUniqueBrandsPerMonth 1.78 1.63 0.15 9.18 0.00 0.91
AvgUniqueSubcategoriesPerMonth 1.52 1.39 0.13 10.35 0.00 0.80
AvgUniqueCategoriesPerMonth 1.27 1.16 0.11 10.82 0.00 0.68
AvgMarginPerMonth 175.95 157.42 18.53 8.18 0.00 76.73
Table A5: Customer-Level Summary Statistics for All Three Groups – Covid adopters, Organic Adopters, and Offline-only Customers for post-online-adoption period (i.e., between May 1st, 2020 and April 30th, 2023).
Variables Mean Std. Min 25% 50% 75% Max Count
Covid adopters
AvgSpendPerMonth 514.47 647.83 0.00 155.32 338.44 655.25 40153.38 32,313
AvgQuantitiesPerMonth 4.88 10.11 0.00 0.97 2.18 4.89 414.11 32,313
AvgOrdersPerMonth 0.89 0.81 0.00 0.36 0.69 1.17 17.81 32,313
AvgUniqueItemsPerMonth 2.28 2.41 0.00 0.78 1.57 2.97 58.62 32,313
AvgUniqueBrandsPerMonth 1.78 1.61 0.00 0.67 1.33 2.41 18.61 32,313
AvgUniqueSubcategoriesPerMonth 1.52 1.23 0.00 0.62 1.21 2.11 12.86 32,313
AvgUniqueCategoriesPerMonth 1.27 0.93 0.00 0.56 1.06 1.78 8.81 32,313
AvgMarginPerMonth 175.95 222.05 -95.36 54.56 116.51 224.52 15880.26 32,313
Organic Adopters
AvgSpendPerMonth 489.47 582.43 0.00 141.99 317.95 629.33 10709.68 12,330
AvgQuantitiesPerMonth 4.11 8.10 0.00 0.83 1.89 4.22 245.86 12,330
AvgOrdersPerMonth 0.89 0.83 0.00 0.33 0.67 1.19 11.03 12,330
AvgUniqueItemsPerMonth 2.04 2.15 0.00 0.67 1.39 2.67 30.58 12,330
AvgUniqueBrandsPerMonth 1.63 1.51 0.00 0.58 1.19 2.19 16.31 12,330
AvgUniqueSubcategoriesPerMonth 1.39 1.16 0.00 0.54 1.08 1.92 9.47 12,330
AvgUniqueCategoriesPerMonth 1.16 0.88 0.00 0.50 0.94 1.61 6.81 12,330
AvgMarginPerMonth 157.42 191.05 -24.34 46.69 101.13 203.75 5028.99 12,330
Offline-only Customers
AvgSpendPerMonth 202.29 330.96 0.00 35.98 99.68 242.19 24573.87 573,934
AvgQuantitiesPerMonth 1.85 4.59 0.00 0.30 0.77 1.85 487.17 573,934
AvgOrdersPerMonth 0.45 0.53 0.00 0.12 0.28 0.58 40.11 573,934
AvgUniqueItemsPerMonth 1.13 1.56 0.00 0.25 0.63 1.39 123.22 573,934
AvgUniqueBrandsPerMonth 0.91 1.09 0.00 0.23 0.55 1.17 48.44 573,934
AvgUniqueSubcategoriesPerMonth 0.80 0.87 0.00 0.21 0.50 1.06 19.90 573,934
AvgUniqueCategoriesPerMonth 0.68 0.70 0.00 0.19 0.44 0.94 10.00 573,934
AvgMarginPerMonth 76.73 120.02 -491.73 14.17 38.99 93.21 9834.37 573,934
B.2 Event 2: Black Friday
In this section, we present the descriptive statistics for customers in the sample for the Black Friday event. We first summarize demographics in Table A6. Next, in Tables A7 and A8, we show customer-level summary statistics of average purchase variables over the pre-adoption and post-adoption period, respectively. Finally, in Table A9, we present more detailed summary statistics (e.g., median and quartiles) for all three groups during the post-online-adoption periods (i.e., between December 1st, 2019 and October 31, 2020). Note that the counts of organic adopters and Black Friday adopters from Table A9 slightly differ from the raw count based on our cohorts’ definition because 79 (2.58% of) organic adopters and 24 (1.74% of) Black Friday adopters churned before 2019-12-01, and therefore are not observed in the post-Black Friday panel. Including these customers does not affect our identification of post-online-adoption ATT, and the estimation results remain the same.
Table A6: Customer Demographics – Summary Statistics of Gender (Panel A) and Geographic State (Panel B).
Categories Black Friday Adopter (%) Organic Adopter (%) Offline-only (%)
Panel A: Gender
No Information 14.13 13.59 15.42
Female 59.42 57.89 46.46
Male 26.45 28.52 38.12
Total 100.00 100.00 100.00
Panel B: Geographic State
São Paulo 57.97 51.91 64.83
Minas Gerais 8.48 13.10 4.69
Rio de Janeiro 7.90 6.99 4.50
Distrito Federal 4.78 4.64 4.38
Zip_State_Missing 4.93 3.76 9.90
Santa Catarina 2.54 3.40 2.60
Bahia 2.61 3.20 0.50
Goiás 2.83 3.17 2.19
Rio Grande do Sul 2.17 3.17 2.10
Espírito Santo 1.96 2.52 0.73
Paraná 1.74 1.99 2.39
Mato Grosso do Sul 1.52 1.83 1.10
Other States 0.58 0.33 0.09
Total 100.00 100.00 100.00
Table A7: Customer-Level Summary Statistics of Age and Monthly Household Income (Panel A), Pre-online-adoption Purchase Behaviors between January 1st 2019 and October 31st 2019 (Panel B). Columns 1 - 3 describe statistics for Black Friday adopters, and Organic adopters, respectively. Column 4 calculates the difference between Black Friday adopters and Organic adopters, and columns 5 and 6 show the two-sample t-test and p-value results. The variable, Tenure, is the number of days between a customer’s first offline purchase in the retailer and the last date of the pre-adoption period of the Black Friday study (i.e., 2019-10-31).
Variable Black Friday Adopters (Mean) Organic Adopters (Mean) Diff t-stats p-value Offline-only (Mean)
Panel A: Numerical Demographic Variables and Other Pre-online-adoption Variable
Age 41.86 42.72 -0.86 -2.49 0.01 47.31
Monthly Household Income 2278.60 2254.31 24.29 0.42 0.68 2046.46
Tenure(Days) 203.25 188.93 14.32 4.51 0.00 204.45
Panel B: Pre-online-adoption Monthly Purchase Behaviors
AvgSpendPerMonth 484.56 472.50 12.06 0.67 0.50 291.87
AvgQuantitiesPerMonth 5.00 4.49 0.50 2.07 0.04 3.00
AvgOrdersPerMonth 1.00 0.99 0.02 0.55 0.58 0.76
AvgUniqueItemsPerMonth 3.06 2.95 0.10 0.99 0.32 1.93
AvgUniqueBrandsPerMonth 2.44 2.39 0.05 0.72 0.47 1.57
AvgUniqueSubcategoriesPerMonth 2.13 2.08 0.04 0.74 0.46 1.41
AvgUniqueCategoriesPerMonth 1.72 1.69 0.03 0.77 0.44 1.19
AvgMarginPerMonth 165.90 168.36 -2.46 -0.36 0.72 99.52
Table A8: Customer-Level Summary Statistics of Post-online-adoption Purchase Behaviors between December 1st 2019 and October 31st 2020 (Panel A). Columns 1 - 2 describe statistics for Black Friday adopters and Organic adopters, respectively. Column 3 calculates the difference between Black Friday adopters and Organic adopters, and columns 4 and 5 show the two-sample t-test and p-value results.
Variable Black Friday Adopters (Mean) Organic Adopters (Mean) Diff t-stats p-value Offline-only (Mean)
Panel A: Post-online-adoption Monthly Purchase Behaviors
AvgSpendPerMonth 417.53 468.23 -50.70 -2.94 0.00 195.21
AvgQuantitiesPerMonth 4.37 4.55 -0.17 -0.59 0.56 2.06
AvgOrdersPerMonth 0.88 0.97 -0.10 -3.36 0.00 0.48
AvgUniqueItemsPerMonth 2.28 2.39 -0.11 -1.33 0.18 1.24
AvgUniqueBrandsPerMonth 1.83 1.90 -0.08 -1.25 0.21 0.99
AvgUniqueSubcategoriesPerMonth 1.56 1.63 -0.07 -1.55 0.12 0.87
AvgUniqueCategoriesPerMonth 1.29 1.35 -0.06 -1.66 0.10 0.74
AvgMarginPerMonth 144.17 157.64 -13.47 -2.31 0.02 73.46
Table A9: Customer-Level Summary Statistics for All Three Groups – Black Friday adopters, Organic Adopters, and Offline-only customers for post-online-adoption period (i.e., between December 1st, 2019 and October 31st, 2020).
Variables Mean Std. Min 25% 50% 75% Max Count
Black Friday adopters
AvgSpendPerMonth 417.53 485.46 0.00 118.34 280.31 536.91 6029.03 1,356
AvgQuantitiesPerMonth 4.37 10.24 0.00 0.82 2.00 4.64 219.18 1,356
AvgOrdersPerMonth 0.88 0.82 0.00 0.27 0.64 1.18 6.91 1,356
AvgUniqueItemsPerMonth 2.28 2.51 0.00 0.64 1.50 3.09 20.27 1,356
AvgUniqueBrandsPerMonth 1.83 1.82 0.00 0.55 1.27 2.55 13.82 1,356
AvgUniqueSubcategoriesPerMonth 1.56 1.40 0.00 0.55 1.18 2.18 8.91 1,356
AvgUniqueCategoriesPerMonth 1.29 1.06 0.00 0.45 1.00 1.91 7.00 1,356
AvgMarginPerMonth 144.17 167.61 -56.75 40.64 96.10 184.10 2184.19 1,356
Organic Adopters
AvgSpendPerMonth 468.23 543.57 0.00 127.57 315.86 605.36 6608.86 2,982
AvgQuantitiesPerMonth 4.55 8.55 0.00 0.82 2.09 4.91 162.82 2,982
AvgOrdersPerMonth 0.97 0.93 0.00 0.36 0.73 1.27 7.82 2,982
AvgUniqueItemsPerMonth 2.39 2.66 0.00 0.67 1.64 3.09 27.00 2,982
AvgUniqueBrandsPerMonth 1.90 1.88 0.00 0.64 1.36 2.55 17.82 2,982
AvgUniqueSubcategoriesPerMonth 1.63 1.45 0.00 0.55 1.27 2.27 12.45 2,982
AvgUniqueCategoriesPerMonth 1.35 1.09 0.00 0.55 1.09 1.91 7.73 2,982
AvgMarginPerMonth 157.64 182.93 -105.01 42.74 105.73 203.51 2349.98 2,982
Offline-only Customers
AvgSpendPerMonth 195.21 330.14 0.00 14.31 86.21 243.99 16616.77 534,871
AvgQuantitiesPerMonth 2.06 5.51 0.00 0.18 0.73 2.09 522.36 534,871
AvgOrdersPerMonth 0.48 0.62 0.00 0.09 0.27 0.64 19.73 534,871
AvgUniqueItemsPerMonth 1.24 1.84 0.00 0.18 0.64 1.55 52.91 534,871
AvgUniqueBrandsPerMonth 0.99 1.31 0.00 0.12 0.55 1.36 23.18 534,871
AvgUniqueSubcategoriesPerMonth 0.87 1.06 0.00 0.09 0.55 1.25 16.09 534,871
AvgUniqueCategoriesPerMonth 0.74 0.85 0.00 0.09 0.45 1.09 9.82 534,871
AvgMarginPerMonth 73.46 120.07 -331.15 5.58 33.33 93.25 6313.29 534,871
B.3 Event 3: Loyalty Program
In this section, we present the descriptive statistics for customers in the sample for the loyalty program analysis. We first summarize demographics in Table A10. Next, in Tables A11 and A12, we show customer-level summary statistics of average purchase variables over the pre-adoption and post-adoption period, respectively. Finally, in Table A13, we present more detailed summary statistics (e.g., median and quartiles) for all three groups during the post-online-adoption periods (i.e., between October 1st, 2023 and March 31, 2024)
Table A10: Customer Demographics – Summary Statistics of Gender (Panel A) and Geographic State (Panel B). Our data providers label around 45% of overall customers as no gender information in this study.
Categories Loyalty Program Adopter (%) Organic Adopter (%) Offline-only (%)
Panel A: Gender
No Information 50.16 48.30 42.85
Female 30.57 32.52 31.02
Male 19.27 19.17 26.13
Total 100.00 100.00 100.00
Panel B: Geographic State
São Paulo 35.99 43.29 55.28
Zip_State_Missing 32.99 19.92 21.93
Minas Gerais 5.30 6.69 3.84
Rio de Janeiro 3.89 4.39 3.15
Santa Catarina 6.25 4.14 3.04
Distrito Federal 2.62 3.86 3.09
Espírito Santo 2.23 2.59 1.28
Ceará 2.04 2.37 0.83
Pernambuco 1.08 2.10 0.94
Bahia 1.21 1.99 1.12
Goiás 1.34 1.83 1.63
Mato Grosso do Sul 0.70 1.22 0.95
Other States 4.34 5.62 2.93
Total 100.00 100.00 100.00
Table A11: Customer-Level Summary Statistics of Age and Monthly Household Income (Panel A), Pre-online-adoption Purchase Behaviors between January 1st, 2023 and May 31st, 2023 (Panel B). Columns 1 - 3 describe statistics for Loyalty Program Adopters, and Organic adopters, respectively. Column 4 calculates the difference between Loyalty Program Adopters and Organic adopters, and columns 5 and 6 show the two-sample t-test and p-value results. The variable, Tenure, is the number of days between a customer’s first offline purchase in the retailer and the last date of the pre-adoption period of the loyalty program study (i.e., 2023-05-31).
Variable Loyalty Program Adopters (Mean) Organic Adopters (Mean) Diff t-stats p-value Offline-only (Mean)
Panel A: Numerical Demographic Variables and Other Pre-online-adoption Variable
Age 44.57 43.43 1.14 2.78 0.01 47.01
Monthly Household Income 1959.34 2072.83 -113.49 -2.24 0.02 1975.35
Tenure(Days) 786.58 827.54 -40.96 -3.09 0.00 915.96
Panel B: Pre-online-adoption Monthly Purchase Behaviors
AvgSpendPerMonth 576.42 444.70 131.71 8.49 0.00 332.57
AvgQuantitiesPerMonth 4.31 3.25 1.06 6.44 0.00 2.62
AvgOrdersPerMonth 1.04 0.83 0.20 9.64 0.00 0.70
AvgUniqueItemsPerMonth 2.55 2.05 0.50 7.96 0.00 1.64
AvgUniqueBrandsPerMonth 2.00 1.64 0.36 8.48 0.00 1.31
AvgUniqueSubcategoriesPerMonth 1.70 1.41 0.29 8.76 0.00 1.14
AvgUniqueCategoriesPerMonth 1.42 1.20 0.22 8.57 0.00 0.98
AvgMarginPerMonth 198.59 164.46 34.12 6.04 0.00 125.81
Table A12: Customer-Level Summary Statistics of Post-online-adoption Purchase Behaviors between October 1st, 2023 and March 31st, 2024 (Panel A). Columns 1 - 2 describe statistics for Loyalty Program Adopters and Organic adopters, respectively. Column 3 calculates the difference between Loyalty Program Adopters and Organic adopters, and columns 4 and 5 show the two-sample t-test and p-value results.
Variable Loyalty Program Adopters (Mean) organic adopters (Mean) Diff t-stats p-value Offline-only (Mean)
Panel A: Post-online-adoption Monthly Purchase Behaviors
AvgSpendPerMonth 696.00 524.96 171.04 9.42 0.00 284.45
AvgQuantitiesPerMonth 5.22 3.58 1.64 7.70 0.00 2.18
AvgOrdersPerMonth 1.36 1.02 0.34 12.57 0.00 0.59
AvgUniqueItemsPerMonth 2.75 2.05 0.69 10.81 0.00 1.36
AvgUniqueBrandsPerMonth 2.16 1.65 0.51 11.62 0.00 1.09
AvgUniqueSubcategoriesPerMonth 1.84 1.42 0.41 12.33 0.00 0.96
AvgUniqueCategoriesPerMonth 1.54 1.22 0.32 12.27 0.00 0.83
AvgMarginPerMonth 223.51 180.19 43.32 6.86 0.00 110.25
Table A13: Customer-Level Summary Statistics for All Three Groups – Loyalty Program Adopters, Organic Adopters, and Offline-only Customers for post-online-adoption period (i.e., between October 1st, 2023 and March 31st, 2024).
Variables Mean Std. Min 25% 50% 75% Max Count
Loyalty Program Adopters
AvgSpendPerMonth 696.00 810.94 0.00 210.82 452.64 876.14 9712.41 1,567
AvgQuantitiesPerMonth 5.22 11.23 0.00 1.00 2.67 5.45 262.33 1,567
AvgOrdersPerMonth 1.36 1.31 0.00 0.50 1.00 1.83 20.50 1,567
AvgUniqueItemsPerMonth 2.75 2.89 0.00 1.00 2.00 3.67 43.33 1,567
AvgUniqueBrandsPerMonth 2.16 1.97 0.00 0.83 1.67 2.83 27.67 1,567
AvgUniqueSubcategoriesPerMonth 1.84 1.48 0.00 0.83 1.50 2.67 15.17 1,567
AvgUniqueCategoriesPerMonth 1.54 1.11 0.00 0.67 1.33 2.17 9.00 1,567
AvgMarginPerMonth 223.51 272.37 -44.82 66.43 144.49 280.82 3800.20 1,567
Organic Adopters
AvgSpendPerMonth 524.96 674.20 0.00 132.98 336.48 678.11 21482.62 16,746
AvgQuantitiesPerMonth 3.58 7.69 0.00 0.67 1.67 3.83 358.33 16,746
AvgOrdersPerMonth 1.02 0.99 0.00 0.33 0.83 1.33 15.17 16,746
AvgUniqueItemsPerMonth 2.05 2.38 0.00 0.67 1.33 2.67 41.67 16,746
AvgUniqueBrandsPerMonth 1.65 1.64 0.00 0.50 1.17 2.17 22.33 16,746
AvgUniqueSubcategoriesPerMonth 1.42 1.25 0.00 0.50 1.17 2.00 14.33 16,746
AvgUniqueCategoriesPerMonth 1.22 0.97 0.00 0.50 1.00 1.67 9.00 16,746
AvgMarginPerMonth 180.19 235.65 -1078.75 45.07 114.10 233.52 7693.62 16,746
Offline-only Customers
AvgSpendPerMonth 284.45 484.92 0.00 36.70 146.34 356.93 37203.48 396,396
AvgQuantitiesPerMonth 2.18 5.39 0.00 0.33 1.00 2.17 859.33 396,396
AvgOrdersPerMonth 0.59 0.69 0.00 0.17 0.40 0.83 50.67 396,396
AvgUniqueItemsPerMonth 1.36 1.90 0.00 0.33 0.83 1.80 114.83 396,396
AvgUniqueBrandsPerMonth 1.09 1.29 0.00 0.25 0.67 1.50 56.00 396,396
AvgUniqueSubcategoriesPerMonth 0.96 1.02 0.00 0.20 0.67 1.33 22.67 396,396
AvgUniqueCategoriesPerMonth 0.83 0.82 0.00 0.17 0.67 1.17 10.67 396,396
AvgMarginPerMonth 110.25 180.70 -690.21 14.02 56.76 138.84 11025.70 396,396
B.4 Loyalty Program Control Variables
Below, we provide the definition of time-varying control variables related to customers’ participation in the loyalty program. These variables ($\boldsymbol{x}_{it}$) are included in the DID and event study analysis for the loyalty program study.
  • $\bullet$
    $isActivated_{it}$: 1 if customer i activates the offer during month $t$ and 0 otherwise. Use to compare within-activators, how much spend differs between months with/without activation.
  • $\bullet$
    $isFirstActivation_{it}$: A binary variable equal to 1 if customer $i$ activates the offer for the first time in month $t$, and 0 otherwise. This variable isolates the novelty (first-time) in the month of initial activation, relative to subsequent activation months.
  • $\bullet$
    $Months\_Since\_First\_Activation_{it}$: A non-negative integer indicating the number of months elapsed since customer $i$’s first activation of the offer, as of month $t$; equals 0 for months prior to the first activation. For example, if the first activation occurs in August 2023, then
    $Months\_Since\_First\_Activation_{i,September\_2023}=1$, and so on.
  • $\bullet$
    $\%Month\_WithActivation_{it}$: The proportion of months in which customer $i$ has activated the offer up to (but not including) month $t$, calculated as the total number of activation months up to $t$ divided by the number of months since first activation (excluding $t$ itself). For the first month of activation, this variable is defined as 0 (since the denominator is 0); all first-month effects are captured by $isFirstActivation_{it}$. Starting from the second month post-activation, this variable reflects the historical activation rate prior to month $t$.
Appendix C Appendix to the Profitability Analysis: Missing Value Imputation
As discussed in the main text, we have data on prices, margins, and costs for each item sold since January 2020; however, we lack cost data for 2019 due to issues with the data provider’s upgrade to a new data management platform in 2020. Thus, we need to impute the missing costs and margins for the 2019 data. We do this in two steps.
Step 1: We impute the cost of each item in 2019 via an inflation-adjusting approach by using the unit cost and the inflation data. Based on conversations with our retail collaborator, we learned that procurement contracts are typically updated on an annual basis and that unit cost increases for the majority of items are largely due to inflation (or in sync with inflation rates). Hence, we calculate the average cost for each item $j$ in each channel $c$ (i.e., online and offline) and each year (i.e., 2020, 2021, and 2022), denoted as $cost_{jc,2020},cost_{jc,2021}$, $cost_{jc,2022}$.  [^32] We focus on cost value over a 36-month period between 2020 and 2022. We do not consider the 2023 cost value as the basis for cost imputation, as we do not have complete cost data for the full year of 2023 and cannot compute the yearly average cost.
We choose Brazil’s Extended National Consumer Price Index (IPCA) in June (the midpoint of the year) as the average level of IPCA for each year, and obtain a set of unit cost estimates for each item $j$ in channel $c$, denoted as $s_{jc}$, by adjusting the ratio of IPCA index in Equation A1.333333We obtain Brazil’s Broad National Consumer Index (IPCA) from the Brazil Institute of Geography and Statistics (IBGE). The IPCA data is available under the Historic Serie section of the following URL:https://www.ibge.gov.br/en/statistics/economic/prices-and-costs.
$s_{jc}=\{cost_{jc,2020}\times\frac{ICPA_{2019}}{ICPA_{2020}},cost_{jc,2021}\times\frac{ICPA_{2019}}{ICPA_{2021}},cost_{jc,2022}\times\frac{ICPA_{2019}}{ICPA_{2022}}\}$ (A1)
We can then take the average of the elements of set $s_{jc}$ for each item-channel combination and calculate the imputed cost of item $j$ in channel $c$ in 2019, as $cost_{jc,2019}$.
Step 2: Since we observe the price, we can then calculate the margin for each item-channel combination in 2019 as
$margin_{jc}=price_{jc}\times(1-AvgTaxRate)-cost_{jc,2019}.$ (A2)
Here, we impute missing margin using an average sales tax rate of 16% for imputing the missing margin. This average sales tax rate, 16% (or one minus average sales tax rate) is obtained by calculating the retailer’s total monthly margin plus cost divided by total monthly sales over a 36-month period between 2020-01-01 and 2022-12-31. This resulting 16% closely matches Brazil’s average sales tax rate of 17% Trading Economics (2025), and our retail collaborator has confirmed that this method is appropriate for imputing missing margin. Finally, we aggregate these margins at the customer-month level to get the customer-month panel data for the profitability measures in 2019.
Appendix D Appendix to Mechanisms, Theoretical Underpinnings and Managerial Implications
D.1 Appendix to the Share of Offline-Online Spend
Table A14: DID Analysis - Offline Spend (Adopters vs. Offline-only)
Study: Covid Black Friday Loyalty Program
DV: Offline Spend
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 280.1 (0.557) 291.7 (0.586) 332.6 (0.800)
[0.000] [0.000] [0.000]
Adopter 177.2 (2.885) 177.8 (9.177) 123.4 (4.420)
[0.000] [0.000] [0.000]
Post -62.90 (0.495) -96.95 (0.426) -51.36 (0.632)
[0.000] [0.000] [0.000]
Adopter_Post -114.3 (2.442) -110.5 (2.553) -116.7 (7.217) -118.0 (6.661) -189.8 (4.585) -116.9 (3.827)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 22,204,608 22,204,608 9,505,999 9,505,999 4,403,449 4,403,449
R2 0.00501 0.36735 0.00899 0.41478 0.00480 0.47768
Within R2 0.00068 0.00016 0.00165
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A15: DID Main Analysis of Offline Spend – Event Adopters vs. Organic Adopters
Study: Covid Black Friday Loyalty Program
DV: Offline Spend
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 432.8 (5.325) 459.1 (10.48) 444.7 (4.415)
[0.000] [0.000] [0.000]
Covid_Adopter 32.84 (6.281)
[0.000]
Post -199.9 (4.462) -205.8 (8.770) -229.8 (4.490)
[0.000] [0.000] [0.000]
Covid_Adopter * Post 32.24 (5.283) 36.13 (5.391)
[0.000] [0.000]
BlackFriday_Adopter 32.01 (20.72)
[0.122]
BlackFriday_Adopter * Post -23.82 (15.26) -7.116 (13.96)
[0.118] [0.611]
Loyalty_Adopter 131.7 (19.08)
[0.000]
Loyalty_Adopter * Post -210.9 (19.02) -12.09 (18.59)
[0.000] [0.515]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 76,842 76,842 197,658 197,658
R2 0.01177 0.36501 0.02459 0.40155 0.05133 0.43555
Within R2 0.00012 $9.15\times 10^{-6}$ 0.01910
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A16: DID Main Analysis of Share of Offline Spend – Event Adopters vs. Organic Adopters
Study: Covid Black Friday Loyalty Program
DV: Share of Offline Spend
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 1.000 ($2.28\times 10^{-15}$) 1.000 ($1.69\times 10^{-16}$) 1.000 ($3.29\times 10^{-17}$)
[0.000] [0.000] [0.000]
Covid_Adopter 0.000 (0.000)
[0.000]
Post -0.5109 (0.0034) -0.4332 (0.0073) -0.4531 (0.0039)
[0.000] [0.000] [0.000]
Covid_Adopter * Post 0.0811 (0.0040) 0.0698 (0.0042)
[0.000] [0.000]
BlackFriday_Adopter 0.000 (0.000)
[0.000]
BlackFriday_Adopter * Post 0.0977 (0.0124) 0.0922 (0.0120)
[0.000] [0.000]
Loyalty_Adopter 0.000 (0.000)
[0.000]
Loyalty_Adopter * Post 0.0002 (0.0115) 0.0018 (0.0114)
[0.987] [0.889]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 949,448 949,448 42,299 42,299 110,339 110,339
R2 0.18119 0.50739 0.25515 0.53685 0.28417 0.57059
Within R2 0.00140 0.00528 0.00904
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A17: Differential Trends in the Share of Offline Spend (Estimation Results for Equation (A3)) – Event Adopters vs. Organic Adopters
DV: Share of Offline Spend
Model: Covid Black Friday Loyalty Program
Variables
Organic_Adopter 0.4833 (0.0084) 0.5930 (0.0285) 0.5662 (0.0093)
[0.000] [0.000] [0.000]
Event_Adopter 0.0807 (0.0119) 0.1424 (0.0404) 0.0306 (0.0132)
[0.000] [0.000] [0.020]
t -0.0020 (0.0004) -0.0108 (0.0048) -0.0020 (0.0031)
[0.000] [0.025] [0.518]
Event_Adopter * t 0.0012 (0.0006) -0.0109 (0.0068) 0.0013 (0.0043)
[0.046] [0.109] [0.763]
Fit statistics
Observations 72 22 12
R2 0.81922 0.69953 0.72620
Adjusted R2 0.81124 0.64945 0.62353
Standard errors in parentheses; p‑values in square brackets.
Table A18: Differential Trend in the Share of Offline Spend by Age – Covid Adopters vs. Organic Adopters
DV: Share of Offline Spend
Model: (1) Senior Adopters (2) Young Adopters
Variables
Organic_Adopter 0.4947 (0.0092) 0.4728 (0.0081)
[0.000] [0.000]
Covid_Adopter 0.0706 (0.0130) 0.0882 (0.0114)
[0.000] [0.000]
t -0.0019 (0.0005) -0.0022 (0.0004)
[0.001] [0.000]
Covid_Adopter * t 0.0020 (0.0006) 0.0004 (0.0006)
[0.002] [0.523]
Fit statistics
Observations 72 72
R2 0.79964 0.82429
Adjusted R2 0.79080 0.81654
Standard errors are shown in parentheses; p-values are shown in square brackets.
Table A14 presents the DID results of offline spend between adopters and offline-only customers. Table A15 and A16 shows the DID results of offline spend and share of offline spend between event adopters and organic adopters. Table A17 presents the estimation results for the linear trend model by comparing a type of event adopters with organic adopters. Table A18 presents the changes in share of offline spend by different age groups, showing senior adopters tend to have a slower rate of shifting their spend to the online channel.
In §7.2.2, we find that the share of offline spending is persistently higher for covid adopters compared to organic adopters in the post-adoption period, and the downward trend is steeper for organic adopters, indicating that covid adopters shift their spending to online channels slower than organic adopters. To formally test whether these differential trends in share of offline spend are systematic, we estimate the following model with linear time trend on the post-adoption monthly panel data for both groups:
$y_{gt}=\zeta+\rho_{0}t+\rho_{1}\mathbbm{I}\{Event\_Adopter_{g}\}+\rho_{2}\mathbbm{I}\{Event\_Adopter_{g}\}\times t+\epsilon_{gt}.$ (A3)
Here, $y_{gt}$ is the fraction of the offline spend relative to the total spend (ranging from 0 to 1) for group $g$ in period $t$. $t$ is the numerical time index, and we set the first post-adoption month of each study as $t=0$. The results for covid, Black Friday and loyalty program studies from this regression are shown in Table A17. The estimate of the difference in trend ($\rho_{2}$) between covid and organic adoption cohorts is 0.0012 (or 0.12%). This finding suggests that although the different reasons (i.e., organic vs. exogenous shock) for adopting online shopping do not seem to have much impact on how much consumers spend with the firm in total, they do cause differences in where they spend it. Figure A1 plots the fraction of offline spending in the post-treatment period separately for older and younger adopters.
Figure A1: Differential Trend Visualization by Age – Covid Adopters vs. Organic Adopters
(a) Senior Adopters (Age $\geq 41$)
(b) Young Adopters (Age <41<41)
D.2 Appendix to the Other Metrics
Finally, to further explore the differences in the purchase behavior of different groups of adopters relative to organic adopters post-online-adoption, we estimate the same DID model (as shown in Equation (3) and (4)) on other metrics of interest (Quantity, Number of orders, Unique items, Brands, Categories, and Subcategories). For each study (Covid, Black Friday, and Loyalty Program), we present results for both total and offline metrics. Specifically, total metrics are reported in Tables A19, A21, and A23, while offline metrics appear in Tables A20, A22, and A24, respectively.
To begin with, we find some differences between covid and organic groups. Covid adopters tend to place fewer orders but purchase a wider variety of items, brands, and categories, compared to organic adopters. This is consistent with their heavier use of the offline channel since consumers who use offline channels generally tend to buy more frequently and also buy a wider variety of products (Chintala, Liaukonytė, and Yang 2023). Next, we find Black Friday adopters consistently exhibit lower purchase overall across all total metrics than organic adopters, but show insignificant differences in offline metrics. This suggests, short-term online promotional event are insufficient to raise adopters’ purchase levels in online channel to those observed among organic adopters. Lastly, we do not observe significant differences in total and offline purchase metrics between Loyalty Program adopters and organic adopters, after controlling for loyalty program’s participation. This suggests that when adopters’ motivation for online adoption is to access long-term discounts (available for both online and offline), their purchase variety and channel utilization remain similar to organic adopters. However, they may be leveraging the loyalty program to maintain similar consumption levels while spending less, thus potentially becoming less profitable to the firm.
Appendix to Additional Covid Analysis
Tables A19 and A20 present DID results for the total and offline metrics under the covid analysis.
Table A19: Covid adopters vs. Organic Adopters Results – Other Total Metrics
DVs: Quantity Orders Unique Items Unique Brands Unique Subcategories Unique Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Covid_Adopter * Post 0.026 (0.099) -0.029 (0.010) 0.039 (0.028) 0.038 (0.020) 0.039 (0.016) 0.032 (0.012)
[0.792] [0.004] [0.163] [0.057] [0.015] [0.008]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949
R2 0.49396 0.37110 0.36977 0.33648 0.31704 0.30595
Within R2 $2.22\times 10^{-7}$ $2.38\times 10^{-5}$ $4.7\times 10^{-6}$ $8.66\times 10^{-6}$ $1.42\times 10^{-5}$ $1.54\times 10^{-5}$
Clustered (Customer) standard‐errors in parentheses; p‑values in square brackets.
Table A20: Covid adopters vs. Organic Adopters Results – Other Offline Metrics
DVs: Quantity Orders Unique Items Unique Brands Unique Subcategories Unique Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Covid_Adopter * Post -0.028 (0.082) 0.031 (0.008) 0.090 (0.027) 0.087 (0.019) 0.074 (0.015) 0.059 (0.012)
[0.627] [0.000] [0.000] [0.000] [0.000] [0.000]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949
R2 0.45211 0.36623 0.34965 0.32551 0.31257 0.30808
Within R2 $4.03\times 10^{-7}$ $4.13\times 10^{-5}$ $2.94\times 10^{-5}$ $5.17\times 10^{-5}$ $5.74\times 10^{-5}$ $6.07\times 10^{-5}$
Clustered (Customer) standard‐errors in parentheses; p‑values in square brackets.
Appendix to Additional Black Friday Analysis
Tables A21 and A22 present the DID estimates of other total metrics and offline metrics of interest under the Black Friday analysis.
Table A21: Black Friday Adopters vs. Organic Adopters Results - Other Total Metrics
DVs: Quantity Orders Unique Items Unique Brands Unique Subcategories Unique Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
BlackFriday_Adopter * Post -0.653 (0.219) -0.114 (0.025) -0.232 (0.076) -0.140 (0.056) -0.125 (0.044) -0.099 (0.033)
[0.002] [0.000] [0.003] [0.014] [0.005] [0.004]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 76,842 76,842 76,842 76,842 76,842 76,842
R2 0.52977 0.42114 0.41307 0.39483 0.37340 0.35736
Within R2 0.00035 0.00070 0.00029 0.00019 0.00023 0.00025
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A22: Black Friday Adopters vs. Organic Adopters Results - Other Offline Metrics
DVs: Quantity Orders Unique Items Unique Brands Unique Subcategories Unique Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
BlackFriday_Adopter * Post -0.211 (0.205) -0.001 (0.022) -0.039 (0.076) 0.013 (0.056) 0.011 (0.044) 0.017 (0.033)
[0.308] [0.967] [0.610] [0.812] [0.804] [0.591]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 76,842 76,842 76,842 76,842 76,842 76,842
R2 0.51966 0.44015 0.40718 0.39489 0.37993 0.37063
Within R2 $5.29\times 10^{-5}$ $1.13\times 10^{-7}$ $9.13\times 10^{-6}$ $1.93\times 10^{-6}$ $1.92\times 10^{-6}$ $8.1\times 10^{-6}$
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Appendix to Additional Loyalty Program Analysis
Tables A23 and A24 present the DID estimates of the other total metrics and offline metrics of interest under the loyalty program analysis.
Table A23: Loyalty Program Adopters vs. Organic Adopters Results - Other Total Metrics
DVs: Quantity Orders Unique Items Unique Brands Unique Subcategories Unique Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Loyalty_Adopter * Post -0.048 (0.201) -0.047 (0.032) -0.071 (0.077) -0.061 (0.053) -0.040 (0.041) -0.026 (0.031)
[0.811] [0.142] [0.356] [0.250] [0.330] [0.402]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 197,658 197,658 197,658 197,658 197,658 197,658
R2 0.56438 0.44676 0.44267 0.40084 0.37005 0.35296
Within R2 0.01429 0.03068 0.02502 0.02647 0.02701 0.02610
Clustered (Customer) standard‐errors are presented in standard brackets; p‑values are shown in square brackets.
Table A24: Loyalty Program Adopters vs. Organic Adopters Results - Other Offline Metrics
DVs: Quantity Orders Unique Items Unique Brands Unique Subcategories Unique Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Loyalty_Adopter * Post 0.141 (0.206) 0.018 (0.026) 0.090 (0.075) 0.087 (0.051) 0.074 (0.039) 0.059 (0.030)
[0.492] [0.488] [0.708] [0.752] [0.608] [0.406]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 197,658 197,658 197,658 197,658 197,658 197,658
R2 0.52637 0.44029 0.43107 0.39543 0.36797 0.35526
Within R2 0.01064 0.02050 0.01867 0.02049 0.02178 0.02205
Clustered (Customer) standard‐errors are presented in standard brackets; p‑values are shown in square brackets.
D.3 Appendix to Forward buying
Now, we discuss how to set up the event study regression for testing the forward buying under Black Friday and loyalty program analysis. For the Black Friday analysis, we define event time 0 as November 2019, given that Black Friday occurred on 2019-11-29. The pre-adoption period is from 2019-01-01 to 2019-10-31 (event time -10 to -1), and the post-adoption period is from 2019-12-01 to 2020-10-31 (event time 1 to 11). By estimating the event study regression in Equation (5), we present the estimation result in Table A25 and show that Black Friday adopters spend significantly higher than organic adopters during the adoption month, but spend significantly lower during the post-adoption periods, which indicates a pattern of forward buying.
Table A25: Event Study for Spend – Black Friday Adopters vs. Organic Adopters
DV: Spend
Model: (1)
Variables
BlackFriday_Adopter $\times$ Lag –10 –39.54 (50.73)
[0.436]
BlackFriday_Adopter $\times$ Lag –9 –44.26 (38.91)
[0.255]
BlackFriday_Adopter $\times$ Lag –8 –26.81 (35.61)
[0.452]
BlackFriday_Adopter $\times$ Lag –7 –25.93 (30.61)
[0.397]
BlackFriday_Adopter $\times$ Lag –6 5.206 (34.45)
[0.880]
BlackFriday_Adopter $\times$ Lag –5 –4.364 (27.55)
[0.874]
BlackFriday_Adopter $\times$ Lag –4 –18.65 (27.25)
[0.494]
BlackFriday_Adopter $\times$ Lag –3 –9.695 (25.51)
[0.704]
BlackFriday_Adopter $\times$ Lag –2 –38.55 (23.94)
[0.107]
BlackFriday_Adopter $\times$ Adoption Month 128.6 (32.46)
[0.000]
BlackFriday_Adopter $\times$ Lead 1 –134.3 (25.36)
[0.000]
BlackFriday_Adopter $\times$ Lead 2 –114.4 (25.65)
[0.000]
BlackFriday_Adopter $\times$ Lead 3 –105.5 (25.58)
[0.000]
BlackFriday_Adopter $\times$ Lead 4 –88.21 (28.63)
[0.002]
BlackFriday_Adopter $\times$ Lead 5 –55.82 (28.28)
[0.048]
BlackFriday_Adopter $\times$ Lead 6 –126.9 (27.95)
[0.000]
BlackFriday_Adopter $\times$ Lead 7 –88.91 (28.21)
[0.002]
BlackFriday_Adopter $\times$ Lead 8 –69.56 (27.24)
[0.011]
BlackFriday_Adopter $\times$ Lead 9 –25.79 (28.13)
[0.359]
BlackFriday_Adopter $\times$ Lead 10 –90.03 (28.42)
[0.002]
BlackFriday_Adopter $\times$ Lead 11 –67.00 (28.53)
[0.019]
Fixed-effects
Customer Yes
YearMonth Yes
Fit statistics
Observations 81,283
R2 0.42086
Within R2 0.00211
Clustered (Customer) standard-errors are presented in standard brackets,
while p‑values are shown in square brackets.
Regarding the loyalty program analysis, for both Loyalty Program adopters and organic adopters, the month of adoption is labeled as event time 0. For example, event time 0 is June 2023 for organic adopters who adopted online shopping in that month, and September 2023 for Loyalty Program adopters who adopted online shopping in that month.
We compare each group’s spending at event time 0 to their most recent pre-adoption outcome, specifically those observed in May 2023. This period serves as a common baseline for pre-treatment behaviors across all adopter groups. As defined in §5.2.3, the post-adoption periods span from October 2023 and March 2024, and we label October 2023 as the event time 1 for both organic and Loyalty Program adopters and label all the following months until March 2024 as event time 6. The event study coefficient at time 1 captures the difference in the monthly spend at October 2023 between two groups relative to the pre-adoption difference in the reference period May 2023.
We first compare the spend of organic and Loyalty Program adopters during their adoption month relative to the pre-adoption reference period (i.e., spend in May 2023), and Table A26 shows that Loyalty Program adopters tend to spend 177.7 MCU more than organic adopters during the adoption month. Next, we notice that both Loyalty Program adopters and organic adopters can activate the monthly loyalty offer during the post-adoption period. To test the post-adoption differences in purchase behavior, we estimate the event study regression in Equation (5) with control variables for the customers’ activations. Table A27 shows that Loyalty Program adopters spend significantly less than organic adopters in most of the post-adoption periods.
Table A26: Event Study for Spend during the Adoption Month - Loyalty Program Adopters vs. Organic Adopters
DV: Spend
Model: (1)
Variables
Loyalty_Adopter $\times$ Lag -5 -11.43 (26.52)
[0.672]
Loyalty_Adopter $\times$ Lag -4 -29.64 (25.37)
[0.245]
Loyalty_Adopter $\times$ Lag -3 0.7290 (24.62)
[0.976]
Loyalty_Adopter $\times$ Lag -2 -25.12 (26.40)
[0.340]
Loyalty_Adopter $\times$ Adoption Month 177.7 (31.54)
[0.000]
Fixed-effects
Customer Yes
Event_Time Yes
Fit statistics
Observations 109,857
R2 0.53391
Within R2 0.00111
Clustered (Customer) standard-errors are presented in standard brackets,
while p‑values are shown in square brackets.
Table A27: Event Study for Spend during the Post-Online-Adoption Periods – Loyalty Program Adopters vs. Organic Adopters
DV: Spend
Model: (1)
Variables
Loyalty_Adopter $\times$ Lag -5 -11.45 (26.52)
[0.670]
Loyalty_Adopter $\times$ Lag -4 -29.64 (25.37)
[0.245]
Loyalty_Adopter $\times$ Lag -3 0.7526 (24.62)
[0.975]
Loyalty_Adopter $\times$ Lag -2 -25.10 (26.40)
[0.340]
Loyalty_Adopter $\times$ Lead 1 -102.5 (29.56)
[0.001]
Loyalty_Adopter $\times$ Lead 2 -34.48 (28.49)
[0.224]
Loyalty_Adopter $\times$ Lead 3 -45.26 (31.24)
[0.144]
Loyalty_Adopter $\times$ Lead 4 -73.56 (30.45)
[0.015]
Loyalty_Adopter $\times$ Lead 5 -69.94 (33.10)
[0.037]
Loyalty_Adopter $\times$ Lead 6 -75.00 (35.29)
[0.032]
Loyalty program controls Yes
Fixed-effects
Customer Yes
YearMonth Yes
Fit statistics
Observations 197,658
R2 0.47130
Within R2 0.02699
Clustered (Customer) standard-errors are presented in standard brackets,
while p‑values are shown in square brackets.
Next, we run the event study regression for offline spend, and Figure A2 presents the event study figure for the offline spend of Black Friday and loyalty program study.
Finally, we follow a similar approach described above to test for covid adopters. Figure A3 shows that there is no forward buying for covid adopters relative to organic adopters.
Figure A2: Event Study for Offline Spend
(a) Black Friday
(b) Loyalty Program
Figure A3: Event Study for Spend – Covid Adopters vs. Organic Adopters during the month of adoption
D.4 Appendix to Managerial Implications
Table A28 presents an additional DID result for profit, which are inputs for the discussion of managerial implications. Instead of estimating Equation (1) and (2), with a single indicator $\mathbbm{I}\{Adopter_{i}\}$, we now include two separate dummies – $\mathbbm{I}\{Event\_Adopter_{i}\}$ and $\mathbbm{I}\{Organic\_Adopter_{i}\}$ – each interacted with the post-adoption indicator. This allows us to estimate the ATT of online adoption on profit for organic adopters and event adopters, relative to offline-only customers.
Table A28: DID Results (All Three Groups) – Profit
Study: Covid Black Friday Loyalty Program
Dependent Var.: Profit
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 92.57 (0.177) 96.13 (0.187) 125.8 (0.288)
[0.000] [0.000] [0.000]
Organic_Adopter 53.18 (1.707) 58.55 (3.523) 38.65 (1.652)
[0.000] [0.000] [0.000]
Covid_Adopter 64.14 (1.085)
[0.000]
Post -10.56 (0.174) -22.83 (0.152) -16.81 (0.238)
[0.000] [0.000] [0.000]
Organic_Adopter * Post 30.83 (1.796) 30.28 (1.903) 26.14 (3.166) 23.57 (3.160) -4.244 (1.803) 18.55 (1.509)
[0.000] [0.000] [0.000] [0.000] [0.019] [0.000]
Covid_Adopter * Post 38.08 (1.147) 39.78 (1.272)
[0.000] [0.000]
BlackFriday_Adopter 65.89 (5.809)
[0.000]
BlackFriday_Adopter * Post 5.621 (4.164) 7.813 (4.014)
[0.177] [0.051]
Loyalty_Adopter 72.78 (6.210)
[0.000]
Loyalty_Adopter * Post -91.07 (7.814) 2.986 (6.798)
[0.000] [0.292]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 22,204,608 22,204,608 9,505,999 9,505,999 4,403,449 4,403,449
R2 0.01237 0.31903 0.00451 0.34750 0.00660 0.44546
Within R2 0.00047 $2.88\times 10^{-5}$ 0.00194
Clustered (Customer) standard‐errors are presented in standard brackets; p‑values are shown in square brackets.
Additional Analysis for ROI/breakeven analyses for promotion:
[…]
When the cost per customer is 50 MCU, if managers naively used the incremental profit from organic adopters, they would compute a breakeven time of 2.70 months (i.e., 50/18.55), whereas the correct breakeven period using Loyalty Program adopters’ incremental profit is 16.74 months (i.e., 50/2.986) – more than six times longer. Therefore, managers should appropriately calibrate their expectations of increased consumer spending when running promotions that encourage the adoption of online shopping.
When the cost per customer is 150 MCU, if managers naively used the incremental profit from organic adopters, they would compute a breakeven time of 8.09 months (i.e., 150/18.55), whereas the correct breakeven period using Loyalty Program adopters’ incremental profit is 50.23 months (i.e., 150/2.986) – more than six times longer. Therefore, managers should appropriately calibrate their expectations for increased consumer spending when running promotions that encourage online shopping adoption.
Appendix E Appendix to Validity and Robustness Checks
E.1 Appendix to Effect Size Measurement
Figure A4 presents the bootstrapped distribution of true ATT for covid adopters relative to organic adopters. Figure A5 presents the bootstrapped distribution of true ATT for Black Friday adopters relative to organic adopters.
Figure A4: True ATT – Covid Adopters vs. Organic Adopters
(a) Spend
(b) Profit
Figure A5: True ATT – Black Friday Adopters vs. Organic Adopters
(a) Spend
(b) Profit
E.2 Appendix to Parallel Trend Tests
E.2.1 Parallel Trend Test for Covid Analysis - Total Spend
Figure A6: Parallel Trend Assumption - Total Spend
(a) Visualization
(b) Adopters vs. Non-Adopters
(c) Covid Adopters vs. organic adopters
Figure A6 provides parallel trend visualization and test for total spend. Figure 6(a) visualizes the average monthly spend of all three groups in both the pre- and post-treatment periods. For each group, this average monthly spend is calculated by dividing the total spend of customers belonging to the group in a month by the number of customers who have not churned by the end of that month. As we can see, the spending of all three groups evolves parallelly during the pre-adoption period from 2019-01-01 to 2019-12-31, and there are no differential trends in spend across the three groups. [^34] Each line in the parallel trend plots shows the average monthly spending for a given group of customers (organic adopters, covid adopters, and offline-only) over the observation period. At any given period $t$, the averaging is done over the set of active consumers in that period. Both at the beginning and end of the observation period, we have a (relatively) small number of customers in all three groups because of customer acquisition and churn: (1) Not all customers in each group initiate their first purchase at the beginning of the time series (e.g., Jan and Feb 2019), i.e., many customers joined after early 2019, and (2) Customers in all three groups continue to churn in the post-treatment period, and many are no longer in our panel towards the end of the observation period (e.g., March and April 2023). Further, the consumers who were with the firm from early 2019 and/or those who did not churn by mid-2023 tend to be loyal consumers (who spend more on average). Since the early and late observations largely consist of such consumers, we observe higher average spending in these time periods. However, the existence of such spikes does not invalidate the parallel trends observed in the pre-treatment period.
Then, Figure 6(b) presents the pre-trend between adopters (both covid and organic adopters) and offline-only customers (i.e., Non-Adopters), which shows some violation of parallel pre-trend. We use the honest DID approach from Rambachan and Roth (2023) and provide bounded ATT estimates for the total spend in Web Appendix §E.3. The incremental spend due to online adoption is still significant and robust, even in the presence of a large violation of parallel trend.
Finally, Figure 6(c) presents the event study estimated coefficients between organic adopters and covid adopters. As we can see, the estimated coefficients are not significantly different from zero, i.e., the parallel pre-trend assumption seems valid.  [^35] For the parallel trend assumption to hold, it is permissible for a small number of pre-treatment coefficients to be statistically significant (provided the majority of these coefficients are statistically insignificant), see Gopalan, Rosinger, and Ahn (2020) and Chapter 5 of Angrist and Pischke (2008). This suggests that the DID estimation of treatment effects was not falsely inflated by any trends before the treatment. As a further robustness check, we also provide bounded ATT estimates for the total spend result between organic adopters and covid adopters in Web Appendix §E.3.
E.2.2 Parallel Trend Tests for Covid Analysis - Profitability Analysis
Figure A7 presents the parallel trend plot for the profit. Figure 7(a) visualizes the trends. Figure 7(b) presents the pre-trend between adopters (both covid and organic adopters) and offline-only customers (i.e., Non-Adopters), which shows some violation of parallel pre-trend. We use the honest DID approach from Rambachan and Roth (2023) and provide bounded ATT estimates for profit in Web Appendix §E.3. The incremental profit due to online adoption is still significant and robust, even in the presence of a large violation of parallel trend.
Finally, Figure 7(c) confirms that the parallel trend assumption for total profit between covid adopters and organic adopters is satisfied since almost all pre-treatment coefficients are insignificant and not statistically different from 0. As a further robustness check, we also provide bounded ATT estimates for the profit result between organic adopters and covid adopters in Web Appendix §E.3.
Figure A7: Parallel Trend Assumption - Profit
(a) Visualization
(b) Adopters vs. Non-Adopters
(c) Covid Adopters vs. Organic Adopters
E.2.3 Additional Parallel Trend Plots for Covid Analysis – Other Metrics
In this section, we validate the parallel trend assumption for DID results in other metrics between organic adopters and covid adopters, presented in Table A19 and A20. Because both organic and covid adopters only purchase offline during the pre-adoption period, the pre-adoption values for total and offline metrics are exactly the same. So here we test for total metrics and same conclusions apply to offline metrics as well. In Figure A8, we present the parallel trend plots for the metrics shown in Table A19. Figure A9 shows the event study estimates test for pre-trend. We can see some significant pre-treatment coefficients for these metrics, but the magnitude of these coefficients are small.
Figure A8: Monthly Average Metrics – Covid, Organic adopters and Offline-only Customers (i.e., Non-Adopters)
(a) Quantity
(b) Orders
(c) Unique Items
(d) Unique Brands
(e) Unique Subcategories
(f) Unique Categories
Figure A9: Parallel Trend for Other Metrics - Covid Adopters vs. Organic Adopters
(a) Quantity
(b) Orders
(c) Unique Items
(d) Unique Brands
(e) Unique Subcategories
(f) Unique Categories
E.2.4 Parallel Trend for Black Friday Analysis
First, we provide a pre-trend test for two main metrics (i.e., spend and profit) between adopters (including both Black Friday adopters and organic adopters) relative to offline-only customers. In total, we have 10 pre-treatment periods (i.e., between January 2019 and October 2019), labeled as -10 to -1, and 11 post-adoption periods (i.e., between December 2019 and October 2020), labeled as 1 to 11. The base period is October 2019, and the event time 0 refers to the adoption month (i.e., November 2019). We exhibit a minor violation during the early stage of the pre-adoption period, but most pre-adoption coefficients are significant. To show the robustness of our main finding in the presence of violation of parallel pre-trend, we show the bounded ATT in Web Appendix §E.3 and confirm that the incremental spend and profit are still positive and significant in the presence of violation of parallel trend.
Figure A10: Event Study – Adopters (pooled by Organic and Black Friday) vs. Offline-only
(a) Spend
(b) Profit
Next, we provide a pre-trend test between Black Friday adopters and organic adopters. Figure A11 visualizes the monthly average metric values for Black Friday adopters, organic adopters and offline-only customers. Figure A12 presents the event study regression estimates for the parallel pre-trend of the Black Friday study.
Figure A11: Monthly Average Metrics - Black Friday, Organic adopters, and Offline-only Customers (i.e., Non-Adopters)
(a) Spend
(b) Profit
(c) Quantity
(d) Orders
(e) Unique Items
(f) Unique Brands
(g) Unique Subcategories
(h) Unique Categories
Figure A12: Parallel Trend for Black Friday Analysis - Total Metrics (Black Friday Adopters vs. Organic Adopters). We present the estimates of event study regression. The pretreatment outcomes of all offline metrics are the same as the total metrics, since all customers only shop offline.
(a) Spend
(b) Profit
(c) Quantity
(d) Orders
(e) Unique Items
(f) Unique Brands
(g) Unique Subcategories
(h) Unique Categories
E.2.5 Parallel Trend for Loyalty Program Analysis
First, we provide a pre-trend test for two main metrics (i.e., spend and profit) between adopters (including both Loyalty Program adopters and Organic adopters) relative to offline-only customers (see Figure A13). We exhibit a violation of the parallel trend assumption. To show the robustness of our main finding in the presence of violation of parallel pre-trend, we show the bounded ATT in Web Appendix §E.3 and confirm that the incremental spend is still positive and significant in the presence of violation of parallel trend.
Figure A13: Event Study – Adopters (pooled by Organic and Loyalty Program) vs. Offline-only
(a) Spend
(b) Profit
Next, we provide a pre-trend test between Loyalty Program adopters and organic adopters. Figure A14 visualizes the monthly average metric values for Loyalty Program adopters, organic adopters, and offline-only customers. Figure A15 presents the event study regression estimates for the parallel pre-trend of the loyalty program online activation study.
Figure A14: Monthly Average Metrics - Loyalty Program, Organic adopters, and Offline-only Customers (i.e., Non-Adopters).
(a) Spend
(b) Profit
(c) Quantity
(d) Orders
(e) Unique Items
(f) Unique Brands
(g) Unique Subcategories
(h) Unique Categories
Figure A15: Parallel Trend for Loyalty Program Analysis - Total Metrics. We present the estimates of event study regression for Loyalty Program Adopters. The pretreatment outcomes of all offline metrics are the same as the total metrics, since all customers only shop offline.
(a) Spend
(b) Profit
(c) Quantity
(d) Orders
(e) Unique Items
(f) Unique Brands
(g) Unique Subcategories
(h) Unique Categories
E.3 Honest DID
In E.2, we see that when comparing adopters and offline-only customers, there are some violations of the parallel trend assumption. There are also some concerns about the parallel trend assumption between covid adopters and organic adopters. To ensure that estimated ATTs of online adoption are robust to the violations in the parallel trends assumption, we perform the Honest Difference-in-Differences (Honest DID) estimation to quantify the extent to which violations in parallel trends can nullify the treatment effect (Rambachan and Roth 2023).
To conduct this sensitivity analysis, we impose the restriction that the change in slope of the differential trend between treatment and control customers across two periods is bounded by:
$\Delta^{SD}:=\{\zeta:|(\zeta_{t+1}-\zeta_{t})-(\zeta_{t}-\zeta_{t-1})|\leq M,~\forall t\}$ (A4)
Here, $\zeta_{t}$ is the difference in trends between two groups at time $t$, and $M$ is the magnitude of maximum allowable change in trend between consecutive periods.
First, we estimate a linear trend model, including all treatment and control customers using all customer-month observations during the pre-adoption periods:
$y_{it}=\gamma_{i}+\eta_{t}+\phi_{e,v}\mathbbm{I}\{Treatment\}_{i}\times t+\epsilon_{it}$ (A5)
Here, $\mathbbm{I}\{Treatment\}_{i}$ refers to the treatment customers. When we compare adopters with offline-only customers, $\mathbbm{I}\{Treatment\}_{i}=1$ is customer $i$ is an adopter. When we compare covid adopters with organic adopters, $\mathbbm{I}\{Treatment\}_{i}=1$ if customer $i$ is a covid adopter. $t$ is the year-month for the pre-adoption period, and $\phi_{e,v}$ is the linear trend coefficient for event $e$ variable $v$. We test a wide range of $M$ values, up to and beyond $\phi_{e,v}$, for spend and profit under each event, and assess to what extent our ATT will be nullified.
We now present the Honest DID estimates for three events.
E.3.1 COVID-19
Table A29: Linear Trend – Adopters (Organic and Covid) vs. Offline-only
DVs: Spend Profit
Model: (1) (2)
Variables
Adopter * t 4.170 (0.377) 1.128 (0.141)
[0.000] [0.000]
Fixed-effects
Customer Yes Yes
YearMonth Yes Yes
Fit statistics
Observations 5,146,359 5,146,359
R2 0.48244 0.39923
Within R2 $6.28\times 10^{-5}$ $3.05\times 10^{-5}$
Clustered (Customer) standard-errors are presented in standard brackets,
while p‑values are shown in square brackets.
Figure A16: Honest DID Estimation - Adopters (Organic and Covid) vs. Offline-only
(a) Total Spend
(b) Total Profit
We first present the Honest DID estimates with varying $M$ values for two main metrics, monthly spend and profit between adopters (i.e., including both organic and covid adopters) and offline-only customers. Figure 16(a) presents the honest DID 95% confidence intervals when we allow for a violation of parallel trends between consecutive periods up to $M$. We test a wide range of $M$ values and conclude that even if the post-treatment violation of parallel trend is twice as large as the pre-treatment trend (i.e., estimated coefficient from the linear trend model in Table A29), our ATT estimates and main effects still hold: adopters spend more and are more profitable than offline-only customers during the post-adoption periods.
Table A30: Linear Trend – Covid Adopters vs. Organic Adopters
DVs Spend Profit
Model: (1) (2)
Variables
Covid_Adopter * t 4.393 (0.8605) 1.415 (0.3183)
[0.000] [0.000]
Fixed-effects
Customer Yes Yes
YearMonth Yes Yes
Fit statistics
Observations 378,387 378,387
R2 0.51305 0.42830
Within R2 0.00012 $7.71\times 10^{-5}$
Clustered (Customer) standard-errors are presented in standard brackets,
while p‑values are shown in square brackets.
Figure A17: Honest DID Estimation - Covid Adopters vs. Organic Adopters
(a) Total Spend
(b) Total Profit
Next, we present the Honest DID estimates between organic adopters and covid adopters. Figure A17 presents the honest DID 95% confidence intervals when we allow for a violation of parallel trends between consecutive periods up to $M$. We test a wide range of $M$ values and conclude that even if the post-treatment violation of parallel trend is twice as large as the pre-treatment trend (i.e., estimated from the linear trend model in Table A30), our ATT estimates and main effects still hold for spend: covid adopters have a similar incremental spend as the organic adopters post-online-adoption.
For the profit, we demonstrate that the main finding remains valid in the presence of a linear pre-trend (i.e., a significant 95% confidence interval when $M=0$ and a similar magnitude of confidence interval as the original effect). This also confirms the robustness of the result, indicating covid adopters are more profitable than organic adopters in the post-adoption period. However, the effect becomes weaker when we assume a more severe post-treatment violation of the parallel trend.
E.3.2 Black Friday
Table A31: Linear Trend – Adopters (Organic and Black Friday) vs. Offline-only
DVs Spend Profit
Model: (1) (2)
Variables
Adopter * t 4.244 (1.571) 1.081 (0.6015)
[0.007] [0.072]
Fixed-effects
Customer Yes Yes
YearMonth Yes Yes
Fit statistics
Observations 3,855,908 3,855,908
R2 0.48754 0.40597
Within R2 $5.24\times 10^{-6}$ $2.25\times 10^{-6}$
Clustered (Customer) standard-errors are presented in standard brackets,
while p‑values are shown in square brackets.
Figure A18: Honest DID Estimation - Adopters (Organic and Black Friday) vs. Offline-only
(a) Total Spend
(b) Total Profit
We now present the Honest DID estimates between Adopters (i.e., including both organic and Black Friday) and Offline-only customers. Figure 18(a) presents the honest DID 95% confidence intervals when we allow for a violation of parallel trends between consecutive periods up to $M$. We test a wide range of $M$ values and conclude that even if the post-treatment violation of parallel trend is twice as large as the pre-treatment trend (i.e., estimated from the linear trend model in Table A31), our ATT estimates and main effects for spend still hold: adopters spend more than offline-only customers during the post-adoption periods. For the profit, we do not see a significant linear pre-trend (see column 2 of Table A31). Also, we demonstrate that the main finding remains valid in the presence of a linear pre-trend and some deviation (i.e., a magnitude of 0.4) from the linear pre-trend (i.e., 1.081 from Table A31). This also confirms the robustness of the result, indicating adopters are more profitable than offline-only customers after the online adoption.
E.3.3 Loyalty Program
Table A32: Linear Trend – Adopters (Organic and Loyalty Program) vs. Offline-only
DVs: Spend Margin
Model: (1) (2)
Variables
Adopter * t 17.51 (1.504) 6.811 (0.578)
[0.000] [0.000]
Fixed-effects
Customer Yes Yes
YearMonth Yes Yes
Fit statistics
Observations 2,073,524 2,073,524
R2 0.55397 0.51055
Within R2 0.00012 0.00012
Clustered (Customer) standard-errors are presented in standard brackets,
while p‑values are shown in square brackets.
Figure A19: Honest DID Estimation – Adopters (Organic and Loyalty Program) vs. Offline-only
(a) Total Spend
(b) Total Profit
We present the Honest DID estimates between adopters (including organic and loyalty program adopters) and offline-only customers. Figure 19(a) presents the honest DID 95% confidence intervals when we allow for a violation of parallel trends between consecutive periods up to $M$. We test a wide range of $M$ values and conclude that even if the post-treatment violation of parallel trend is twice as large as the pre-treatment trend (i.e., estimated from the linear trend model in Table A32), our ATT estimates and main effects still hold: adopters spend more than offline-only customers after the online adoption. However, under this robustness check, the incremental profit associated with adopters (pooled by organic adopters and Loyalty Program adopters) is not significantly greater than that of offline-only customers during the post-adoption period, even when we assume a more severe post-treatment violation of the parallel trend.
E.4 Appendix to DID Models with Propensity Score Matching and Inverse Propensity Score Weighting
In this section, we provide a set of robustness checks under PSM and IPTW. To conduct this analysis, we label the event adopters as 1 and organic adopters as 0 and model the propensity score of customer $i$ being an event adopter. The propensity score model is specified as follows:
$PropensityScore_{i}=f(X_{i}^{T}\theta),$ (A6)
where $f(\cdot)$ is the logit link function where $f(X_{i}^{T}\theta)=\frac{1}{1+\exp(-X_{i}^{T}\theta)}$. $X_{i}\in\mathcal{R}^{M\times 1}$ is a column vector with $M$ covariates of customer $i$. These covariates include $\textit{Gender}_{i},\textit{Age}_{i},\textit{Tenure(Days)}_{i}$, $\textit{AverageSpendPerMonth}_{i}$, etc. $\theta\in\mathcal{R}^{M\times 1}$ is a column vector of coefficients. Table A33 in Web Appendix E.4.1 presents the propensity score model estimation for covid, Black Friday, and loyalty program analysis, respectively. Figure A20 presents the distribution of propensity scores before matching, and we observe some differences between event adopters and organic adopters. After matching, Figure A21 suggests that event adopters and organic adopters become more overlapped.
Next, we use these propensity scores to augment our DID analysis in two ways.
  • $\bullet$
    First, we perform a customer-level propensity score matching and then re-run the DID analysis. Matching based on propensity score is much more efficient than matching based on a set of covariates (Rosenbaum and Rubin 1985). Based on the predicted propensity score of each customer $i$, we use MatchIt package in R (Stuart et al. 2011) to perform nearest-neighbor matching for each event adopter. We then run the DID analysis on the newly matched sample (see Table A35, A36, and A37 in Web Appendix E.4.2). We find that the results from this exercise are consistent with the main findings.
  • $\bullet$
    Next, we use a slightly different approach to control for any potential selection issues. Specifically, we use an inverse propensity of treatment weight-adjusted (IPTW) DID regression. Here, the sample is the same as the one from the main analysis, but each observation is inversely weighted by its propensity score. We then perform a two-sample t-test on all the pre-treatment variables based on the weighted sample and confirm that the two samples are statistically insignificant on all the variables (see Table A34 for the weighted sample of organic adopters and event adopters, respectively in the Web Appendix E.4.3). The DID results of event adopters and organic adopters from this analysis are shown in Table A38, A39 and A40 for covid, Black Friday, and Loyalty Program analysis, respectively in Web Appendix E.4.3. Again, we find that the results and conclusions are quite similar to those in the main section.
E.4.1 Propensity Score Model Estimation
Table A33: Propensity Score Logit Model – Event Adopters vs. Organic Adopters
DVs:
Covid_Adopter BlackFriday_Adopter LoyaltyProgram_Adopter
Gender_NoInformation -0.044 (0.027) 0.095 (0.088) -0.031 (0.042)
[0.103] [0.278] [0.459]
Gender_Female 0.052 (0.018) 0.098 (0.062) -0.050 (0.035)
[0.004] [0.114] [0.152]
LogTenureDays 0.243 (0.009) 0.175 (0.031) -0.087 (0.026)
[0.000] [0.000] [0.001]
AvgSpendPerMonth -0.0001 (0.0001) 0.0003 (0.0002) 0.002 (0.0001)
[0.317] [0.133] [0.000]
AvgQuantityPerMonth 0.010 (0.002) 0.008 (0.006) 0.0001 (0.003)
[0.000] [0.183] [0.974]
AvgUniqueOrdersPerMonth -0.147 (0.018) -0.121 (0.063) 0.174 (0.028)
[0.000] [0.055] [0.000]
AvgUniqueItemsPerMonth 0.036 (0.013) 0.022 (0.043) -0.066 (0.021)
[0.006] [0.610] [0.002]
AvgUniqueBrandsPerMonth -0.086 (0.022) -0.031 (0.070) -0.0001 (0.042)
[0.000] [0.659] [0.998]
AvgUniqueSubcategoriesPerMonth 0.003 (0.025) 0.081 (0.080) 0.315 (0.055)
[0.905] [0.311] [0.000]
AvgUniqueCategoriesPerMonth 0.088 (0.026) -0.040 (0.089) -0.096 (0.059)
[0.001] [0.655] [0.103]
AvgUniqueVisitedStoresPerMonth 0.289 (0.041) 0.284 (0.137) -0.074 (0.067)
[0.000] [0.038] [0.272]
AvgMarginPerMonth 0.0002 (0.0001) -0.001 (0.0005) -0.006 (0.0003)
[0.045] [0.045] [0.000]
Constant -1.063 (0.410) -1.191 (0.203) -0.079 (0.189)
[0.010] [0.000] [0.672]
Age Yes Yes Yes
Household Income Yes Yes Yes
Observations 46,045 4,441 18,313
Log Likelihood -48,790.450 -3,983.461 -19,285.020
Akaike Inf. Crit. 97,622.890 8,006.922 38,610.040
Standard‐errors in standard brackets; p‑values in square brackets.
  • Note: AvgUniqueVisitedStoresPerMonth measures the average number of distinct stores a customer visits per month during the pre-adoption period under each study.
Table A33 presents the propensity score model estimates for covid adopters, Black Friday adopters, and Loyalty Program adopters versus organic adopters. Figure A20 presents the distribution of propensity scores of event adopters and organic adopters under three studies. After performing nearest-neighbor matching, A21 shows that the distribution of propensity scores between event adopters and organic adopters becomes more overlapped.
Figure A20: Propensity Score Model Predictions - Before Matching
(a) Covid
(b) Black Friday
(c) Loyalty Program
Figure A21: Propensity Score Model Predictions - After Matching
(a) Covid
(b) Black Friday
(c) Loyalty Program
Table A34: Customer-level Summary Statistics of Pre-Adoption Variables for Event Adopters and Organic Adopters: Re-weighted Sample Based on Inverse Propensity Score. Columns 2 and 3 show the weighted mean of variables of event adopters and organic adopters, with weights being the inverse of propensity.
Variable Event Adopters mean Organic adopters mean Diff T-stats P-value
Panel A: Covid Analysis
LogTenure(Days) 5.12 5.12 -0.00 -0.40 0.69
AvgSpendPerMonth 459.22 461.68 -2.47 -0.42 0.67
AvgQuantitiesPerMonth 4.72 4.91 -0.19 -1.67 0.09
AvgOrdersPerMonth 0.95 0.95 0.00 0.04 0.97
AvgUniqueItemsPerMonth 2.90 2.90 0.00 0.01 1.00
AvgUniqueBrandsPerMonth 2.32 2.33 -0.00 -0.04 0.97
AvgUniqueSubcategoriesPerMonth 2.03 2.03 -0.00 -0.10 0.92
AvgUniqueCategoriesPerMonth 1.65 1.65 -0.00 -0.09 0.93
AvgMarginPerMonth 162.25 162.72 -0.47 -0.22 0.83
Panel B: Black Friday Analysis
LogTenure(Days) 5.01 5.01 0.00 0.02 0.99
AvgSpendPerMonth 477.58 477.41 0.17 0.01 0.99
AvgQuantitiesPerMonth 4.75 4.70 0.05 0.21 0.83
AvgOrdersPerMonth 1.00 1.00 0.00 0.01 1.00
AvgUniqueItemsPerMonth 3.00 3.00 0.00 0.01 0.99
AvgUniqueBrandsPerMonth 2.42 2.42 -0.00 -0.00 1.00
AvgUniqueSubcategoriesPerMonth 2.10 2.10 -0.00 -0.01 0.99
AvgUniqueCategoriesPerMonth 1.70 1.70 -0.00 -0.00 1.00
AvgMarginPerMonth 166.97 167.08 -0.11 -0.02 0.99
Panel C: Loyalty Program Analysis
LogTenure(Days) 6.45 6.45 0.00 0.14 0.89
AvgSpendPerMonth 514.74 535.10 -20.36 -1.38 0.17
AvgQuantitiesPerMonth 3.73 3.85 -0.12 -0.70 0.48
AvgOrdersPerMonth 0.93 0.94 -0.01 -0.42 0.67
AvgUniqueItemsPerMonth 2.28 2.29 -0.02 -0.27 0.78
AvgUniqueBrandsPerMonth 1.80 1.81 -0.01 -0.37 0.71
AvgUniqueSubcategoriesPerMonth 1.54 1.55 -0.01 -0.39 0.70
AvgUniqueCategoriesPerMonth 1.30 1.31 -0.01 -0.38 0.70
AvgMarginPerMonth 180.94 184.54 -3.60 -0.70 0.49
E.4.2 DID Results with PSM
Table A35, A36 and A37 present the DID estimates of Equation (3) of key purchase metrics based on the matched sample under covid, Black Friday, and loyalty program analysis. All results are consistent with the main findings.
Table A35: Covid Adopters vs. Organic Adopters – PSM + DID
DVs: Spend Offline Spend Share of Offline Spend Profit
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 438.1 (5.946) 438.1 (5.946) 1.000 (1.94e-15) 147.4 (1.901)
[0.000] [0.000] [0.000] [0.000]
Covid_Adopter 27.61 (6.816) 27.61 (6.816) 0.000 (0.000) 9.298 (2.181)
[0.000] [0.000] [0.000] [0.000]
Post 84.05 (5.916) -202.1 (4.983) -0.5091 (0.0038) 20.28 (1.994)
[0.000] [0.000] [0.000] [0.000]
Covid_Adopter * Post -8.837 (6.768) -1.842 (6.910) 34.42 (5.729) 38.08 (5.813) 0.0793 (0.0044) 0.0693 (0.0046) 7.237 (2.294) 9.258 (2.423)
[0.192] [0.790] [0.000] [0.000] [0.000] [0.000] [0.002] [0.000]
Fixed‐effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 1,689,232 1,689,232 1,689,232 1,689,232 901,177 901,177 1,689,232 1,689,232
R2 0.00125 0.39081 0.01127 0.36431 0.17947 0.50579 0.00122 0.29503
Within R2 1.72e-07 0.00012 0.00125 2.59e-05
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A36: Black Friday Adopters vs. Organic Adopters – PSM + DID
DVs: Spend Offline Spend Share of Offline Spend Profit
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 455.4 (17.04) 455.4 (17.04) 1.000 (8.42e-17) 149.7 (5.392)
[0.000] [0.000] [0.000] [0.000]
BlackFriday_Adopter 35.72 (24.69) 35.72 (24.69) 0.000 (0.000) 12.36 (7.925)
[0.148] [0.148] [0.000] [0.124]
Post 11.79 (14.44) -201.8 (13.40) -0.4275 (0.0127) 7.025 (4.835)
[0.414] [0.000] [0.000] [0.147]
BlackFriday_Adopter * Post -83.50 (18.99) -78.26 (17.94) -27.88 (18.31) -16.68 (17.37) 0.0921 (0.0162) 0.0959 (0.0159) -24.23 (6.380) -22.46 (6.431)
[0.000] [0.000] [0.128] [0.337] [0.000] [0.000] [0.000] [0.001]
Fixed‐effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 42,329 42,329 42,329 42,329 23,221 23,221 42,329 42,329
R2 0.00137 0.41012 0.02553 0.41414 0.23976 0.52090 0.00065 0.34764
Within R2 0.00099 5.74e-05 0.00671 0.00058
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A37: Loyalty Program Adopters vs. Organic Adopters – PSM + DID
DVs: Spend Offline Spend Share of Offline Spend Profit
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 464.3 (5.119) 464.3 (5.119) 1.000 (2.44e-18) 170.0 (1.864)
[0.000] [0.000] [0.000] [0.000]
Loyalty_Adopter 112.1 (19.25) 112.1 (19.25) 0.000 (0.000) 28.59 (6.478)
[0.000] [0.000] [0.000] [0.000]
Post -41.09 (5.787) -245.4 (5.191) -0.4546 (0.0044) -24.27 (2.034)
[0.000] [0.000] [0.000] [0.000]
Loyalty_Adopter * Post -244.3 (22.01) -52.62 (20.60) -198.8 (19.86) -8.541 (19.41) -0.0009 (0.0119) 0.0013 (0.0117) -83.07 (7.797) -16.32 (7.228)
[0.000] [0.011] [0.000] [0.660] [0.940] [0.912] [0.000] [0.024]
Loyalty program controls Yes Yes Yes Yes Yes Yes Yes Yes
Fixed‐effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 159,333 159,333 159,333 159,333 90,778 90,778 159,333 159,333
R2 0.05562 0.47539 0.05221 0.44214 0.28560 0.56883 0.04528 0.44853
Within R2 0.02722 0.01915 0.00922 0.02278
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
E.4.3 DID Results with IPTW
Table A38, A39 and A40 present the DID estimates of Equation (3) of key purchase metrics with inverse propensity score weighting under covid, Black Friday, and loyalty program analysis. All results are consistent with the main findings.
Table A38: Covid Adopters vs. Organic Adopters - IPTW + DID
DVs Spend Offline Spend Share of Offline Spend Profit
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 436.7 (5.445) 436.7 (5.445) 1.000 ($1.68\times 10^{-15}$) 146.7 (1.731)
[0.000] [0.000] [0.000] [0.000]
Covid_Adopter 15.72 (6.310) 15.72 (6.310) 0.000 (0.000) 5.898 (2.016)
[0.014] [0.014] [0.000] [0.004]
Post 89.00 (5.416) -198.3 (4.544) -0.5071 (0.0035) 22.29 (1.833)
[0.000] [0.000] [0.000] [0.000]
Covid_Adopter * Post -13.12 (6.310) -6.305 (6.550) 34.22 (5.316) 37.96 (5.465) 0.0731 (0.0041) 0.0651 (0.0043) 4.667 (2.144) 6.651 (2.307)
[0.040] [0.315] [0.000] [0.000] [0.000] [0.000] [0.030] [0.008]
Fixed-effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 1,783,949 1,783,949 949,448 949,448 1,783,949 1,783,949
R2 0.00126 0.38696 0.01100 0.36110 0.18052 0.50768 0.00105 0.29041
Within R2 $2.31\times 10^{-6}$ 0.00014 0.00125 $1.53\times 10^{-5}$
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A39: Black Friday Adopters vs. Organic Adopters - IPTW + DID
DVs: Spend Offline Spend Share of Offline Spend Profit
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 471.5 (11.07) 471.5 (11.07) 1.000 ($1.37\times 10^{-15}$) 156.7 (3.548)
[0.000] [0.000] [0.000] [0.000]
BlackFriday_Adopter 5.388 (19.85) 5.388 (19.85) 0.000 (0.000) 3.074 (6.679)
[0.784] [0.784] [0.000] [0.646]
Post 5.951 (9.359) -210.2 (8.970) -0.4277 (0.0074) 4.135 (3.139)
[0.532] [0.000] [0.000] [0.186]
BlackFriday_Adopter * Post -74.02 (15.36) -68.89 (14.30) -16.17 (15.14) -4.429 (14.12) 0.0845 (0.0127) 0.0848 (0.0122) -22.85 (5.303) -20.64 (5.231)
[0.000] [0.000] [0.288] [0.765] [0.000] [0.000] [0.000] [0.000]
Fixed-effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 76,842 76,842 76,842 76,842 42,299 42,299 76,842 76,842
R2 0.00123 0.40586 0.02482 0.39667 0.25280 0.53426 0.00074 0.34948
Within R2 0.00067 $3.58\times 10^{-6}$ 0.00455 0.00042
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A40: Loyalty Program Adopters vs. Organic Adopters - IPTW + DID
DVs Spend Offline Spend Share of Offline Spend Margin
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 445.6 (4.422) 445.6 (4.422) 1.000 ($1.64\times 10^{-14}$) 164.7 (1.627)
[0.000] [0.000] [0.000] [0.000]
Loyalty_Adopter 130.8 (19.08) 130.8 (19.08) 0.000 (0.000) 33.90 (6.413)
[0.000] [0.000] [0.000] [0.000]
Post -30.58 (5.047) -230.5 (4.497) -0.4530 (0.0039) -21.23 (1.788)
[0.000] [0.000] [0.000] [0.000]
Loyalty_Adopter * Post -253.8 (20.82) -52.53 (19.43) -210.1 (19.03) -11.43 (18.60) $9.51\times 10^{-5}$ (0.0115) 0.0017 (0.0114) -86.53 (7.352) -15.22 (6.835)
[0.000] [0.009] [0.000] [0.546] [0.994] [0.888] [0.000] [0.024]
Loyalty program controls Yes Yes Yes Yes Yes Yes Yes Yes
Fixed-effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 197,658 197,658 197,658 197,658 110,339 110,339 197,658 197,658
R2 0.05633 0.47155 0.05137 0.43591 0.28420 0.57046 0.04567 0.44366
Within R2 0.02694 0.01910 0.00903 0.02249
Clustered (Customer) standard-errors are presented in standard brackets, while p‑values are shown in square brackets.
E.4.4 Doubly Robust DID for Covid Analysis
Given the potential concern of parallel trend assumption under COVID-19 analysis, we seek an alternative identification assumption and conduct a doubly robust DID for the covid analysis. We now show that the main conclusions remain consistent by using the doubly robust DID (Sant’Anna and Zhao 2020). We provide the ATT estimates for each metric based on the package did in R provided by Callaway and Sant’Anna (2021). Table A41 presents the doubly robust DID estimates for all metrics. Figure A22 and A23 below present the event study figure under doubly robust DID results for main metrics (i.e., spend and profit) and other metrics, respectively. We can see that the parallel pre-trends are also well satisfied, conditional on customer demographics and pre-adoption purchase variables. The estimation results are consistent with the main findings.
Table A41: Doubly Robust DID – Covid Adopters vs. Organic Adopters
DVs ATT SE T-stat P-value CI- CI+
Spend -5.963 7.814 -0.763 0.445 -21.279 9.353
Offline Spend 26.893 8.339 3.225 0.001 10.549 43.238
Share of Offline Spend 0.083 0.004 20.550 0.000 0.075 0.091
Profit 11.623 2.942 3.951 0.000 5.856 17.390
Quantity 0.246 0.324 0.760 0.447 -0.389 0.881
Orders -0.034 0.013 -2.579 0.010 -0.059 -0.008
Unique Items 0.079 0.042 1.875 0.061 -0.004 0.162
Unique Brands 0.054 0.028 1.907 0.057 -0.002 0.110
Unique Subcategories 0.047 0.022 2.141 0.032 0.004 0.091
Unique Categories 0.028 0.017 1.719 0.086 -0.004 0.061
Offline Spend 26.893 8.339 3.225 0.001 10.549 43.238
Offline Quantity 0.367 0.234 1.570 0.116 -0.091 0.824
Offline Orders 0.039 0.010 3.783 0.000 0.019 0.060
Offline Unique Items 0.178 0.037 4.778 0.000 0.105 0.251
Offline Unique Brands 0.134 0.026 5.108 0.000 0.082 0.185
Offline Unique Subcategories 0.106 0.021 5.051 0.000 0.065 0.147
Offline Unique Categories 0.074 0.016 4.617 0.000 0.042 0.105
Figure A22: Doubly Robust Event Study for the Main Metrics – Covid Adopters vs. Organic Adopters
(a) Spend
(b) Profit
Figure A23: Doubly Robust Event Study for the Other Total Metrics – Covid Adopters vs. Organic Adopters. The pretreatment outcomes of all offline metrics are the same as the total metrics, since all customers only shop offline.
(a) Quantity
(b) Orders
(c) Unique Items
(d) Unique Brands
(e) Unique Subcategories
(f) Unique Categories
E.5 Appendix to Alternate definitions for the organic adopter and covid adopter cohorts
E.5.1 Appendix to Covid Analysis - Re-defining the Organic Adopter Cohort
In this section, we provide a robustness check by redefining organic adopters as those customers who were offline-only before 2019-10-31 and made their first online purchase between 2019-11-01 to 2020-02-29. The definition of covid adopters remains the same. Under this new definition, we continue to have 33,246 covid adopters and a larger number of 25,958 organic adopters. As before, we exclude the periods used for defining these two cohorts (2019-11-01 to 2020-04-30) and re-run the same set of DID models as in the main analysis. Table A42 presents the DID results between covid adopters and organic adopters. Table A43 and A44 present DID results for other total and offline metrics, respectively. We find that all the results are consistent with the main analysis.
Table A42: Covid Adopters vs. Organic Adopters – Redefining Organic Adopter Group
DVs: Spend Offline Spend Share of Offline Spend Profit
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 450.4 (3.916) 450.4 (3.916) 1.000 (3.24e-16) 151.2 (1.288)
[0.000] [0.000] [0.000] [0.000]
Covid_Adopter 27.56 (5.291) 27.56 (5.291) 0.000 (0.000) 9.284 (1.727)
[0.000] [0.000] [0.000] [0.000]
Post 67.35 (4.220) -214.8 (3.423) -0.4980 (0.0024) 14.89 (1.397)
[0.000] [0.000] [0.000] [0.000]
Covid_Adopter * Post -4.580 (5.459) 10.03 (5.944) 34.76 (4.587) 44.76 (4.773) 0.0682 (0.0032) 0.0580 (0.0034) 8.835 (1.839) 13.20 (2.082)
[0.402] [0.091] [0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Fixed-effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 2,178,690 2,178,690 2,178,690 2,178,690 1,153,061 1,153,061 2,178,690 2,178,690
R2 0.00079 0.39039 0.01289 0.36518 0.16191 0.50935 0.00099 0.30924
Within R2 5.66e-06 0.00021 0.00108 6.31e-05
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A43: Covid Adopters vs. Organic Adopters (Other Total Metrics) – Redefining Organic Adopter Group
DVs: Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Covid_Adopter * Post 0.211 (0.090) -0.007 (0.008) 0.090 (0.024) 0.077 (0.017) 0.073 (0.013) 0.059 (0.010)
[0.019] [0.382] [0.000] [0.000] [0.000] [0.000]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 2,178,690 2,178,690 2,178,690 2,178,690 2,178,690 2,178,690
R2 0.46518 0.37879 0.37581 0.34501 0.32459 0.31320
Within R2 1.37e-05 1.45e-06 2.86e-05 3.97e-05 5.59e-05 6.20e-05
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A44: Covid Adopters vs. Organic Adopters (Other Offline Metrics) – Redefining Organic Adopter Group
DVs: Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Covid_Adopter * Post 0.113 (0.076) 0.036 (0.007) 0.115 (0.024) 0.106 (0.017) 0.091 (0.013) 0.073 (0.010)
[0.137] [0.000] [0.000] [0.000] [0.000] [0.000]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 2,178,690 2,178,690 2,178,690 2,178,690 2,178,690 2,178,690
R2 0.44370 0.36979 0.35206 0.33015 0.31714 0.31261
Within R2 7.65e-06 6.34e-05 5.52e-05 8.80e-05 9.95e-05 1.10e-04
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
E.5.2 Appendix to Covid Analysis - Re-defining the Covid Adopter Cohort
In this section, we provide an alternative definition of covid adopters cohort for the covid analysis. We provide a robustness check by redefining covid adopters as those customers who were offline-only before 2019-12-31 and made their first online purchase between 2020-05-01 and 2020-06-30 (i.e., two months later than the exact onset of COVID-19). The definition of organic adopters remains the same as the definition in §5.2.1. Under this definition, we continue to have 12,799 organic adopters but now have 27,448 covid adopters. As before, we exclude the periods used for defining these two cohorts (i.e., 2020-01-01 to 2020-06-30), re-run the same set of DID models as in the main analysis, and present the results in Table A45. Table A46 and A47 present DID results for other total and offline metrics. We find that all results are consistent with the main analysis.
Table A45: Covid Adopters vs. Organic Adopters – Re‐defining Covid Adopters Group
DVs: Spend Offline Spend Share of Offline Spend Profit
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 432.8 (2.87) 432.8 (2.17) 1.00 (0.002) 145.7 (1.10)
[0.000] [0.000] [0.000] [0.000]
Covid_Adopter -30.9 (3.42) -30.9 (2.59) 0.000 (0.000) -9.07 (1.31)
[0.000] [0.000] [0.000] [0.000]
Post 88.8 (3.23) -199.3 (2.45) -0.510 (0.002) 20.5 (1.23)
[0.000] [0.000] [0.000] [0.000]
Covid_Adopter * Post 3.90 (3.86) 7.16 (6.69) 70.1 (2.93) 70.2 (5.50) 0.086 (0.002) 0.078 (0.004) 9.26 (1.48) 9.36 (2.31)
[0.323] [0.286] [0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Fixed‐effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 1,461,440 1,461,440 1,461,440 1,461,440 744,927 744,927 1,461,440 1,461,440
R2 0.00205 0.37338 0.00919 0.36248 0.18744 0.51561 0.00112 0.27316
Within R2 3.38e-06 0.00055 0.00202 3.41e-05
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A46: Covid Adopters vs. Organic Adopters – Other Total Metrics (Re‐defining Covid Adopters Group)
DVs: Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Covid_Adopter * Post 0.085 (0.103) 0.001 (0.011) 0.154 (0.029) 0.125 (0.021) 0.111 (0.016) 0.083 (0.013)
[0.411] [0.927] [0.000] [0.000] [0.000] [0.000]
Fixed‐effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 1,461,440 1,461,440 1,461,440 1,461,440 1,461,440 1,461,440
R2 0.46377 0.36292 0.35646 0.32818 0.30849 0.29795
Within R2 3.11e-06 5.72e-08 9.05e-05 0.00011 0.00013 0.00012
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A47: Covid Adopters vs. Organic Adopters – Other Offline Metrics (Re‐defining Covid Adopters Group)
DVs: Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6)
Variables
Covid_Adopter * Post 0.473 (0.086) 0.092 (0.009) 0.299 (0.028) 0.240 (0.020) 0.201 (0.016) 0.154 (0.012)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Fixed‐effects
Customer Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 1,461,440 1,461,440 1,461,440 1,461,440 1,461,440 1,461,440
R2 0.44248 0.35480 0.33845 0.31801 0.30476 0.30009
Within R2 0.00016 0.00044 0.00039 0.00047 0.00050 0.00049
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
E.5.3 Appendix to Covid Analysis - Comparing Covid Adopters (March/April 2020) vs. Later Covid Adopters (May/June 2020)
Table A48 compares covid adopters (i.e., adopt during March/April 2020) and later covid adopters (i.e., adopt during May/June 2020). Although the later covid adopters spend more than the covid adopters by 9.96 MCU, it is not large in magnitude with a p-value of 0.048. So we conclude there is no major difference in post-adoption spend, share of offline spend (i.e., channel utilization), and profit.
Table A48: Covid Adopters (March/April) vs. Later Covid Adopters (May/June)
DVs: Spend Offline Spend Share of Offline Spend Profit
Model: (1) (2) (3) (4) (5) (6) (7) (8)
Variables
Constant 465.7 (1.690) 465.7 (1.340) 1.00 (0.0010) 156.7 (0.594)
[0.000] [0.000] [0.000] [0.000]
Later_Covid_Adopter -63.7 (2.530) -63.7 (2.010) 0.000 (0.000) -20.0 (0.890)
[0.000] [0.000] [0.000] [0.000]
Post 78.3 (1.930) -164.0 (1.530) -0.423 (0.001) 27.9 (0.676)
[0.000] [0.000] [0.000] [0.000]
Later_Covid_Adopter * Post 14.4 (2.880) 9.96 (5.040) 34.9 (2.290) 30.4 (4.270) -0.001 (0.002) -0.0002 (0.003) 1.82 (1.010) 0.070 (1.770)
[0.000] [0.048] [0.000] [0.000] [0.617] [0.947] [0.072] [0.968]
Fixed-effects
Customer Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes
Fit statistics
Observations 2,240,950 2,240,950 2,240,950 2,240,950 1,170,421 1,170,421 2,240,950 2,240,950
R2 0.00241 0.39921 0.00832 0.36673 0.16862 0.50555 0.00231 0.35185
Within R2 8.12e-06 0.00011 9.28e-09 3.00e-09
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
E.6 Appendix to Using Natural Log Scale Dependent Variables
In this section, we provide a robustness check for using natural log transformed scale of the dependent variables (i.e., ln(dependent variable + 1)), and conclude that the direction of effects are consistent with the main finding. Note that the profit variable can take both positive and negative values, so we do not consider natural log transformation of profit. First, we present the result for spend between adopters and offline-only customers in Table A49, and show that the results are consistent with Table 3.
Table A49: DID Analysis - Spend (in Natural Log Scale) (Adopters vs. Offline-only)
Study: Covid Black Friday Loyalty Program
DV ln(Spend + 1)
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 2.824 (0.0026) 2.946 (0.0027) 2.898 (0.0029)
[0.000] [0.000] [0.000]
Adopter 0.7498 (0.0104) 0.7398 (0.0320) 0.4664 (0.0146)
[0.000] [0.000] [0.000]
Post -0.9089 (0.0025) -1.048 (0.0023) -0.4705 (0.0030)
[0.000] [0.000] [0.000]
Adopter * Post 0.7266 (0.0108) 0.6883 (0.0104) 0.6734 (0.0316) 0.5742 (0.0302) 0.3381 (0.0197) 0.6099 (0.0190)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 22,204,608 22,204,608 9,505,999 9,505,999 4,403,449 4,403,449
R2 0.02929 0.30744 0.03104 0.36627 0.01150 0.29925
Within R2 0.00087 0.00010 0.00208
Clustered (Customer) standard-errors are presented in standard brackets, while p-values are shown in square brackets.
Next, we present DID results for all metrics of comparing event adopters and organic adopters by three events.
Covid Analysis
Table A50: DID Analysis of Total Metrics (Natural Log Scale) – Covid Adopters vs. Organic Adopters
DVs: Spend Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6) (7)
Variables
Covid_Adopter * Post 0.008 (0.023) 0.010 (0.008) -0.007 (0.004) 0.010 (0.006) 0.010 (0.006) 0.011 (0.005) 0.010 (0.005)
[0.728] [0.211] [0.080] [0.096] [0.096] [0.028] [0.046]
Fixed‐effects
Customer Yes Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949
R2 0.28677 0.38486 0.34049 0.33495 0.32103 0.31157 0.30123
Within R2 $2.16\times 10^{-7}$ $4.09\times 10^{-6}$ $6.85\times 10^{-6}$ $5.62\times 10^{-6}$ $7.22\times 10^{-6}$ $86\times 10^{-6} 9.9$$94\times 10^{-6}$
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A51: DID Analysis of Offline Metrics (Natural Log Scale) – Covid Adopters vs. Organic Adopters
DVs: Spend Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6) (7)
Variables
Covid_Adopter * Post 0.148 (0.021) 0.025 (0.008) 0.017 (0.004) 0.029 (0.006) 0.029 (0.005) 0.027 (0.005) 0.025 (0.005)
[0.000] [0.002] [0.000] [0.000] [0.000] [0.000] [0.000]
Fixed‐effects
Customer Yes Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949 1,783,949
R2 0.30476 0.36449 0.34627 0.33410 0.32530 0.31930 0.31278
Within R2 $9.57\times 10^{-5}$ $2.79\times 10^{-5}$ $5.60\times 10^{-5}$ $5.29\times 10^{-5}$ $6.33\times 10^{-5}$ $6.46\times 10^{-5}$ $6.52\times 10^{-5}$
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A50 and A51 provide DID results between covid adopters and organic adopters with natural log scale dependent variables for total and offline metrics, respectively.
Black Friday Analysis
Table A52 and A53 provide DID results between Black Friday and organic adopters with natural log scale dependent variables for total and offline metrics, respectively.
Table A52: DID Analysis of Total Metrics (Natural Log Scale) – Black Friday Adopters vs. Organic Adopters
DVs: Spend Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6) (7)
Variables
BlackFriday_Adopter * Post -0.268 (0.062) -0.084 (0.021) -0.050 (0.011) -0.060 (0.017) -0.050 (0.015) -0.046 (0.014) -0.043 (0.013)
[0.000] [0.000] [0.000] [0.001] [0.001] [0.001] [0.001]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 76,842 76,842 76,842 76,842 76,842 76,842 76,842
R2 0.32002 0.41568 0.38717 0.38110 0.36995 0.35915 0.34685
Within R2 0.00049 0.00050 0.00069 0.00036 0.00030 0.00029 0.00031
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A53: DID Analysis of Offline Metrics (Natural Log Scale) – Black Friday Adopters vs. Organic Adopters
DVs: Spend Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6) (7)
Variables
BlackFriday_Adopter * Post 0.082 (0.059) 0.003 (0.020) 0.005 (0.010) 0.009 (0.017) 0.013 (0.015) 0.014 (0.014) 0.013 (0.013)
[0.170] [0.882] [0.639] [0.582] [0.383] [0.308] [0.325]
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 76,842 76,842 76,842 76,842 76,842 76,842 76,842
R2 0.35766 0.42251 0.41348 0.39672 0.38879 0.38097 0.37182
Within R2 $5.17\times 10^{-5}$ $5.66\times 10^{-7}$ $6.82\times 10^{-6}$ $9.35\times 10^{-6}$ $2.27\times 10^{-5}$ $2.87\times 10^{-5}$ $3.05\times 10^{-5}$
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Loyalty Program Analysis
Table A54 and A55 provide DID results between loyalty program and organic adopters with natural log scale dependent variables for total and offline metrics, respectively.
Table A54: DID Analysis of Total Metrics (Natural Log Scale) – Loyalty Program Adopters vs. Organic Adopters
DVs: Spend Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6) (7)
Variables
Loyalty_Adopter * Post -0.151 (0.064) -0.029 (0.021) -0.022 (0.012) -0.025 (0.017) -0.025 (0.015) -0.021 (0.014) -0.017 (0.012)
[0.017] [0.158] [0.077] [0.144] [0.101] [0.146] [0.173]
Loyalty program controls Yes Yes Yes Yes Yes Yes Yes
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 197,658 197,658 197,658 197,658 197,658 197,658 197,658
R2 0.31155 0.42241 0.39069 0.37743 0.35922 0.34595 0.33365
Within R2 0.02144 0.02661 0.02799 0.02767 0.02723 0.02660 0.02525
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
Table A55: DID Analysis of Offline Metrics (Natural Log Scale) – Loyalty Program Adopters vs. Organic Adopters
DVs: Spend Quantity Orders Unique Items Unique Brands Subcategories Categories
Model: (1) (2) (3) (4) (5) (6) (7)
Variables
Loyalty_Adopter * Post -0.029 (0.065) 0.011 (0.021) 0.005 (0.011) 0.003 (0.017) 0.002 (0.015) 0.003 (0.014) 0.005 (0.012)
[0.662] [0.606] [0.665] [0.866] [0.890] [0.826] [0.696]
Loyalty program controls Yes Yes Yes Yes Yes Yes Yes
Fixed-effects
Customer Yes Yes Yes Yes Yes Yes Yes
YearMonth Yes Yes Yes Yes Yes Yes Yes
Fit statistics
Observations 197,658 197,658 197,658 197,658 197,658 197,658 197,658
R2 0.32577 0.41498 0.38925 0.38051 0.36537 0.35334 0.34367
Within R2 0.02343 0.02335 0.02347 0.02427 0.02436 0.02433 0.02379
Clustered (Customer) standard‐errors are presented in standard brackets, while p‑values are shown in square brackets.
E.7 Appendix to Handling Zero-Inflated Dependent Variables
In this section, we outline the procedure and results of the two-part hurdle model and the estimation results. The two-part hurdle model includes two separate estimation steps:
  • $\bullet$
    Step 1: Estimate a linear probability model with outcome variable $\mathbbm{I}\{Spend_{it}\}$ with all independent variables specified in Equation (1) and (2) when comparing adopters vs. offline-only, and independent variables specified in Equation (3) and (4) when comparing event adopters vs. organic adopters;
  • $\bullet$
    Step 2: Estimate Equation (1) and (2) conditional on customers making non-zero spend when comparing adopters vs. offline-only, and independent variables specified in Equation (3) and (4) when comparing event adopters vs. organic adopters.
Next, we present the estimation results for adopters vs. offline-only customers and for each event Adopters vs. organic adopters.
Adopters vs. Offline-only
Table A56 presents the estimation result of linear probability model of $\mathbbm{I}\{Spend_{it}\}$ between adopters and offline-only customers. Table A57 presents the DID estimation results conditional on customer-month with non-zero spend between adopters and offline-only customers.
Table A56: Linear Probability Model with DID Specifications – Adopters vs. Offline-only
Study: Covid Black Friday Loyalty Program
DV $\mathbbm{I}\{Spend_{it}\}$
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 0.4792 (0.0004) 0.5002 (0.0004) 0.4766 (0.0004)
[0.000] [0.000] [0.000]
Adopter 0.0947 (0.0015) 0.0942 (0.0047) 0.0586 (0.0021)
[0.000] [0.000] [0.000]
Post -0.1627 (0.0004) -0.1810 (0.0004) -0.0786 (0.0005)
[0.000] [0.000] [0.000]
Adopter * Post 0.1098 (0.0016) 0.1033 (0.0016) 0.1085 (0.0048) 0.0910 (0.0046) 0.0570 (0.0030) 0.0933 (0.0029)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 22,204,608 22,204,608 9,505,999 9,505,999 4,403,449 4,403,449
R2 0.02892 0.28000 0.03368 0.34101 0.01031 0.26334
Within R2 0.00073 $9.3\times 10^{-5}$ 0.00156
Clustered (Customer) standard-errors are presented in standard brackets, while p-values are shown in square brackets.
Table A57: DID Estimates for Spend (Conditional on Purchase) – Adopters vs. Offline-only
Study: Covid Black Friday Loyalty Program
DV Spend
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 584.6 (0.9550) 583.1 (0.9740) 697.9 (1.448)
[0.000] [0.000] [0.000]
Adopter 212.2 (4.176) 206.7 (12.83) 154.2 (7.012)
[0.000] [0.000] [0.000]
Post 101.9 (1.061) 26.81 (0.8289) 8.823 (1.252)
[0.000] [0.000] [0.000]
Adopter * Post 127.7 (4.204) 135.4 (3.953) 52.78 (10.73) 65.40 (9.385) -52.10 (8.243) 50.28 (6.334)
[0.000] [0.000] [0.000] [0.000] [0.000] [0.000]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 8,187,536 8,187,536 3,744,951 3,744,951 1,939,763 1,939,763
R2 0.02034 0.47201 0.00178 0.50834 0.00501 0.62202
Within R2 0.00084 $4.37\times 10^{-5}$ 0.00150
Clustered (Customer) standard-errors are presented in standard brackets, while p-values are shown in square brackets.
Event Adopters vs. Organic Adopters
Table A58 presents the estimation result of linear probability model of $\mathbbm{I}\{Spend_{it}\}$ between event adopters and organic adopters. Table A59 presents the DID estimation results conditional on customer-month with non-zero spend between event adopters and organic adopters.
Table A58: Linear Probability Model with DID Specifications – Event Adopters vs. Organic Adopters
Study: Covid Black Friday Loyalty Program
DVs: $\mathbbm{I}\{Spend_{it}\}$
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 0.5557 (0.0029) 0.5889 (0.0057) 0.5297 (0.0022)
[0.000] [0.000] [0.000]
Covid_Adopter 0.0244 (0.0033)
[0.000]
Post -0.0451 (0.0031) -0.0569 (0.0060) -0.0162 (0.0030)
[0.000] [0.000] [0.000]
Covid_Adopter * Post -0.0100 (0.0036) 0.0039 (0.0034)
[0.005] [0.251]
BlackFriday_Adopter 0.0171 (0.0098)
[0.073]
BlackFriday_Adopter * Post -0.0488 (0.0100) -0.0361 (0.0095)
[0.000] [0.000]
Loyalty_Adopter 0.0634 (0.0076)
[0.000]
Loyalty_Adopter * Post -0.1187 (0.0093) -0.0202 (0.0095)
[0.000] [0.032]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Observations 1,783,949 1,783,949 76,842 76,842 197,658 197,658
R2 0.00210 0.25272 0.00569 0.28908 0.03886 0.27038
Within R2 $2.36\times 10^{-6}$ 0.00036 0.01574
Clustered (Customer) standard-errors are presented in standard brackets, while p-values are shown in square brackets.
Table A59: DID Estimates for Spend (Conditional on Purchase) – Event Adopters vs. Organic Adopters
Study: Covid Black Friday Loyalty Program
DVs: Spend
Model: (1) (2) (3) (4) (5) (6)
Variables
Constant 778.8 (7.845) 779.6 (14.49) 839.6 (7.031)
[0.000] [0.000] [0.000]
Covid_Adopter 23.90 (9.168)
[0.009]
Post 237.0 (7.804) 103.2 (13.31) -30.62 (8.003)
[0.000] [0.000] [0.000]
Covid_Adopter * Post -9.354 (9.147) -3.982 (8.373)
[0.316] [0.623]
BlackFriday_Adopter 30.84 (29.04)
[0.288]
BlackFriday_Adopter * Post -75.22 (21.45) -64.19 (19.68)
[0.000] [0.001]
Loyalty_Adopter 132.3 (27.63)
[0.000]
Loyalty_Adopter * Post -232.3 (27.67) -36.03 (23.92)
[0.000] [0.126]
Loyalty program controls Yes Yes
Fixed-effects
Customer Yes Yes Yes
YearMonth Yes Yes Yes
Fit statistics
Standard-Errors Customer
Observations 949,448 949,448 42,299 42,299 110,339 110,339
R2 0.00803 0.44326 0.00248 0.45643 0.02928 0.56886
Within R2 $7.04\times 10^{-7}$ 0.00047 0.01280
Clustered (Customer) standard-errors are presented in standard brackets, while p-values are shown in square brackets.
[…]